A Complete Bibliography of Publications in Synthese, 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.02

Title word cross-reference

#30 [61].

∗ [548, 348, 551]. ^{++} [126]. ^{e} [181]. E [398]. Γ [398]. ω [502]. Π_{1}^{1} [1045]. T [651].

-Admissibilty [398]. -complex [1045]. -Inconsistency [502]. -Maximin [398].

20th [309].

=Woods [71].

abduction [976]. Abductive [262]. Abilities [155]. ability [960]. Abnormal [62]. Abnormality [942, 1045]. aboutness [842]. Abstract [270, 987]. Abstracting [657]. Absurdity [608]. Acceptance [476, 159].

Acceptibility [574]. access [840]. accessible [778]. Accidentally [344]. Account [630, 704, 189, 508, 709, 780, 703, 696, 889]. Acquaintance [771]. across [184]. Action [614, 715, 143, 800]. Actions [804, 1101]. activation [723]. Active [144]. activities [677]. Activity [423]. Acts [136, 473, 147, 1006]. Acts-Principle [473]. Adam [126]. Adaptive [62, 815, 1045, 1099, 1002]. Addressing [965]. adequacy [960]. Adequate [971]. adjointness [1091]. Admissibilty [398]. aesthetic [993]. affect [1006]. Affective [50]. Affine [80]. after [1084, 338]. again [311, 977]. Against [1074, 139, 627, 408, 249, 141, 547, 1125, 1131]. agency [804]. Agent [378, 404, 380]. agentive [861]. Agents [296, 1000, 1109]. Aggregating [406]. Aggregation [296, 938, 696]. agreements [1164]. AI [1130]. aims [247]. Albert [324]. Alethic [491]. Alethically [197]. Alexander [713]. Alfred [91]. Algebras [81]. Algorithmic [28]. algorithmics [962]. Alleged [359, 637, 811]. Alper [202]. Alternating [614, 443, 767]. Alternating-Time [614, 443]. Alternative [143, 63, 456, 963]. Amartya [397]. Ambiguity [359]. Ambit [107]. Amending [541]. Among [369]. analogue [1114]. analogy [1040]. Analysis [522, 658, 6, 372, 579, 246, 127, 273, 244, 424, 546, 892, 960]. Analytic [255, 967]. analytic/synthetic [967]. Analyticity [185, 627]. Anaphoric [712]. Android [4]. Animal [8]. Animals [20, 930]. Anomalies [408]. Anomalous [568]. Answer [525, 960]. Answers [632]. Anti [836, 830, 1044, 1172]. Anti-luck [836, 830]. anti-realism [1172]. anti-realists [1044]. Antirealist [799]. Any [321, 46, 1068]. Anymore [360]. Anything [620, 1136]. aporia [1169]. appear [812]. Appendix [305]. Application [272, 412, 69]. Applications [236, 70, 290]. Applied [121, 14, 237]. Apprehension [324]. Approach [262, 472, 694, 443, 11, 114, 567, 275, 76, 316, 1127, 1162, 1161, 950, 456]. Approached [587]. Approaches [245, 712, 490]. Appropriateness [900]. Appropriation [229, 320]. Arbitrary [615]. Archetypal [417]. architecture [861]. Arguing [272]. Argument [280, 693, 126, 470, 688, 274, 504, 798, 139, 694, 627, 319, 249, 618, 436, 354, 299, 997, 699, 1155, 1022, 1151, 696, 1131, 1021, 543, 1124]. Argument- [299]. argument-based [696]. Argumentation [118, 505]. Arguments [519, 644, 737, 1068, 1030, 1040, 1125]. Arguments* [634]. Aristotelian [108, 1028]. Aristotle [56, 277]. Arithmetic [733, 66, 542]. Arithmetical [400]. Arithmetik [541]. arrow [1090, 240, 107]. Arthur [641]. Article [60]. articulation [464]. artifact [1075]. artifactual [1095]. Artificial [1130, 119, 1118]. Ascent [846]. ascription [853, 847]. Ascriptions [532, 850]. aspects [804]. Assertability [47]. assertion [841]. assertion-candidate [841]. Assessing [893]. assessments [1028]. Assigning [898, 400]. Assumption [145]. Assumptions [449, 450]. asymmetry [1090]. attempt [940]. Attention [49]. Attitudes [284, 1160, 869, 846]. Aufbau [194]. Author [367, 389, 414, 488, 513, 539, 562, 599, 624, 649]. Authors [258, 302, 346, 368, 390, 415, 489, 563, 600, 625, 650, 514]. Autonomy [369, 104]. averse [825]. avidity [702]. Avoiding [768, 285, 894]. awareness [861]. Away [454, 615]. Axiom [571]. Axiomatizations [250, 511]. Axiomatizing [827]. Axioms [546, 80].

B [918]. Babies [4]. Back [230]. background [1022]. Backgrounds [59]. Backward [312, 1105]. Bad [1146, 1145, 1143, 797, 1044, 1144, 1147]. balance [759]. Base [705]. Based [245, 443, 612, 250, 595, 36, 1000, 1036, 1161, 950, 1017, 551, 696]. Bases [70]. basic [120]. basis [960]. Battle [653]. Baumann [821]. Bayes [893]. Bayesian [939, 783, 801, 222, 694, 873, 810, 296, 770, 431, 786, 1047]. Bayesianism [787, 959]. Be [339, 197, 9, 469, 913, 1069, 1170, 813, 738, 903]. bearing [985]. beautiful [993]. Beauty [796, 785, 384, 873, 1079, 210, 979, 870]. Bechtel [1066]. becoming [1059, 1166]. Begging [695]. beginning [559]. Begriffsschrift [292]. Behavior [6, 288]. behind [673]. Being [38]. Belief [36, 532, 549, 344, 160, 551, 728, 507, 696, 5, 377, 158, 550, 536, 159, 1000, 909, 1001, 789, 1162, 1077, 759, 825]. Belief-based [551]. beliefs [760, 972, 756, 461, 758]. Believe [46]. Believing [122, 910, 906]. Bell [949, 141]. Benacerraf [1156]. bench [671]. Benchmarks [119]. benefit [962]. Best [187, 221, 864, 463, 806]. Bets [384]. Between [308, 398, 568, 670]. Beyond [542, 1089, 1088]. BI [1052]. Bias [360]. Bickle [675]. Big [1144, 1044]. Bilking [319]. Billboards [1026]. bio [1038]. bio-cognitive [1038]. Biochemical [496]. Biochemistry [752]. biological [1014, 1095, 898, 1089]. Biology [3, 664]. Bipolar [465]. Birth [470]. bites [1044]. Bivalence [530]. Black [529]. Body [631, 105, 1116]. Boghossian [1016]. Bogus [283]. Bohmian [741]. Bohr [408]. Bois [544]. Bois-Reymonds [544]. Boltzmann [253]. Bolzano [637]. bombs [1026]. Bonjour [570, 844]. Book [255]. Bookkeeping [361]. Books [257, 979]. Boolean [1127]. Bootstrap [605]. Bostrom [470]. Boundaries [123, 356, 935]. Boundary [537, 628]. Bounds [819, 1109]. Bovens [476]. Bradley [952]. Brain [423, 150, 427, 422, 560, 208, 254, 723, 857, 859, 721]. Brain-Function [422]. Branching [223, 647, 1103, 973]. Branching-Time [647]. Breakfast [130]. Bridge [208, 254, 722]. Bridger [202]. Bridging [1108]. Brief [195, 982]. Broken [17]. brooked [1042]. Brother [1144]. Brouwer [331, 327]. Brute [509]. Brutes [277]. bug [1044]. build [857]. building [1133]. built [765]. Bundling [751]. business [973].

C [218]. C. [436]. CAD [1088]. Calculi [241]. Calculus [423, 925, 337]. Calibrating [764]. Can [307, 1069, 408, 436, 962, 9, 833, 917]. candidate [841]. Candor [501, 500]. cannot [680]. Cantorian [249]. Cao [309]. Capacities [182]. Carnap [839, 328, 194, 437, 338]. Carol [405]. Cartesian [780]. cartographic [1097]. Cartwright [660]. Case [452, 496, 353, 138, 602, 516, 208, 254, 927, 1116, 759, 1012, 821, 758]. casino [1086]. cat [672]. categorical [1125]. Categories [911, 108, 1048]. Category [582, 248]. Category-specified [582]. Causal [420, 319, 626, 622, 225, 369, 422, 198, 601, 1127, 898, 965, 964]. Causality [222]. Causally [481]. Causation [658, 322, 713, 819, 11, 223, 199, 621, 878, 1043]. Cause [455, 127, 199, 949]. cellular [674]. Century [177, 178, 309]. certainty [1042]. Ceteris [182, 1050]. chains [898]. Challenge [101, 471, 1067]. Challenges [584, 1077]. chance [1044, 1022]. Chances [220]. Change [691, 428, 432, 348, 759, 1159]. changes [1163]. Changing [934, 756]. Changizi [427]. Character [93, 840]. Characterization [379, 828]. charm [1113]. Checking [443, 244]. Cheers [207]. chemical [763]. chemistry [766, 762]. chemists [765]. Cheshire [672]. children [847]. Choice [503, 553, 396, 1161, 951]. Choices [161]. Church [729, 881]. circle [1009]. Circularity [774]. claims [965]. Clarify [140]. class [791]. Classical [165, 591, 577, 988, 1091, 75]. classical-quantum [1091]. Classifying [889]. Clauses [182]. Close [423]. Closed [319, 1107]. Closer [534]. Closure [661, 921]. coalition [1005]. coexistents [75]. Coffman [921]. Cognition [3, 819]. Cognitive [148, 108, 559, 137, 517, 23, 855, 720, 722, 1038, 856, 1155, 756, 942, 678]. Cognitively [576]. Cognitivism [915, 1157, 701]. Coherence [523, 806, 728, 436, 520, 13, 784, 811, 808, 807, 812, 896]. coherent [1033]. coherentism [810]. coherentist [809]. Cohomology [72]. Coincidence [581]. Collective [159, 800, 827]. Color [145, 22]. Colour [254, 230]. Combined [180]. come [786]. Comment [724, 214]. Comments [1149, 476, 824]. Commitment [340, 1030]. commitments [1167]. Common [941, 949, 950]. common-common-cause-type [949]. Communication [35, 1109, 1008, 1003, 824]. communications [823]. communitarianism [1016]. Commutative [38]. Compactification [78]. company [1144, 1146, 1145, 1143]. Comparative [185, 919]. comparators [861]. Compared [406]. Comparing [245, 378]. Comparison [250, 195]. Comparisons [31]. Compartment [621]. Competence [283]. Complementarity [680, 681]. Completeness [30, 87, 779]. Complex [632, 1019, 1045]. Complexity [365, 250]. Composition [55]. Compositionality [471, 98, 97, 775]. Compound [775]. Compulsion [10]. computation [723]. Computational [693, 423, 718, 610, 722]. Computationalism [729]. Computer [1119, 961, 35, 1122, 1117, 721, 1115]. computers [1089]. Computing [1113, 989]. Conceivability [662, 1168, 1157, 876]. Concept [102, 558, 364, 297, 426, 912, 1017]. Conception [294]. Concepts [173, 688, 247, 168, 20, 588, 23, 252, 1144, 1140, 900, 766, 748, 1096]. Conceptual [309, 652, 1141, 892, 1073]. Concerning [272, 160, 320, 440, 747]. Conclusiveness [457]. Concrete [270]. condition [966]. Conditional [579, 523, 189, 339]. Conditionalization [38, 1079]. Conditionals [1152, 1029, 968, 919]. Conditions [537, 211, 123, 457, 847].

Condorcet [1016]. conducive [808, 807, 812]. Confirmation [605, 770, 431, 787, 808]. Conformity [475]. Confounding [424, 965]. Conjecture [206, 71]. Conjoining [508]. conjunction [1150]. Conjunctions [769]. Connecting [36]. Connectionism [139, 152]. Connectionist [151]. connections [890]. Connectives [646]. Connexive [115]. conscious [848, 936]. Consciousness [425, 148, 103, 279, 872, 722, 1171, 937, 944, 874]. Consciousness- [425]. Consequences [161, 610]. Conservation [52]. Conservatism [419, 975, 793]. considerations [1179]. Consistency [332, 506, 337]. Consolidation [752, 173]. consonance [836]. Constitution [270, 897, 341]. Constraints [537, 843, 700]. constructing [764]. Constructions [261]. Constructive [596, 545, 80, 646, 247, 760]. Contemporary [85, 1155]. Content [528, 840, 736, 135, 279, 622, 603, 1141]. Contents [193, 259, 301, 345, 366, 413, 388, 487, 512, 538, 561, 598, 623, 648, 232]. Context [775, 872]. Contexts [705]. Contextual [472]. Contextualism [382, 163, 753, 1169, 1025]. Contra [660, 1176]. contract [865]. Contractarian [638]. Contracting [286]. Contraction [691]. Contrast [879, 398]. Contrastive [424, 891]. Contribution [409]. Control [426, 909, 1030, 853]. Controlling [847]. Controversy [140]. convection [878]. Convention [529, 651, 554]. Conventional [381]. Conventionalism [338]. Conversational [381]. Convex [398]. Conveying [37]. Cooperation [117, 1101]. Copernicanism [1018]. Copping [44]. Core [684, 847]. Corrected [12]. Correction [511]. correlations [973, 1041]. Correspondence [838, 94, 343, 1091]. Cosmic [920]. Could [913, 339]. Counterexample [377]. counterexamples [969]. Counterfactual [127]. Counterfactuals [435, 689, 732, 769]. counterparts [1071]. counterpossibles [1168]. Counting [418, 1073, 933, 1144, 1002]. Covariance [584]. Covenants [802]. Cowie [313]. Cracking [537]. Cracks [537]. Craig [984, 455, 981, 987]. Cramer [626]. credit [835, 1098]. Critical [844, 116, 427, 309, 571, 112]. Critique [105, 299, 1043]. cross [1116]. cross-model [1116]. CTM [1137]. current [674]. cybernetics [1089].

D. [936]. Darwinian [204]. Data [70, 498, 495, 1053, 966, 965]. Dates [211]. Davidson [280, 492, 916]. Debate [361, 1069, 679, 680]. Decision [691, 525, 938, 383, 552, 398, 12, 106, 792, 1106]. decision-theoretic [792]. Decision-Theoretical [691]. Decisions [548, 442, 790]. Dedekind [342, 546]. Deduction [592, 1055, 948]. Deduction* [593]. Default [518, 523, 516, 671]. Defeasible [118, 195, 809]. Defeated [730]. Defeats [635]. Defective [38]. Defence [739, 139, 189, 652, 899]. Defending [661]. Defense [727, 470, 630, 533, 95, 342, 92]. Definition [590, 93, 364, 92]. deflated [1180]. Deflating [459]. deflationary [1092]. Deflationism [96, 995]. Defusing [1032]. Democracy [207]. Democratic [632]. Demonstrative [137, 408]. Demopoulos [255]. Dempster [36]. Denominations [374]. Denoting [688]. deontologism [905]. Dependent [622]. Derivations [592, 949]. descendants [984]. Descriptions [180]. Descriptus [130]. deserve [835]. design [1017, 1089]. desire [789]. determinacy [842]. determines [1014]. Determining [1107]. Determinism [52, 1068, 1091]. Deterministic [175, 1127]. Developing [155]. Development [309, 1084]. Developmental [3, 850, 889]. Diachronic [979, 510, 54]. Diagnosis [196, 814]. Diagrammatic [437]. Dialectic [119]. Dialectical [695, 939]. Dialogical [113, 114, 76, 115]. Dialogue [117, 320, 112, 35, 120]. dictum [969]. Different [560, 1003]. digital [1074]. Dilemma [392, 799, 696]. Dilemmas [312]. Dimension [527]. Dimensional [385, 352, 464]. dimensions [974]. directed [852]. directions [964]. Disappearance [608]. Discrete [755, 1113]. discursive [696]. Discussed [119]. Discussion [116, 456]. Disjunctions [769]. Disjunctivism [1169]. Dismal [409]. disorder [465]. Dispensability [798]. disposition [479]. Dispositional [998, 946]. Dispositions [579, 189, 480, 481, 531, 858, 882]. Dispute [118]. Disputes [478]. Disputors [119]. Disquotation [448]. Dissolution [308]. Dissolve [307]. Distance [755]. Distinction [133, 753, 1128, 1008, 1011, 1007]. distinctions [1073, 967]. Distinguishing [825]. Distracted [468]. Diversity [935]. Divine [294]. Division [24, 822, 874]. DNA [735]. Do [1094, 261, 408, 1140, 557, 786]. Does [620, 42, 268, 1115, 812]. Dogmas [568]. Doing [720]. Domains [537, 108, 607, 75]. Don’t [129, 220, 835]. Doom [785]. Doomsday [126, 470, 299]. Double [528, 1148]. doubt [1042]. Doubts [196, 163]. down [834]. Downward [878]. Doxastic [340, 754, 907, 909, 757, 828]. dream [839]. Dretske [467]. Drivers [468]. dual [678]. Duality [647, 723]. Duplication [145]. Dutch [979]. Dynamic [750, 177, 1103, 757, 712, 118, 550, 531, 1001, 1108, 1163, 1102, 1004]. dynamic-epistemic [1163]. Dynamical [526]. Dynamics [170, 202, 572, 760, 1164, 756, 1006].

E-Type [712]. early [952, 255]. Earth [43, 634, 44]. easy [1098]. Economics [409, 800]. Economy [370]. ed [255]. editor [805, 1062, 421, 717]. Editorial [303, 1111, 999, 142, 347, 564, 725, 862, 375]. Editors [1]. Effect [199, 863]. Effectiveness [69, 767]. Effects [289]. Ehrenfest [373]. Eight [119]. Eklund [214]. Elegy [688]. Elementary [236, 242, 77]. Elephant [1137]. eliminate [820]. Elimination [982, 677, 963]. Embedded [819]. Emergence [496, 684, 685, 766, 153, 54, 679, 664, 669, 680, 670, 665, 668, 678]. emergence-reduction [679, 680]. emergent [1037]. Emergentism [713, 671]. emotion [889]. emotions [1076]. Empirical [1128, 658, 170, 495, 819, 299, 1116, 1083, 1080, 1084]. Empiricism [508, 46]. emulation [858]. End [420, 826]. Endedness [297]. Ends [533, 926]. enduring [1055]. Energy [52, 1090, 859, 741]. engineering [1081, 1082]. ensure [917]. Entailments [184]. Enterprise [402]. Enthymemes [505]. Entities [307]. Entitlement [794, 1174, 1175, 1177]. Entrenchment [691, 14]. entropic [1090]. Enumerative [93]. Envelope [494]. Environment [2, 622, 1014, 1107]. Environment-Dependent [622]. environmental [1087]. Environments [300]. Epistemic [376, 9, 632, 750, 110, 533, 773, 443, 14, 926, 286, 753, 439, 285, 689, 419, 695, 919, 316, 947, 1082, 1108, 348, 844, 832, 1103, 1057, 975, 793, 1059, 1032, 1163, 1102, 957, 758]. Epistemological [429, 766, 1130]. Epistemologically [560]. epistemologies [836]. epistemologists [834]. Epistemology [783, 774, 324, 4, 17, 638, 5, 947, 61, 905, 1040, 699, 1080, 830]. EPR [1041]. EPR-like [1041]. Epsilon [337]. Equality [68]. Equilibria [57]. Equivalence [635, 799]. Equivalent [896]. Erdos [71]. Erdos-=Woods [71]. Ernst [370]. Erotetic [276]. Erratum [573, 656, 692]. Error [963, 964, 1013, 1171, 956, 961]. Error-statistical [963, 961]. errors [965]. Escape [261]. Essay [427, 186]. Essays [255, 981]. Essence [294]. Essential [1091, 928]. Essentialism [482, 483, 204]. Establishing [71]. Ethical [498]. Ethics [416, 909, 1077]. Euclid [1042]. Euclidean [272]. Evaluating [939, 1064, 803]. Evans [1071]. Eve [126]. Events [435, 224, 631]. everything [835]. Evidence [574, 442, 1154, 1128, 1031, 1064]. evidential [968]. Evolution [2, 317, 517, 381]. Evolution-Theoretic [517]. Evolutionary [526, 380, 1034]. exactly [913]. Exactness [980]. Examination [750, 112]. example [741]. exaptation [855]. except [1136]. Exclusion [369, 1093, 1177]. Existence [739, 1042, 1126, 1117]. Existential [56, 450, 1070]. Experience [411, 228, 144, 468, 146]. Experiences [38]. Experiment [412, 140]. Experimental [518, 735, 962, 1012, 1067, 965]. Experiments [188, 1119, 1115]. Expertise [58]. Explain [495, 719]. Explained [915]. Explanation [452, 453, 630, 187, 221, 172, 451, 458, 454, 559, 495, 622, 457, 186, 200, 864, 891, 806, 1130, 718]. Explanation-Seeking [457]. Explanations [456, 1130]. Explanatoriness [455]. Explanatory [171, 901, 13, 208, 254, 897, 1118, 894]. Explicating [675, 674]. Explicativism [637]. Exploration [556]. Explorations [112]. expression [746]. Expressive [30]. expressivity [1003]. Extend [114]. Extending [542, 1102]. extensions [962]. Externalism [690, 278, 884, 570, 885, 840, 887, 886]. Extrinsic [374]. Extruding [269].

faces [990]. Fact [777]. Factors [424]. Facts [425, 509]. Factualized [638]. Faculties [410]. faculty [1157]. Failure [52]. Fake [199]. Fallacies [694, 1056]. Fallacious [394, 1016]. Fallacy [394, 162, 729, 1150]. Farbung [349]. Fast [235]. Favorable [300]. Feasibility [227, 767]. Federalism [681]. Feet [427]. fellows [1019]. Feminists [360]. Fiction [420, 705]. Field [293, 304, 309, 652]. Figures [272]. Finding [677]. Findings [1083]. Fine [504, 141, 1095]. fine-grained [1095]. Finitism [337]. First [213, 639, 423, 945, 427]. First-Order [213]. First-Person [639, 945]. Fit [601]. fitness [1014]. Fixed [241]. Fixed-Point [241]. Fleas [373]. flow [1090, 1060]. fluid [878]. Focus [1149]. Fodor [1141, 23]. folding [764]. Folk [491, 992]. follow [1083]. following [1179]. Foole [312]. forbid [1094].

Foreword [609, 1166]. forgetting [1110]. Form [228]. Formal [14, 234, 273, 635, 205, 79, 744]. Formalization [733, 1076, 971]. Formalizing [950]. formally [969]. Formation [138, 377]. Forms [34, 417, 1037]. formulation [958]. Foundation [590, 238]. Foundational [972]. foundationalism [877]. Foundations [548, 234, 638, 517, 740, 743, 1073, 348, 937, 993, 90]. foundherentism [877]. Four [478, 974]. Fraassen [411]. frame [665]. Framework [554, 431, 1084, 1047]. Framing [289, 383, 938]. framings [1083]. Frank [714]. Fréchet [270]. Free [503, 859, 553, 80]. Free-energy [859]. freedom [907]. Frege [255, 540, 180, 541, 636, 255, 571, 1046, 543]. Fregean [328, 1146, 542, 1148]. Frequencies [221]. Function [291, 183, 422, 857, 894]. Functional [188, 854, 668, 855, 247, 1095]. Functionalism [491]. Functionality [372]. Functionalization [490]. Functions [36, 292, 329, 1017, 195, 898, 1161, 1075]. Fundamental [305, 377]. Fundamentalism [660]. Fundamentality [305]. Funny [973]. Further [112]. Future [629, 785, 860, 1028]. Fuzzy [275, 13].

Gaifman [401]. Galileo [269]. Game [34, 416, 655, 380, 61, 1162, 954]. game-theoretic [954]. game-theoretical [61]. Games [379, 548, 242, 506, 1163, 941, 1004]. Gap [208, 254, 897]. Gaps [263]. Gavagai [977]. General [482, 584, 236, 590, 428, 1058, 93, 1037, 1017, 665, 1131]. Generalization [387, 953]. Generalized [482, 452, 1161, 1143]. generalizing [460]. genomics [670]. Gentzen [336]. genuine [914]. Genus [186]. geometrical [1090]. Geometry [272, 238, 250, 80, 1042, 924, 511]. get [833]. gets [675]. Gettier [382, 438]. Giving [615]. Global [163, 1090, 1129, 1087]. GNY [615]. goal [760]. God [294]. Gödel [839, 328, 252]. Good [43, 357, 1068, 797, 1147, 910, 993]. Gott [299]. Graded [551]. Grades [902]. grained [1095]. graphic [954]. Gravitation [295]. Gray [529, 688]. Green [635]. Gricean [133]. Group [69, 296]. Growing [235]. Grows [3]. Growth [434]. Grue [635]. Grundgedanke [156]. Grundgesetze [541]. Guest [805]. Guide [617].

H [1042]. Haack [877]. Hall [927, 821]. Halonen [456]. Handling [518]. Harmonious [984]. Harmonising [948]. Harré [681, 680]. Harvard [714]. Haugeland [341]. Hawthorne [476]. head [1134]. Heap [710]. Hégélien [181]. Heidegger [887]. Heisenberg [408]. Hell [286, 285]. Henrich [936]. Here [359]. Heritage [328]. Heuristic [91]. heuristics [676]. Hidden [77]. Hierarchies [447]. Higher [245, 713, 279, 577]. Higher-Order [245, 279]. Hilary [21]. Hilbert [540, 544, 1126, 330, 332, 333, 334, 545, 337]. Hintikka [395, 213, 456]. Hintikka-approach [456]. Historical [569, 732, 120]. History [24, 822, 642, 237, 647, 982]. Hofmann [926]. Hold [620]. Holism [170]. Homeless [576]. honest [676]. Honesty [499]. Honey [43]. Honor [255, 981]. hoping [1105]. Horgan [465, 870]. hot [1121]. how-questions [1020]. Howard [403]. Human [715, 519, 520, 663, 1019]. Hume [1144, 698, 751, 699, 702, 701, 571, 700]. Humean [969, 703]. Husserl [87, 229, 779, 341, 228, 887, 884, 327]. Husserlian [722]. Hybrid [641, 796, 1163]. hylozoism [1089]. Hyperbolic [250, 511]. Hypotheses [425, 430, 963]. Hypothesis [735, 920].

IBE [1047]. Idea [157]. idealisation [1050]. idealism [886]. Idealization [1112, 362]. Ideas [423, 65, 684]. Identical [418]. Identity [278, 224, 32, 105, 158, 208, 254, 1071, 742, 932, 930]. Idiolectical [690]. If [611, 765, 1052, 90]. If-Logic [611]. ignorant [9]. II [1175]. III [1175]. I’ll [46]. Illuminates [173]. illusion [848, 944]. imagery [888]. Imagination [849]. Imaginations [103]. immanence [885, 887]. Immanent [423]. Immanuel [21]. Immediate [324]. immortality [849]. immunity [1171]. Imperfect [111]. implementation [857, 998]. implication [682, 683]. Implications [293, 1082, 1091]. Implicature [521, 891, 1012]. Implicit [364, 674, 1030]. Imply [42]. Import [56]. Importance [519]. Impossibility [492, 406, 811, 817]. Impossible [663, 911]. improvement [997]. Incan [243]. inclosure [875]. Include [452]. Inclusivist [567]. incoherence [465]. Incommensurability [495, 321]. Incomplete [7, 1160]. Incompleteness [396, 946]. Inconsistencies [537]. Inconsistency [502, 285]. Inconsistent [452]. Indefinitely [312]. Independence [610, 1180]. Independent [66, 552]. Indeterminacy [1071, 48, 967]. indeterminate [825]. Indeterministic [317, 601]. Index [367, 389, 414, 488, 539, 562, 599, 624, 649, 513]. Indexical [532, 1029]. Indexicals [191, 706]. Indian [165]. Indication [709]. Indifference [460, 958]. Indirect [136]. indiscriminability [980]. Indispensability [798, 215, 349, 461]. Individualism [690, 1023]. Individuation [462, 168, 279, 728]. Induction [350, 312, 162, 575, 408, 635, 1105, 699]. Inductive [474, 241, 430, 431, 731]. inequalities [949]. inexactness [980]. Inference [36, 864, 206, 187, 221, 417, 655, 422, 1104, 879, 806, 1016, 964]. Inferences [441, 1155]. Inferential [135, 1061]. infidelity [991]. Infinite [726, 295, 768, 973, 1041]. Infinitesimal [238]. Infinitesimals [273, 249]. Infinitism [351, 947, 1021]. inflation [920]. Influence [714, 330, 194]. Inform [47]. Informal [744, 694]. Information [412, 206, 37, 467, 28, 111, 1053, 1073, 1057, 1058, 1163, 1060, 1061, 1054]. informational [899, 1056]. informativeness [460]. informed [1059]. Inheritance [516, 524]. Inner [248]. Innocent [96]. Innovation [267]. inquiry [1162]. Ins [644]. Insights [533]. insofar [895]. institutional [1164]. Instructions [258, 302, 346, 368, 390, 415, 489, 514, 563, 600, 625, 650]. instrumental [1075]. integration [855, 854]. Integrity [501]. Intellectual [499, 909]. intelligent [1100]. intelligible [895]. Intensionality [32]. intention [940, 1173, 853, 761]. Intentional [263, 847]. Intentionalism [466]. Intentionality [155]. Intentions [1106, 800, 760, 1163]. inter [677]. inter-level [677]. Interaction [1038, 1100, 1107, 1037]. interactive [1036, 1106]. Interactivism [1035]. interactivist [1039]. interesting [1068]. interface [1012]. Internal [21]. Internalism [886, 701, 1131, 957]. International [24, 822, 289]. Interpolation [988, 987, 989, 984, 990]. Interpolations [981]. Interpretation [318, 304, 363, 626, 467, 342, 705, 247, 1130, 936]. Interpretations [523, 247]. Interpreted [526]. Interpreters [571]. interrogative [1162]. Interruptus [130]. Intransitivity [466]. Introducing [422, 776]. Introduction [260, 445, 697, 421, 555, 391, 616, 845, 548, 1024, 929, 1135, 219, 1051, 497, 88, 242, 586, 1165, 326, 1142, 981, 477, 167, 142, 904, 829, 515, 955, 149, 1062, 1035, 717, 1075, 805]. Introspection [851]. Introspective [1110]. Intuition [83, 84, 252]. Intuition* [82]. Intuitionism [274, 547]. Intuitionistic [110, 1054]. Intuitive [273]. Invalidity [274]. Invariance [16]. Inversion [50]. inverted [876]. Investigating [422]. investigations [1155]. Irrationality [7]. Irrealism [734]. irreducibility [723]. Irreducible [365]. Irrelevance [801]. Irresponsibility [9]. is-ought [969]. Issue [749, 1035, 783]. issues [1121, 684]. Iterated [312]. itself [454].

J. [1042]. Jaakko [395]. Jeffrey [1079, 38, 81]. Jerry [1141, 23]. Joe [126]. John [1149]. Joint [49]. Judgment [730, 938, 827, 696]. Judgments [406]. jungle [1086]. Jurors [442]. Jury [442]. Just [360]. justice [702]. Justification [818, 350, 589, 735, 595, 316, 809, 1116, 844, 972]. Justified [536]. justifies [867].

Kant [21]. Keep [1105, 1139]. Keynes [81]. Kim [1093, 685]. Kind [352]. Kinds [321, 153, 493, 1095, 912]. kinematics [789]. Klein [661]. know [835, 872, 996]. know-how [872, 996]. Knowability [653]. Knowing [19]. Knowledge [614, 1081, 470, 535, 363, 70, 434, 551, 510, 612, 184, 153, 402, 1128, 1167, 1013, 1151, 1176, 847, 942, 996, 461, 941, 628]. Knowledge-based [551]. knowledge-beliefs [461]. knowledge-how [1128]. knowledge-that [1128]. knowledge-that/knowledge-how [1128]. Known [184]. Knuth [240]. Kolak [214, 936]. Kripke [881]. Kripkenstein [751]. Kuhnian [816].

L [78]. L. [59]. labor [874]. Lagrangian [652]. Lambda [241]. Lambert [1042]. Land [604]. Language [734, 843, 250, 1157, 778, 1155, 827, 951, 747]. Languages [242, 111, 577, 959]. Languages- [242]. Large [240]. Later [84, 762]. lawhood [912]. Lawlessness [484]. Laws [537, 589, 483, 492, 526, 362, 660, 484, 486, 182, 1094, 1050, 485, 767]. lazy [9]. Learning [1117, 380, 1108]. Leibniz [952, 238]. Leibnizian [318, 374]. Length [239, 1172]. lens [961]. Leonard [255]. Leslie [299]. Lessons [141]. Let [469]. level [677]. levels [994]. Levi [285]. Levy [927]. Lewis [436]. Lewisian [774, 769]. Liar [387]. Liberalism [1177]. Life [43, 936]. Light [246]. Like [776, 765, 146, 1041]. Likelihood [36, 770]. Likelihood-based [36]. Likelihoodism [787]. likelihoods [1154]. Limit [432]. Limited [400]. Limits [411, 103, 545, 1157]. Line [299, 551]. Linguistic [1008, 707, 283].

Link [173]. Linsky [255]. List [407]. Literacy [178]. Literature [426]. Liu [724]. local [1090]. Localist [119]. Localization [559]. Locating [1047]. Location [645, 796]. Löf [545]. Logic [614, 1100, 24, 822, 270, 113, 641, 548, 549, 611, 227, 213, 31, 185, 117, 261, 750, 523, 177, 110, 815, 788, 474, 443, 114, 612, 115, 456, 610, 645, 431, 615, 112, 550, 613, 35, 646, 79, 1159, 1005, 1107, 65, 982, 983, 984, 986, 1108, 1099, 1103, 1020, 757, 1015, 120, 940, 1058, 1059, 1163, 1102, 90, 1061, 1004, 1054, 758]. Logical [839, 298, 372, 589, 438, 1056, 519, 627, 1057, 72, 228, 74, 506, 655, 423, 925, 517, 97, 554, 646, 1006, 1156, 1076, 746, 1157, 1048, 1049, 827]. logicism [1126]. Logico [156]. Logics [376, 62, 67, 378, 823, 118, 647, 988, 1045, 824]. Long [173]. Long-Term [173]. Look [583, 423, 765]. Looking [86, 203]. Loop [626]. Loops [420, 319, 850]. Lost [58]. LOT [1137]. Lottery [166]. lover [993]. Lovers [357]. LTP [173]. luck [836, 837, 832, 830, 834]. lucky [833]. Lumps [581].

MacBeath [385]. MacFarlane [1149]. Mach [370]. machines [1094, 1117]. Maddy [508]. Made [605]. Magnetic [422]. Maintenance [506]. Majority [1002]. make [797, 1018, 965]. Makers [687]. Making [525, 687, 1063, 676, 911, 898, 1139]. Man [8]. Mandelstam [59]. Manifest [274]. Manifestation [471]. Manifolds [72, 779]. Many [68, 1116, 990]. many-body [1116]. Many-Valued [68]. Margin [1013]. Mark [427]. markers [860]. Martin [545]. Martin-Löf [545]. Masses [768]. Massive [855]. Master [354]. Match [273]. materiality [1115]. Mathematical [262, 708, 569, 307, 263, 83, 269, 261, 246, 271, 508, 267, 266, 842, 84, 864, 943, 1126]. mathematician [911]. Mathematics [121, 892, 69, 268, 85, 781, 237, 591, 63, 744, 90, 255]. matter [1115]. matters [931, 856]. Maximin [398]. Maxwell [253]. Mayo [962]. Mcculloch [423]. McDowell [1176]. McKinsey [184]. me [849]. Meaning [690, 734, 332, 587, 381, 27, 902, 841]. Meaningfulness [27]. Means [533, 826, 238, 926]. Means-end [826]. Measurability [486]. Measure [13, 954]. Measurement [190, 704]. Measures [728, 806, 900, 807, 896]. Measuring [784]. mechanical [567]. mechanics [741]. mechanism [673]. Mechanisms [537]. mechanistic [719]. Medieval [324, 506]. Meditations [427]. Mein [156]. Memoriam [218]. Memories [436]. Memory [752, 379, 871, 1171, 793, 1109]. Mental [155, 322, 860, 819, 147, 888]. Mere [431]. Meredith [642]. Mereological [1070]. Merging [206, 696, 1002]. Meta [947]. Meta-epistemology [947]. Metacognition [852]. metaethics [701]. Metalogic [335]. Metamathematics [91]. Metaphysical [353, 105, 932]. Metaphysically [663]. Metaphysics [453, 469, 318, 306, 253, 617, 819, 204, 156, 604, 361, 936, 1095, 952]. metarepresentation [852]. Method [708, 15]. Methodological [169, 957, 1065, 676]. Methodology [24, 822, 774, 914]. methods [939, 1122, 1120]. Might [601, 765]. Milk [43]. Millennium [559]. Mind [752, 19, 6, 2, 3, 631, 105, 557, 144, 423, 560, 254, 674, 854, 703, 852, 678, 883, 675, 758, 208]. Mind/ [208]. Mind/Brain [254]. minds [756]. Mini [300]. Mini-Environments [300]. Minimal [691, 827, 1045]. Minimalism [387, 281, 95, 1180]. Minimalist [141]. Minimally [62]. Miracle [121]. Miroirs [318]. mirror [1063]. misidentification [1171]. Mistakes [323]. Modal [113, 549, 739, 646, 1107, 914, 973]. Modality [619, 41, 325, 253, 978]. Model [752, 733, 796, 124, 373, 442, 443, 47, 244, 174, 26, 960, 577, 1039, 1036, 985, 1116, 900, 1133, 987]. Modeling [537, 559, 1112]. Modelling [151, 167, 1004]. Models [62, 41, 537, 593, 71, 14, 248, 175, 362, 695, 686, 719, 1042, 1116, 764, 1117, 1067, 960]. Moderate [866]. Modern [386, 273]. Modes [776, 296]. Modest [905]. modular [923]. modularization [989]. Molding [731]. molecular [674]. molecule [675]. Moment [647]. Moment/History [647]. Monism [280, 568]. Monotone [241]. Monotonic [517]. Monsters* [83]. Monty [927, 821]. mood [841]. Moore [32]. Moorean [608]. Moral [43, 634, 1034, 44, 701, 1033, 993, 1151, 903]. most [1019]. Motor [426]. Moulines [666]. Multi [378, 380, 994]. Multi-Agent [378, 380]. multi-scale [994]. Multideductive [79]. multiflora [321]. Multiple [715, 730, 1065, 1066, 1064]. multiplicity [1067]. Mumford [1124]. Mundale [1066]. murder [795]. Must [738]. My [659, 716, 758]. myopia [860]. Mystery [283, 944].

N [30, 207]. Naiveté [652]. Name [321]. Names [134]. Naming [493]. Narrators [861]. Nash [828, 57]. Nativism [707]. Natural [196, 537, 320, 352, 592, 493, 912, 948]. Naturalism [734, 353, 700, 957]. Naturalist [197, 956]. naturalistic [698]. Naturalizing [928]. Nature [496, 483, 567, 10, 992, 809, 1150]. navigating [1083]. Necessary [100, 890]. Necessities [662]. Necessity [372, 190, 483, 594, 651]. need [813]. Negation [277, 596]. Negative [782]. Neglected [715]. neighbourhood [1078]. Neil [274]. Neo [542, 1146, 1148]. Neo-Fregean [542, 1146, 1148]. Nervous [423]. Networks [441, 496, 222, 526]. Neumann [379]. Neural [441, 526]. neuroethics [937]. neurological [937]. Neuronal [425, 723]. neurons [1063]. neuropsychology [856]. Neuroscience [1066, 427, 674, 720, 722, 856, 718, 1067]. Neuroscientific [426]. neutrality [460]. Neutralizing [627]. Never [327]. Newcomb [358, 790]. Newtonian [52, 203, 432]. Neyman [962]. Nézondet [71]. No [470, 420, 473, 1068, 1042, 1072]. Nominals [775]. Non [915, 117, 38, 55, 383, 98, 517, 1141, 1090, 988, 980, 1067, 965]. non-classical [988]. Non-Cognitivism [915]. Non-Commutative [38]. Non-Compositionality [98]. non-conceptual [1141]. Non-Cooperation [117]. non-entropic [1090]. non-experimental [965]. Non-Monotonic [517]. non-reductionist [1067]. Non-Sequential [383]. non-transitivity [980]. Non-Unique [55]. Nonconceptual [148]. Nonidealized [160]. Nonindexical [1025]. Nonmonotonic [441, 519, 524]. Nonmonotonicity [520]. Nonphysical [557]. Nonreductive [606]. Nonreferring [56]. Normative [577, 956]. normativity [885]. notation [240]. Note [1, 29, 347]. Notes [596]. Notice [427, 309, 844]. Notion [595]. Notions [818, 298, 87]. Novel [749]. novelty [1122]. Number [165, 546]. Numbers [542, 251, 788, 240].

Object [556]. Objection [629, 1144]. Objective [959]. Objectively [601]. Objectives [715, 120]. Objectivity [365, 165, 138, 781]. Objects [308, 138, 268, 462, 1048, 1049]. objects- [462]. Obligation [612, 853]. Obligationes [506, 655]. obligations [1006]. Observation [49]. observational [965]. Observe [428]. OCC [1076]. Ockham [277, 1060]. off [675]. old [1121, 874]. Omitting [498]. On-line [551]. One [737, 208, 1179, 910, 828, 19]. one-person [828]. One-Sided [737]. Ones [271]. Ontic [517]. Ontological [305, 310, 991, 1023, 660, 766]. Ontology [630, 73, 40, 183, 665, 267, 98, 666, 1036, 1074, 1085]. Open [297, 1151]. Open-Endedness [297]. Operation [30, 1094]. Operationalism [59]. operations [855]. operator [950]. operator-based [950]. Operators [644]. Optimization [42]. or-Implicature [521]. Order [245, 213, 279, 552, 577, 1038, 1113, 874]. Order-Independent [552]. Ordinal [246]. Ordinals [597]. Origins [691, 702]. Oscillating [568]. Other [321]. ought [969]. Our [103, 491, 436, 601, 212]. outcomes [804]. Outs [644]. over-generalization [953]. Overlap [131]. overriding [524]. Overview [518].

Papineau [944]. Paraconsistency [70, 76, 112, 743]. Paraconsistent [65, 67]. Paradox [360, 355, 32, 502, 275, 553, 166, 1105, 881, 460, 461]. Paradoxes [281, 184, 875, 696, 1158]. parallel [75]. Parfaits [318]. Paribus [182, 1050]. paribus-laws [1050]. Paris [332]. Part [322, 551]. Part-Whole [322]. Partial [69, 342, 81]. Partially [418]. Particles [77]. Particulars [567]. Partition [1040]. Pascal [726]. Past [510, 920, 871]. pathology [868]. pathway [675]. pathways [674]. Patterns [169, 976]. Pauli [408]. payoff [941]. Pearson [962]. Pedestrian [66]. Perception [528, 143, 440, 508, 858]. Perceptual [410, 776, 970, 58, 146, 980, 1176]. perfect [795]. Permission [503, 553]. Permutability [383]. permutation [543]. Person [639, 828, 945]. personal [932, 935]. Persons [930]. Perspective [293, 569, 632, 639, 3, 517, 380, 1012, 120]. perspectives [664]. perspectivism [681]. Persuasiveness [264]. Pessimistic [162]. Pettit [407]. Phenomenal [559, 602, 840, 874]. Phenomenalism [308]. Phenomenological [216, 885]. phenomenology [722, 924, 888]. Philipp [714]. Philosophant [181]. philosopher [956, 1166]. Philosophical [293, 370, 585, 1153, 59, 795, 1122, 937]. Philosophicus [156]. Philosophy [752, 24, 61, 18, 25, 33, 39, 45, 53, 176, 192, 201, 209, 217, 231, 256, 822, 420, 255, 559, 409, 178, 59, 85, 58, 84, 800, 943, 1146, 986, 1121, 765, 762, 678, 748]. Phono [627]. Phono-Logical [627]. Physical [307, 269, 271, 486, 54, 369, 84, 230, 79, 923, 669, 1043, 668]. Physical-Mathematical [269]. Physicalism [50, 322, 662, 682, 683].

Physicalists [713]. Physics [386, 69, 306, 420, 59, 606, 1094, 1116, 664, 1136, 747]. Pickpocket [602]. Pictures [585, 1153]. Pinch [636]. Pit [613]. Pitts [423, 925]. Place [496, 35, 106, 677]. Plane [80]. Plantinga [300]. Platonic [993]. Platonism [308]. Plausible [895]. Plea [147, 1132, 1049]. Plural [29]. Pluralism [660, 1065, 675]. Pluralist [490]. pluralistic [992]. plurality [1096]. Poincaré [229]. Point [627, 937, 241]. Points [68]. Political [290]. Pollock [195]. pooling [966]. positive [1061]. Possession [615]. Possibilistic [518]. possibilities [1072]. Possibility [773, 732, 1087]. Possible [642, 463, 911]. Post [181, 89, 670, 178]. Post-genomics [670]. Post-Hégélien [181]. Post-Literacy [178]. Post-Tarskian [89]. Postcript [88]. Potency [619]. potential [1118]. Potentiation [173]. Poverty [313]. Power [269]. Practical [693, 340, 803, 935]. Practice [337, 943]. Pragmatic [553, 1009]. Pragmatics [133, 135, 131, 753, 1132, 1008, 1010, 1011, 1012, 1007]. Pragmatism [285, 333]. Praise [9]. précis [820]. precondition [852]. predicate [959]. Predicates [77, 946, 745, 1114, 953]. Predication [1097]. Predictability [434]. Predictions [749, 430, 797, 1018]. Preemption [224]. Preface [233, 640, 287, 109]. Preferability [439]. Preference [1000, 1161]. Preference-based [1000, 1161]. preferences [1101, 825, 1006]. premises [850]. Presentation [776]. Presuppositions [607]. Primary [1173]. Principle [473, 571, 158, 770, 958, 723, 1093]. Principles [74, 377, 381, 782]. Prior [641, 642, 647]. priors [786]. Prisoner [312]. prisoners [460]. Pritchard [831]. Private [522, 277, 1155, 1003]. privileged [840]. Probabilistic [1001, 523, 474, 11, 728, 223, 524, 1102]. Probabilities [400, 473, 398, 601, 927, 821, 964]. Probability [711, 339, 572, 288, 689, 225]. Problem [494, 61, 387, 412, 418, 382, 358, 556, 715, 438, 628, 498, 116, 631, 728, 332, 565, 1014, 462, 791, 873, 950, 998, 821, 464, 1151, 672, 944, 1156, 953]. Problem* [570]. Problems [61, 18, 25, 33, 39, 45, 53, 176, 192, 201, 209, 217, 231, 256, 544, 504, 634, 626, 14, 383, 982, 790, 703, 1098, 1106, 747]. Process [452, 451, 1037, 1036]. process-based [1036]. Processes [198, 741]. Program [216, 168, 167, 334, 545, 337, 1148]. Programme [542, 333]. Programming [242]. Programs [376, 551, 265]. Projectible [1114]. Projective [495]. prolegomena [760, 550]. Pronouns [29, 712]. Proof [708, 569, 165, 590, 246, 586, 594, 597, 588, 591, 615]. Proof-theoretic [590, 586, 594, 588, 591]. Proofs [263, 329, 265, 266, 587, 264, 337, 1172]. Propensities [221]. Propensity [224, 222, 792]. Proper [537]. Properties [478, 418, 559, 369]. Property [422, 763]. Proportionality [522]. proposal [456]. Proposals [425]. Proposition [654, 742]. Propositional [815, 72, 284, 610, 996, 869, 846, 1099]. Propositions [532, 704, 284, 712, 657]. Prospect [290]. prospective [1085]. Prospects [688, 703]. protein [764]. Proto [633]. Proto-Rigidity* [633]. Protocols [444, 118, 615, 1067]. Provability [114]. Proving [245, 615]. Provirus [735]. provisional [75]. Proxytypes [707]. pseudo [903]. pseudo-question [903]. psychic [1117]. Psychological [48, 1155, 104, 843]. Psychologism [516, 335]. Psychology [699, 992, 1034, 664, 702]. Psychophysical [492, 723]. Public [1003, 823, 824]. Pust [873]. putative [903]. Putnam [105, 21, 282]. Puzzle [440].

Q [78]. QFT [305]. Qualia [50, 557]. Qualitative [525, 526]. qualities [1153, 1173]. quandary [1170]. Quantification [40]. Quantified [1005]. Quantifier [607, 250, 80]. Quantifier-Free [80]. Quantifiers [869]. Quantitative [605, 521]. Quantum [293, 304, 306, 73, 745, 567, 652, 65, 1091, 75, 741, 126]. Quantum-mechanical [567]. Quasi [64, 743]. Quasi-Set [64]. Quasi-Truth [64, 743]. Quasianalysis [437]. Question [204, 695, 1151, 903]. Questions [632, 456, 457, 27, 1020, 1148, 960]. Questions- [632]. Quine [446, 529, 869, 627, 895]. Quipus [243].

Radical [363, 1172, 184]. Radiobiological [140]. Ramsey [240, 433]. Randomization [264]. Randomness [350]. Rank [470]. Ranking [195]. Rational [6, 133, 57, 8, 476, 197, 903]. Rationalist [738]. Rationality [15, 161, 212, 475, 16, 510, 42, 432, 1105, 794, 1080, 1174]. razor [1060]. re} [771, 532]. Reactive [244]. Read [327]. Reading [141, 651]. Real [542, 1178]. Realism [774, 304, 124, 306, 508, 105, 660, 314, 439, 311, 741, 864, 914, 879, 866, 899, 838, 1172, 901, 970, 916]. Realist [21, 738]. Realistic [469]. Realists [653, 1044]. Reality [270, 643]. Realization [104, 1066, 1064, 1065]. Really [749, 275, 317, 790, 931, 1115]. Reason [429, 510, 82, 10]. Reasonable [7, 69]. Reasoned [396]. Reasoning [522, 262, 1053, 272, 518, 785, 269, 519, 400, 523, 474, 1101, 551, 273, 516, 520, 517, 195, 803, 577, 1109, 1019, 746, 1152, 524, 1060]. Reasons [903, 910]. Rebutting [969]. Received [257]. Recognition [151, 655]. Reconsidered [571, 1071, 1023]. Reconsidering [312]. Reconstruction [437]. Reconstructions [133]. redeployment [855]. Reducing [674]. Reduction [173, 669, 374, 174, 887, 369, 672, 104, 198, 679, 664, 685, 680, 766, 670, 665, 668]. Reductionism [556, 676, 674]. reductionist [677, 1067]. Reductive [994]. redux [1138]. Reference [355, 137, 99, 1158, 875, 1014, 1092, 791, 820]. Reflection [447, 46]. Reflections [540, 172, 4, 458, 416, 571]. reflexive [1084]. reflexivity [565]. Reformed [579]. Refutations [507]. Refuted [254]. Regained [166]. Regress [570, 1021]. Reification [644]. Rejoinder [312]. Relating [118]. Relation [730, 918]. relational [787]. Relations [826, 374, 289, 205]. Relationships [422]. Relative [726, 1024, 1030, 1027, 1003]. relativism [1033, 1034, 1031, 1032, 1029]. relativistic [863]. Relativity [305, 428, 385, 432, 22, 580]. relativized [666]. Relativizing [1123]. Relevant [481, 1058]. Reliabilism [410]. Reliabilist [324]. Reliability [807, 686, 316]. Remarks [272, 73, 985, 936, 38]. remembering [851]. Remodelling [306]. repair [1046]. replies [1077]. Reply [681, 921, 310, 673, 669, 685, 680, 667, 677, 666, 675, 1066, 873, 926].

Representation [693, 639, 179, 156, 153, 157, 154, 698, 894, 1097, 874]. Representational [267]. Representationalism [876, 888]. Representations [150]. Representing [155]. Repsychologising [654]. reputations [802]. Requirement [739, 584]. requirements [801, 903]. Rereading [336]. Research [216]. Resolution [245, 358]. Resolve [494, 680]. Resolved [166, 1069]. Resolving [679]. Resource [173]. Resources [400]. Response [7, 212, 1175, 1124]. Responsibility [906, 754, 853]. restatement [792]. Restorative [709]. restored [1149]. Results [71, 406, 1116, 1003]. Rethinking [530]. Retraction [116]. Retrospective [458]. Return [230, 673, 672]. Revelation [254]. Review [255, 60, 427]. Revising [1160, 912]. Revision [549, 630, 5, 550, 1000, 1001, 1162, 761]. Revisited [589, 508, 334, 629, 816, 869, 543]. Revolutions [816]. Reymonds [544]. rhetoric [1082]. Riddle [635]. riddles [544]. Riemannian [238]. Rigging [709]. Right [429]. Rights [34]. rigidity [1168]. Rigidity* [633]. rigidly [953]. Rigor [264]. Rigorous [1116, 238]. Risk [288, 865, 825, 106]. risk-averse [825]. Road [210, 983]. robustness [954]. Roger [655]. Role [510, 760, 986, 756, 951, 678]. Room [932, 1137]. Roots [370, 324, 672]. Rosa [321]. Rotation [269]. Routine [286]. routines [846]. Rovane [405]. Rueger [669]. Rule [759, 1000, 1179]. rule-based [1000]. rule-following [1179]. Rules [429, 383, 552, 398, 106, 1178]. Run [583]. Ruphy [667]. Russell [255, 180, 730, 952, 194, 654]. Russellian [330].

S [291]. S-Shaped [291]. Safe [709]. Safety [832, 831]. Said [132]. Salmon [218]. Salt [636]. Same [874, 442, 1121]. same-order [874]. Sameness [38]. Samuel [713]. Sartwellian [533]. SAT [443]. SAT-Based [443]. Saunders [310]. Saxony [324]. Say [129]. Scalar [891, 521, 1012]. scale [759, 994]. scandal [1055]. Scenarios [276]. Sceptical [1072, 1169]. Scepticism [196, 184, 46, 1175, 1176]. Schaffner [673]. Schema [799]. scheme [875]. Schemes [505]. Schick [393]. Schlick [228]. School [330, 59]. Science [24, 822, 148, 617, 501, 409, 290, 167, 456, 23, 35, 681, 1081, 740, 743, 986, 765, 1155, 1080, 678, 748, 1087, 989]. science- [681]. Sciences [234, 559, 912, 669, 668]. Scientific [818, 482, 412, 15, 124, 472, 168, 172, 340, 14, 321, 507, 738, 457, 186, 879, 838, 901, 894, 960]. Scope [234]. Search [425, 276]. second [1113]. second-order [1113]. secondary [1153, 1173]. Secret [659, 716]. Secrets [615]. Security [444]. see [146]. Seeing [411, 519]. Seeking [457]. Seem [557]. Seidenfeld [399]. selecting [960]. Selection [361]. Self [148, 796, 1167, 510, 853, 603, 875, 909, 861, 850, 1091, 871, 1171, 852, 847, 820]. self-adjointness [1091]. Self-ascription [853]. self-ascriptions [850]. self-awareness [861]. Self-Consciousness [148, 1171]. self-control [909, 853]. self-directed [852]. Self-Knowledge [510, 1167]. Self-location [796]. self-reference [875, 820]. Self-Verification [603]. Semantic [124, 314, 100, 91, 592, 92, 885]. Semantics [101, 642, 190, 1027, 378, 772, 444, 826, 242, 586, 131, 753, 51, 297, 769, 588, 591, 595, 61, 1008, 1011, 1012, 464, 1007, 989, 133]. Semantics/ [133]. semantics/pragmatics [1008, 1011, 1012, 1007]. Sen [397]. sensations [780]. Sense [558, 268, 898, 950, 1139, 1063]. Sensitivity [727]. sensorimotor [146]. Sensory [38]. Sentence [742, 66]. Sentences [47, 49]. Sentential [98]. Separate [949]. Separate- [949]. Separation [74]. Sequential [383]. Series [735]. Serious [469]. Set [64, 107, 542, 508, 364, 297, 247, 813]. Set-Concept [364]. set-structure [813]. Sets [525, 406, 398, 251, 827]. Setting [169, 819]. Severity [574]. Sex [130]. Shafer [36]. Shannon [1073, 467]. Shaped [2, 291]. Shared [813]. Sharp [123, 356]. Shelter [576]. Shift [91]. shifted [1079]. shortcuts [1056]. shotgun [1026]. show [836]. Showing [372]. Sided [737]. Siécle [181]. Sight [420]. Signaling [580]. Signalling [111]. Signature [425]. Significance [137, 314]. signifying [953]. Similarity [629]. Simon [310]. Simple [549, 579, 66, 31, 185]. Simplified [71]. simulacra [1113]. simulated [1114]. simulation [1121, 1122, 961]. simulations [1119, 1115, 1117]. simultaneous [1004]. Single [927, 821]. Single-case [927, 821]. Six [377]. skeptical [461]. Skepticism [382, 661, 921, 1069, 814, 699]. Skill [858]. skolem [459]. Sleeping [796, 873, 1079, 979, 870]. Slow [235]. Small [208, 997]. smile [673]. Smith [706, 164]. Social [2, 226, 767, 843, 951, 865]. societies [1118]. Society [237]. Soft [374]. Software [226]. Solution [553, 875, 1105]. solves [998]. Solving [710]. somatic [860]. Some [62, 985, 533, 747, 71, 1006, 936, 272]. Soon [46]. Sorites [275]. Sort [1078]. sortal [745]. Sortals [631]. Sosa [844]. Source [178]. Soviet [59]. Space [386, 755, 866, 1057, 973]. space-time [866]. space-times [973]. Spaces [270]. spatial [858]. Spatiotemporal [435]. speaker [1030]. Special [783, 385, 1035, 912]. Species [204, 974]. Specification [7]. specified [582]. spectra [876]. Spectroscopic [408]. Speech [136, 608, 1006]. Spinoza [294]. Spite [495]. Spontaneously [17]. Square [385]. Stability [315, 485]. Staccato [583]. Staggering [450]. Standard [733, 12]. Statements [400, 582]. States [50, 847]. Static [531]. Statistical [748, 956, 961, 963]. Statistics [253, 567]. Statues [581]. Status [652]. Stein [403]. stew [1121]. Stimulation [422]. Stimulus [313]. Stock [726]. strand [1179]. Strange [581]. Strata [429]. Strategic [498, 1163]. strategies [828]. strategy [1045]. Strength [519, 703, 524]. Strengthening [801]. strict [682, 683]. Strikes [230]. Strong [503, 662]. Structural [262, 304, 306, 660, 439, 205, 156, 916, 311, 866, 899]. Structuralism [943, 172, 167, 174, 85, 342, 462, 1125, 1156]. Structuralist [169, 173]. Structure [306, 371, 97, 476, 923, 972, 843, 998, 813, 1133, 994, 960]. Structured [173, 712]. Structured-Propositions [712]. Structures [307, 537, 69, 175, 85, 81, 1127, 939, 1103]. Studies [518, 35]. Study [496, 353, 138, 288, 759, 758]. style [893]. Styles [271]. sub [723]. sub-threshold [723]. Subhuman [277]. subject [934, 932]. subjectivity [934]. Subjects [56, 933]. Submodels [236]. substance [952, 1180].

Substantive [353, 584]. Success [659, 716, 686]. sufficient [966]. Sums [238]. Superassertibility [799, 566]. Supererogatory [125]. superfluous [998]. Superluminal [580]. Superluminality [125]. Supertask [203]. Supertasks [52]. Supervenience [123, 11, 105, 1129]. support [968]. Surprise [750]. survey [967]. Surveyability [569]. Swyneshed [655]. Syllogistic [522]. Symmetries [17]. Symmetry [16]. Synchrnoic [937]. Synchronic [873, 54]. Synonymy [185]. Syntactic [139]. Syntactical [160]. Syntax [839, 733]. Synthese [783]. Synthesis [310, 767]. synthetic [967]. System [752, 52, 295, 93]. Systematic [630]. Systematicity [1138]. Systems [378, 526, 244, 546, 380, 1109, 1114, 970].

Tacit [778]. Tait [255]. Take [220]. Taking [726]. tale [1120]. tamed [1145]. Tarski [446, 298, 449, 241, 450, 93, 651, 282, 91]. Tarskian [89, 94, 93]. task [759]. Taxonomic [321]. Techniques [244]. Technology [1080, 1085, 1087, 1084]. Teleofunctions [630]. Teleological [630, 780]. Temin [735]. Temporal [614, 191, 443, 51, 604, 527, 1103, 1018]. ten [762]. Tennant [274]. Tense [643, 51, 880]. Tenseless [211]. tension [928]. Term [173, 239]. Terms [379, 191, 352, 356]. Terrestrial [269]. Testability [500]. Testimonial [363]. testimonies [896]. Testimony [212]. Testing [187, 962]. Tests [433]. that/knowledge [1128]. Their [234, 485, 834, 960]. them [615]. Theorem [245, 160, 107, 141, 984, 817, 241, 987]. theorems [983]. Theoretic [124, 190, 704, 107, 517, 380, 79, 792, 954, 590, 586, 594, 588, 591]. Theoretical [691, 167, 61, 814]. Theories [386, 124, 170, 4, 495, 14, 74, 597, 79, 797, 923, 985, 779, 893, 309]. Theory [293, 412, 452, 548, 688, 64, 305, 304, 373, 7, 715, 69, 165, 590, 451, 730, 138, 454, 495, 340, 246, 416, 290, 362, 467, 491, 94, 507, 174, 156, 297, 343, 423, 28, 545, 597, 432, 100, 1037, 770, 26, 431, 12, 652, 35, 230, 1076, 792, 1092, 962, 247, 760, 985, 968, 1073, 698, 1162, 1108, 858, 542, 146, 951, 240, 852, 1046, 987, 874, 758, 761]. Theory-Commitment [340]. There [216, 66, 1068, 913, 1136]. Thesis [729, 881]. things [913]. think [1050]. Thinking [837, 271, 849, 936]. third [720]. Thompson [611]. Thought [370, 188, 771, 229, 140, 527, 603, 945]. Thoughts [20, 4, 1046]. Threat [408]. Threats [154]. Three [1077, 1125, 974, 460]. threshold [723]. Tian [309]. Tib [950]. Tibbles [950]. Time [614, 386, 880, 443, 385, 440, 223, 647, 1109, 1090, 1094, 866, 860, 722, 767, 385]. Times [164, 973]. TMS [422]. Today [152]. Together [620]. Token [565]. Token-reflexivity [565]. Tokens [211, 164, 134]. tolerance [1010]. too [791]. Tool [273]. Tools [134]. Topological [74]. Topology [236]. tortoises [1117]. touchstone [896]. track [675]. track-Reply [675]. Tractarian [30, 777, 1048]. Tractatus [372, 578, 156]. tradeoffs [1133]. Trajectories [224]. Transactional [626]. transcendence [446]. transcendental [886]. Transcranial [422]. transfer [1116, 1084]. transferability [842]. Transformations [611, 288, 1106]. Transformative [383, 552]. transition [759, 741]. transitivity [980]. Translations [67]. Transmission [363, 184].

Transparency [254, 1013]. Transtheoretical [175]. travel [860]. Treatments [160]. trends [740]. Trivalent [772]. Trouble [351]. True [344, 446]. Truetemp [410]. Truly [536]. Trumping [1028]. Trust [436, 865]. Truth [1024, 64, 372, 102, 211, 329, 447, 578, 448, 89, 210, 491, 94, 93, 343, 98, 100, 99, 687, 128, 97, 566, 686, 92, 490, 743, 922, 880, 1027, 1030, 838, 1173, 812, 1158, 1123]. truth-conducive [812]. Truth-Makers [687]. Truth-Making [687]. Truthlikeness [15]. truthmaker [782]. Truthmakers [978, 890]. truths [782]. Trying [494]. Tuning [504]. Turing [729, 881]. Turn [177, 1083, 1080, 1084]. Turning [1134]. Twentieth [177, 178]. Twin [43, 634, 44]. Two [818, 494, 272, 836, 87, 584, 385, 203, 406, 464, 21, 352, 153, 1098, 296, 299, 119, 154, 995, 913, 983, 804, 1148, 1022, 398, 1120]. two-argument [1022]. Two-Dimensional [385, 352, 464]. Two-Envelope [494]. Type [712, 250, 545, 949]. Types [241, 995].

ugly [1147]. Unattended [468]. Unbound [712]. Unbounded [443]. uncertain [1154]. Uncertainty [48, 900, 941]. uncomputable [1113]. Uncovering [1127]. Undecidability [815, 28, 1099]. Underdetermination [170, 408, 895]. Understanding [472, 509, 363, 402, 778]. Unfinkable [882]. unfounded [1124]. Ungrounded [618, 1124]. Unification [171, 200, 79]. Union [24, 822]. Unique [55]. Units [361]. Unity [777, 234, 935, 1067]. Univers [318]. Universals [205]. Universe [620]. University [714]. unnatural [953]. Unnecessary [220]. Unobservable [411]. Unqualified [358]. Unsafe [535]. Unsolvability [438]. Update [444, 1104]. updating [873]. Use [128, 1075, 894, 945, 398]. Using [140, 1022, 1045, 965]. Utilities [726]. Utility [291]. utterance [1123]. utterance-truth [1123].

V [571]. v2.0 [858]. vacuity [894]. Vague [772, 47, 63, 356, 900]. Vaguely [772]. Vagueness [123, 31, 48, 354, 731, 1070, 465]. valid [969]. Validity [588, 26, 868]. Value [726, 582, 993, 1174]. Valued [68]. Values [592]. variable [524]. variable-strength [524]. Variables [77]. Varieties [186, 419, 947]. Vector [316]. Vehicles [279]. verdict [1177]. Verification [603]. Verificationists [653]. Verifying [1109]. Verisimilitude [418, 736, 817]. vernacular [446]. version [1068]. Versions [299]. Versus [271, 653, 398, 235, 686, 744, 949, 455]. Very [157]. Via [587]. View [124, 324, 627, 594, 927, 932, 937, 721, 671]. Views [143]. Vinci [1088]. Vindicated [915]. Violated [182]. Virtual [318, 325]. Virtue [38, 908, 909, 836]. Virtues [622, 975, 942]. Vision [254, 1148]. Volume [193, 259, 232, 301, 345, 366, 413, 388, 487, 538, 561, 598, 623, 648, 512]. voluntarism [908]. Voluntary [426]. Vs [262, 579, 83, 681, 644].

W [255]. W. [255]. Wager [726]. Walter [925]. Wang [724]. Warrant [344, 922, 566, 1116, 901]. Wave [183]. Wavelength [230]. Way [711, 63, 720]. Ways [272, 203, 808]. Weak [810]. weddings [1026]. Wedge [299]. Went [23]. Wesley [218]. Weyl [229]. Whence [702]. Where [23, 786, 675, 442, 757]. Which [66, 808, 101]. While [615]. White [529]. whodunit [795]. Whole [322]. Whom [392]. Whose [101]. Why-Questions [456]. will [848]. William [981]. Williams [196]. Williamson [354]. Wimsatt [677]. Wissenschaftslogik [986]. within [883]. Without [185, 428, 284, 470, 922, 687, 1180, 686, 883]. Wittgenstein [255, 839, 372, 915, 730, 1015, 331, 1172, 156, 84, 608]. Word [151]. Words [122, 4]. Work [734, 714]. working [911]. World [19, 163, 601, 1027, 883]. Worlds [642, 560, 463]. Worries [831, 1072]. worry [840]. Wright [1176, 799]. Wrong [215, 23].

XX [181].

Yablo [502]. years [762]. Yes [1019]. Yu [309].

Z [604]. Z-Land [604]. Zeno [863]. Zero [1010]. ZFC [66]. zombie [1134]. Zombies [602, 663].

References

Bishop:2000:EN

[1] Michael Bishop, Richard Samuels, and Stephen Stich. Editors’ note. Synthese, 122(1–2):1, February 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

      content/pdf/10.1023/A%3A1005255608590.pdf.

Cummins:2000:HSE

[2] Denise Dellarosa Cummins. How the social environment shaped the evolution of mind. Synthese, 122(1–2):3–28, February 2000. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

      link.springer.com/content/pdf/10.1023/A%3A1005263825428.pdf.

Griffiths:2000:HMG

[3] Paul E. Griffiths and Karola Stotz. How the mind grows: A developmental perspective on the biology of cognition. Synthese, 122(1–2): 29–51, February 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

      0964 (electronic). URL http://link.springer.com/content/pdf/10.
      1023/A%3A1005215909498.pdf.

Glymour:2000:AEB

[4] Clark Glymour. Android epistemology for babies: Reflections on words, thoughts and theories. Synthese, 122(1–2):53–68, February 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005267926337.pdf.

Pollock:2000:BRE

[5] John L. Pollock and Anthony S. Gillies. Belief revision and epistemology. Synthese, 122(1–2):69–92, February 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005220010407.pdf.

Chater:2000:RAM

[6] Nick Chater and Mike Oaksford. The rational analysis of mind and behavior. Synthese, 122(1–2):93–131, February 2000. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005272027245.pdf.

Dawes:2000:TIR

[7] Robyn M. Dawes. A theory of irrationality as a ‘reasonable’ response to an incomplete specification. Synthese, 122(1–2):133–163, February 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005224211316.pdf.

Sosa:2000:MRA

[8] Ernest Sosa and David Galloway. Man the rational animal? Synthese, 122(1–2):165–178, February 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005276228154.pdf. See notice [844].

Bishop:2000:PEI

[9] Michael A. Bishop. In praise of epistemic irresponsibility: How lazy and ignorant can you be? Synthese, 122(1–2):179–208, February 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005228312224.pdf.

Taylor:2000:WNC

[10] Kenneth A. Taylor. What in nature is the compulsion of reason? Synthese, 122(1–2):209–244, February 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005280329063.pdf.

Kim:2000:SCP

[11] Sungsu Kim. Supervenience and causation: A probabilistic approach. Synthese, 122(3):245–259, March 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005282128866.pdf.

Vallentyne:2000:SDT

[12] Peter Vallentyne. Standard decision theory corrected. Synthese, 122(3): 261–290, March 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005247705806.pdf.

Schoch:2000:FME

[13] Daniel Schoch. A fuzzy measure for explanatory coherence. Synthese, 122(3):291–311, March 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005230112028.pdf.

Klee:2000:PFM

[14] Robert Klee. Problems with formal models of epistemic entrenchment as applied to scientific theories. Synthese, 122(3):313–320, March 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005290330684.pdf.

Bonilla:2000:TRS

[15] Jesus P. Zamora Bonilla. Truthlikeness, rationality and scientific method. Synthese, 122(3):321–335, March 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005269826141.pdf.

Kruse:2000:ISR

[16] Michael Kruse. Invariance, symmetry and rationality. Synthese, 122(3): 337–357, March 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005048811720.pdf.

Kosso:2000:ESB

[17] Peter Kosso. The epistemology of spontaneously broken symmetries. Synthese, 122(3):359–376, March 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005213321366.pdf.

Anonymous:2000:PPa

[18] Anonymous. Problems of philosophy. Synthese, 122(3):377, March 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017231305454.pdf.

Bernecker:2000:KWK

[19] Sven Bernecker. Knowing the world by knowing One’s mind. Synthese, 123(1):1–34, April 2000. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1005239420827.pdf.

Glock:2000:ATC

[20] Hans-Johann Glock. Animals, thoughts and concepts. Synthese, 123(1): 35–64, April 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005295521736.pdf.

Moran:2000:HPI

[21] Dermot Moran. Hilary Putnam and Immanuel Kant: Two ‘internal realists’? Synthese, 123(1):65–104, April 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005273927958.pdf.

Ross:2000:RC

[22] Peter W. Ross. The relativity of color. Synthese, 123(1):105–129, April 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005265725232.pdf.

Stainton:2000:JFC

[23] Robert J. Stainton and Christopher Viger. Jerry A. Fodor, concepts: Where cognitive science went wrong. Synthese, 123(1):131–151, April 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005204225362.pdf.

Anonymous:2000:IUH

[24] Anonymous. International union of history and philosophy of science division of logic, methodology and philosophy of science. Synthese, 123(1): 153–162, April 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1017277701972.pdf.

Anonymous:2000:PPb

[25] Anonymous. Problems of philosophy. Synthese, 123(1):163, April 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017299722292.pdf.

Stephanou:2000:MTV

[26] Yannis Stephanou. Model theory and validity. Synthese, 123(2):165–193, May 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005217809302.pdf.

Verheggen:2000:MMQ

[27] Claudine Verheggen. The meaningfulness of meaning questions. Synthese, 123(2):195–216, May 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005243504897.pdf.

Raatikainen:2000:AIT

[28] Panu Raatikainen. Algorithmic information theory and undecidability. Synthese, 123(2):217–225, May 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005298819345.pdf.

Cartwright:2000:NPP

[29] H. M. Cartwright. A note on plural pronouns. Synthese, 123(2):227–246, May 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005205406576.pdf.

Cheung:2000:TOE

[30] Leo K. C. Cheung. The tractarian operation n and expressive completeness. Synthese, 123(2):247–261, May 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005286219444.pdf.

Everett:2000:SLC

[31] Theodore J. Everett. A simple logic for comparisons and vagueness. Synthese, 123(2):263–278, May 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005283218535.pdf.

Jacquette:2000:IIM

[32] Dale Jacquette. Identity, intensionality, and Moore’s paradox. Synthese, 123(2):279–292, May 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005234102606.pdf.

Anonymous:2000:PPc

[33] Anonymous. Problems of philosophy. Synthese, 123(2):293, May 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017251806362.pdf.

Fleurbaey:2000:RGF

[34] Marc Fleurbaey and Martin Van Hees. On rights in game forms. Synthese, 123(3):295–326, June 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005232419918.pdf.

Walton:2000:PDT

[35] Douglas Walton. The place of dialogue theory in logic, computer science and communication studies. Synthese, 123(3):327–346, June 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005237527730.pdf.

Aickin:2000:CDS

[36] Mikel Aickin. Connecting Dempster–Shafer belief functions with likelihood-based inference. Synthese, 123(3):347–364, June 2000. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/content/pdf/10.1023/A%3A1005287422506.
pdf.

Graham:2000:CI

[37] Peter J. Graham. Conveying information. Synthese, 123(3):365–392, June 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005162716568.pdf.

Lange:2000:JCD

[38] Marc Lange. Is Jeffrey conditionalization defective by virtue of being non-commutative? Remarks on the sameness of sensory experiences.

Synthese, 123(3):393–403, June 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005207627094.pdf.

Anonymous:2000:PPd

[39] Anonymous. Problems of philosophy. Synthese, 123(3):405, June 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017203923201.pdf.

Lavine:2000:QO

[40] Shaughan Lavine. Quantification and ontology. Synthese, 124(1–2):1–43, July 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005257423415.pdf.

Blanchette:2000:MM

[41] Patricia A. Blanchette. Models and modality. Synthese, 124(1–2):45–72, July 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005261524323.pdf.

Mongin:2000:DOI

[42] Philippe Mongin. Does optimization imply rationality? Synthese, 124(1– 2):73–111, July 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005150001309.pdf.

Copp:2000:MHG

[43] David Copp. Milk, honey, and the good life on moral twin Earth. Synthese, 124(1–2):113–137, July 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005278727197.pdf.

Horgan:2000:CMT

[44] Terry Horgan and Mark Timmons. Copping out on moral twin Earth. Synthese, 124(1–2):139–152, July 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005234212937.pdf.

Anonymous:2000:PPe

[45] Anonymous. Problems of philosophy. Synthese, 124(1–2):153, July 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017256008180.pdf.

Vickers:2000:BIS

[46] John M. Vickers. I believe it, but soon i’ll not believe it any more: Scepticism, empiricism, and reflection. Synthese, 124(2):155–174, August 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005213608394.pdf.

Kyburg:2000:WVS

[47] Alice Kyburg. When vague sentences inform: A model of assertability. Synthese, 124(2):175–191, August 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005200520511.pdf.

Hark:2000:UVP

[48] Michel Ter Hark. Uncertainty, vagueness and psychological indeterminacy. Synthese, 124(2):193–220, August 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005251425277.pdf.

Modee:2000:OSJ

[49] Johan Modée. Observation sentences and joint attention. Synthese, 124 (2):221–238, August 2000. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1005241707530.pdf.

Campbell:2000:PQI

[50] Neil Campbell. Physicalism, qualia inversion, and affective states. Synthese, 124(2):239–255, August 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005237026544.pdf.

Mozersky:2000:TTS

[51] Joshua M. Mozersky. Tense and temporal semantics. Synthese, 124(2): 257–279, August 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005268831005.pdf.

Alper:2000:WNS

[52] J. S. Alper, M. Bridger, J. Earman, and J. D. Norton. What is a Newtonian system? The failure of energy conservation and determinism in supertasks. Synthese, 124(2):281–293, August 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005257429769.pdf. See

comments [202].

Anonymous:2000:PPf

[53] Anonymous. Problems of philosophy. Synthese, 124(2):295, August 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017212225018.pdf.

Rueger:2000:PED

[54] Alexander Rueger. Physical emergence, diachronic and synchronic. Synthese, 124(3):297–322, September 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005249907425.pdf.

Meirav:2000:NUC

[55] A. Meirav. Non-unique composition. Synthese, 124(3):323–342, September 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005212811794.pdf.

Carson:2000:AEI

[56] Scott Carson. Aristotle on existential import and nonreferring subjects. Synthese, 124(3):343–360, September 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005270116956.pdf.

Risse:2000:WRA

[57] Mathias Risse. What is rational about Nash equilibria? Synthese, 124(3): 361–384, September 2000. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1005259701040.pdf.

Richeimer:2000:HPL

[58] Joel Richeimer. How philosophy lost perceptual expertise. Synthese, 124(3):385–406, September 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005247308438.pdf.

Pechenkin:2000:OPS

[59] A. A. Pechenkin. Operationalism as the philosophy of Soviet physics: The philosophical backgrounds of L. I. Mandelstam and his school. Synthese, 124(3):407–432, September 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005248403672.pdf.

Hintikka:2000:RA

[60] Jaakko Hintikka and Paul Bohan-Broderick. Review article. Synthese, 124(3):433–445, September 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005288120487.pdf.

Anonymous:2000:PPP

[61] Anonymous. Problems of philosophy; problem #30: An epistemology for game-theoretical semantics. Synthese, 124(3):447, September 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017264209088.pdf.

Batens:2000:MAM

[62] Diderik Batens. Minimally abnormal models in some adaptive logics. Synthese, 125(1–2):5–18, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005208424356.pdf.

VanBendegem:2000:AMV

[63] Jean Paul Van Bendegem. Alternative mathematics: The vague way. Synthese, 125(1–2):19–31, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005248104792.pdf.

Bueno:2000:QTQ

[64] Otávio Bueno. Quasi-truth in quasi-set theory. Synthese, 125(1–2):33– 53, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005238423022.pdf.

Chiara:2000:PIQ

[65] Maria Luisa Dalla Chiara and Roberto Giuntini. Paraconsistent ideas in quantum logic. Synthese, 125(1–2):55–68, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005296018904.pdf.

Doria:2000:TSP

[66] Francisco Antonio Doria. Is there a simple, pedestrian arithmetic sentence which is independent of ZFC? Synthese, 125(1–2):69–76, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005244002066.pdf.

DOttaviano:2000:PLT

[67] Itala M. Loffredo D’Ottaviano and Hércules de Araújo Feitosa. Paraconsistent logics and translations. Synthese, 125(1–2):77–95, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005298624839.pdf.

Drossos:2000:MVP

[68] Costas Drossos and Daniele Mundici. Many-valued points and equality. Synthese, 125(1–2):97–101, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005291917995.pdf.

French:2000:REM

[69] Steven French. The reasonable effectiveness of mathematics: Partial structures and the application of group theory to physics. Synthese, 125(1–2):103–120, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005246608001.pdf.

Grant:2000:APD

[70] John Grant and V. S. Subrahmanian. Applications of paraconsistency in data and knowledge bases. Synthese, 125(1–2):121–132, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005231601157.pdf.

Guillaume:2000:SME

[71] Marcel Guillaume. Simplified models establishing some of Nézondet’s results on Erdős-=Woods conjecture. Synthese, 125(1–2):133–146, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005227515270.pdf.

Kouneiher:2000:PML

[72] J. Kouneiher and A. P. M. Balan. Propositional manifolds and logical cohomology. Synthese, 125(1–2):147–154, January 2000. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005275414361.pdf.

Krause:2000:RQO

[73] Décio Krause. Remarks on quantum ontology. Synthese, 125(1–2):155– 167, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005235125284.pdf.

Mortensen:2000:TSP

[74] Chris Mortensen. Topological separation principles and logical theories. Synthese, 125(1–2):169–178, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005263111635.pdf.

Paty:2000:QCD

[75] Michel Paty. The quantum and classical domains as provisional parallel coexistents. Synthese, 125(1–2):179–200, January 2000. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005288605275.pdf.

Rahman:2000:DAP

[76] Sahid Rahman and Walter A. Carnielli. The dialogical approach to paraconsistency. Synthese, 125(1–2):201–232, January 2000. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005294523930.pdf.

SantAnna:2000:EPH

[77] Adonai S. Sant’Anna. Elementary particles, hidden variables, and hidden predicates. Synthese, 125(1–2):233–245, January 2000. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005272827990.pdf.

Sette:2000:CQ

[78] Antonio Mario Sette and José Carlos Cifuentes. Compactification of l(q). Synthese, 125(1–2):247–252, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005286322113.pdf.

deSouza:2000:MLT

[79] Edelcio G. de Souza. Multideductive logic and the theoretic–formal unification of physical theories. Synthese, 125(1–2):253–262, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005254826656.pdf.

Suppes:2000:QFA

[80] Patrick Suppes. Quantifier-free axioms for constructive affine plane geometry. Synthese, 125(1–2):263–281, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005242507092.pdf.

Tsuji:2000:PSJ

[81] Marcelo Tsuji. Partial structures and Jeffrey–Keynes algebras. Synthese, 125(1–2):283–299, January 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005250725748.pdf.

Parsons:2000:RI

[82] Charles Parsons. Reason and intuition*. Synthese, 125(3):299–315, December 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005204911291.pdf.

Feferman:2000:MIV

[83] Solomon Feferman. Mathematical intuition vs. mathematical monsters*. Synthese, 125(3):317–332, December 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005223128130.pdf.

Steiner:2000:MIP

[84] Mark Steiner. Mathematical intuition and physical intuition in Wittgenstein’s later philosophy. Synthese, 125(3):333–340, December 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005118023258.pdf.

Reck:2000:SSC

[85] Erich H. Reck and Michael P. Price. Structures and structuralism in contemporary philosophy of mathematics. Synthese, 125(3):341–383, December 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005203923553.pdf.

Carbone:2000:LIO

[86] A. Carbone and S. Semmes. Looking from the inside and from the outside. Synthese, 125(3):385–416, December 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005244027724.pdf.

DaSilva:2000:HTN

[87] Jairo José Da Silva. Husserl’s two notions of completeness. Synthese, 125(3):417–438, December 2000. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005265017902.pdf.

Hintikka:2001:IP

[88] J. Hintikka. Introduction and postcript. Synthese, 126(1–2):1–15, January 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005287229088.pdf.

Hintikka:2001:PTT

[89] Jaakko Hintikka. Post-Tarskian truth. Synthese, 126(1–2):17–36, January 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005235624149.pdf.

Sandu:2001:ILF

[90] Gabriel Sandu and Tapani Hyttinen. IF logic and the foundations of mathematics. Synthese, 126(1–2):37–47, January 2001. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005231507310.pdf.

Sinaceur:2001:ATS

[91] Hourya Sinaceur. Alfred Tarski: Semantic shift, heuristic shift in metamathematics. Synthese, 126(1–2):49–65, January 2001. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005268531418.pdf.

Wolenski:2001:DSD

[92] Jan Woleński. In defense of the semantic definition of truth. Synthese, 126(1–2):67–90, January 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005243814579.pdf.

Mou:2001:ECT

[93] Bo Mou. The enumerative character of Tarski’s definition of truth and its general character in a Tarskian system. Synthese, 126(1–2):91–122, January 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005254704585.pdf.

Moreno:2001:TTC

[94] Luis Fernández Moreno. Tarskian truth and the correspondence theory. Synthese, 126(1–2):123–148, January 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005223221423.pdf.

Horwich:2001:DM

[95] Paul Horwich. A defense of minimalism. Synthese, 126(1–2):149–165, January 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005279406402.pdf.

Halbach:2001:HID

[96] Volker Halbach. How innocent is deflationism? Synthese, 126(1–2):167– 194, January 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005275222332.pdf.

Sher:2001:TLS

[97] Gila Sher. Truth, logical structure, and compositionality. Synthese, 126(1–2):195–219, January 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005237812821.pdf.

Puntel:2001:TSN

[98] Lorenz B. Puntel. Truth, sentential non-compositionality, and ontology. Synthese, 126(1–2):221–259, January 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005283218810.pdf.

Schantz:2001:TR

[99] Richard Schantz. Truth and reference. Synthese, 126(1–2):261–281, January 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005287609128.pdf.

Rumfitt:2001:STN

[100] Ian Rumfitt. Semantic theory and necessary truth. Synthese, 126(1–2): 283–324, January 2001. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1005283412921.pdf.

Boulter:2001:WCW

[101] Stephen Boulter. Whose challenge? Which semantics? Synthese, 126(1– 2):325–337, January 2001. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1005246919602.pdf.

Culina:2001:CT

[102] Boris Culina. The concept of truth. Synthese, 126(1–2):339–360, January 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005237331006.pdf.

Dietrich:2001:CLO

[103] Eric Dietrich and Anthony Gillies. Consciousness and the limits of our imaginations. Synthese, 126(3):361–381, March 2001. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005270501025.pdf.

Schweizer:2001:RRP

[104] Paul Schweizer. Realization, reduction and psychological autonomy. Synthese, 126(3):383–405, March 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005203928401.pdf.

Marras:2001:PCM

[105] Ausonio Marras. On Putnam’s critique of metaphysical realism: Mind– body identity and supervenience. Synthese, 126(3):407–426, March 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005230727492.pdf.

Weirich:2001:RPD

[106] Paul Weirich. Risk’s place in decision rules. Synthese, 126(3):427– 441, March 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005240226961.pdf.

Guenin:2001:STA

[107] Louis M. Guenin. The set theoretic ambit of Arrow’s theorem. Synthese, 126(3):443–472, March 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005255924462.pdf.

Hacking:2001:ACC

[108] Ian Hacking. Aristotelian categories and cognitive domains. Synthese, 126(3):473–515, March 2001. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1005221431872.pdf.

Rahman:2001:P

[109] Shahid Rahman and Helge Rückert. Preface. Synthese, 127(1–2):1–6, April 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1010314319474.pdf.

Hintikka:2001:ILE

[110] Jaakko Hintikka. Intuitionistic logic as epistemic logic. Synthese, 127(1– 2):7–19, April 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1010357829038.pdf.

Sandu:2001:SLI

[111] Gabriel Sandu. Signalling in languages with imperfect information. Synthese, 127(1–2):21–34, April 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1010362218565.pdf.

VanBendegem:2001:PDL

[112] Jean Paul Van Bendegem. Paraconsistency and dialogue logic critical examination and further explorations. Synthese, 127(1–2):35–55, April 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1010310101727.pdf.

Blackburn:2001:MLD

[113] Patrick Blackburn. Modal logic as dialogical logic. Synthese, 127(1–2): 57–93, April 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1010358017657.pdf.

Nortmann:2001:HED

[114] Ulrich Nortmann. How to extend the dialogical approach to provability logic. Synthese, 127(1–2):95–103, April 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1010349800818.pdf.

Rahman:2001:DCL

[115] Shahid Rahman and Helge Rückert. Dialogical connexive logic. Synthese, 127(1–2):105–139, April 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1010351931769.pdf.

Krabbe:2001:PRC

[116] Erik C. W. Krabbe. The problem of retraction in critical discussion. Synthese, 127(1–2):141–159, April 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1010318403544.pdf.

Gabbay:2001:NCD

[117] Dov Gabbay and John Woods. Non-cooperation in dialogue logic. Synthese, 127(1–2):161–186, April 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1010370420383.pdf.

Prakken:2001:RPD

[118] Henry Prakken. Relating protocols for dynamic dispute with logics for defeasible argumentation. Synthese, 127(1–2):187–219, April 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1010322504453.pdf.

Vreeswijk:2001:EDB

[119] Gerard A. W. Vreeswijk. Eight dialectic benchmarks discussed by two artificial localist disputors. Synthese, 127(1–2):221–253, April 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1010374523108.pdf.

Lorenz:2001:BOD

[120] Kuno Lorenz. basic objectives of dialogue logic in historical perspective. Synthese, 127(1–2):255–263, April 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1010367416884.pdf.

Colyvan:2001:MAM

[121] Mark Colyvan. The miracle of applied mathematics. Synthese, 127(3): 265–278, June 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1010309227321.pdf.

Cappelen:2001:BW

[122] Herman Cappelen and Josh Dever. Believing in words. Synthese, 127(3): 279–301, June 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1010372531110.pdf.

Ebbs:2001:VSB

[123] Gary Ebbs. Vagueness, sharp boundaries, and supervenience conditions. Synthese, 127(3):303–323, June 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1010353810702.pdf.

Chakravartty:2001:SMT

[124] Anjan Chakravartty. The semantic or model–theoretic view of theories and scientific realism. Synthese, 127(3):325–345, June 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1010359521312.pdf.

Monton:2001:SS

[125] Bradley Monton and Brian Kierland. Supererogatory superluminality. Synthese, 127(3):347–357, June 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1010399110654.pdf.

Bostrom:2001:DAA

[126] Nick Bostrom. The Doomsday argument: Adam & Eve, UN^{++}, and Quantum Joe. Synthese, 127(3):359–387, June 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1010350925053.pdf. See

comment [470].

Kvart:2001:CAC

[127] I. Kvart. The counterfactual analysis of cause. Synthese, 127(3):389–427, June 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1010398724481.pdf.

Seigfried:2001:TU

[128] Hans Seigfried. Truth & use. Synthese, 128(1–2):1–13, July 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1010309510255.pdf.

Bach:2001:YDS

[129] Kent Bach. You don’t say? Synthese, 128(1–2):15–44, July 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1010353722852.pdf.

Taylor:2001:SBD

[130] Kenneth A. Taylor. Sex, breakfast, and Descriptus Interruptus. Synthese, 128(1–2):45–61, July 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1010349621943.pdf.

Manor:2001:OPS

[131] Ruth Manor. On the overlap of pragmatics and semantics. Synthese, 128 (1–2):63–73, July 2001. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1010305806922.pdf.

Recanati:2001:WS

[132] François Recanati. What is said. Synthese, 128(1–2):75–91, July 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1010383405105.pdf.

Garcia-Carpintero:2001:GRR

[133] Manuel García-Carpintero. Gricean rational reconstructions and the Semantics/ pragmatics distinction. Synthese, 128(1–2):93–131, July 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1010301706013.pdf.

Pelczar:2001:NTN

[134] M. W. Pelczar. Names as tokens and names as tools. Synthese, 128(1–2): 133–155, July 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1010304825913.pdf.

Hinzen:2001:PIC

[135] Wolfram Hinzen. The pragmatics of inferential content. Synthese, 128(1– 2):157–181, July 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1010362521497.pdf.

Asher:2001:ISA

[136] Nicholas Asher and Alex Lascarides. Indirect speech acts. Synthese, 128 (1–2):183–228, July 2001. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1010340508140.pdf.

Loeffler:2001:DRC

[137] Ronald Loeffler. Demonstrative reference and cognitive significance. Synthese, 128(3):229–244, September 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1011918626632.pdf.

Hauser:2001:OOC

[138] Kai Hauser. Objectivity over objects: A case study in theory formation. Synthese, 128(3):245–285, September 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1011994619998.pdf.

Guarini:2001:DCA

[139] Marcello Guarini. A defence of connectionism against the “syntactic” argument. Synthese, 128(3):287–317, September 2001. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1011905917986.pdf.

Shrader-Frechete:2001:UTE

[140] Kristan Shrader-Frechete. Using a thought experiment to clarify a radiobiological controversy. Synthese, 128(3):319–342, September 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1011961527452.pdf.

Muller:2001:AMR

[141] Thomas Müller and Tomasz Placek. Against a minimalist reading of Bell’s theorem: Lessons from fine. Synthese, 128(3):343–379, September 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1011971808127.pdf.

Myin:2001:EI

[142] Erik Myin. Editorial introduction. Synthese, 129(1):1–2, October 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1012663122859.pdf.

Hurley:2001:PAA

[143] Susan Hurley. Perception and action: Alternative views. Synthese, 129 (1):3–40, October 2001. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1012643006930.pdf.

Noe:2001:EAM

[144] Alva Noë. Experience and the active mind. Synthese, 129(1):41–60, October 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1012695023768.pdf.

Myin:2001:CDA

[145] Erik Myin. Color and the duplication assumption. Synthese, 129(1): 61–77, October 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1012647207838.pdf.

ORegan:2001:WIL

[146] J. Kevin O’Regan and Alva noë. What it is like to see: A sensorimotor theory of perceptual experience. Synthese, 129(1):79–103, October 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1012699224677.pdf.

Proust:2001:PMA

[147] Joëlle Proust. A plea for mental acts. Synthese, 129(1):105–128, October 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1012651308747.pdf.

Bermudez:2001:NSC

[148] José Luis Bermúdez. Nonconceptual self-consciousness and cognitive science. Synthese, 129(1):129–149, October 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1012603425585.pdf.

Vacariu:2001:I

[149] Mihai Vacariu, Edmund T. Rolls, and Gabriel Vacariu. Introduction. Synthese, 129(2):151, November 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1013067728072.pdf.

Rolls:2001:RB

[150] Edmund T. Rolls. Representations in the brain. Synthese, 129(2):153– 171, November 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1013059525140.pdf.

McLeod:2001:CMW

[151] Peter McLeod, David C. Plaut, and Tim Shallice. Connectionist modelling of word recognition. Synthese, 129(2):173–183, November 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1013003423323.pdf.

Plunkett:2001:CT

[152] Kim Plunkett. Connectionism today. Synthese, 129(2):185–194, November 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1013099222414.pdf.

Rantala:2001:KRT

[153] Veikko Rantala. Knowledge representation: Two kinds of emergence. Synthese, 129(2):195–209, November 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1013063626048.pdf.

Wheeler:2001:TTR

[154] Michael Wheeler. Two threats to representation. Synthese, 129(2):211– 231, November 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1013055424231.pdf.

Bogdan:2001:DMA

[155] Radu J. Bogdan. Developing mental abilities by representing intentionality. Synthese, 129(2):233–258, November 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1013051306484.pdf.

Parvu:2001:MGL

[156] Ilie Pârvu. “Mein Grundgedanke Ist … ” the structural theory of representation as the metaphysics of Wittgenstein’s Tractatus Logico– Philosophicus. Synthese, 129(2):259–274, November 2001. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1013047205576.pdf.

Vacariu:2001:TVI

[157] Gabriel Vacariu, Dalia Terhesiu, and Mihai Vacariu. Toward a very idea of representation. Synthese, 129(2):275–295, November 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1013019621505.pdf.

Spencer:2001:BPI

[158] Cara Spencer. Belief and the principle of identity. Synthese, 129(3):297– 318, December 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1013174018923.pdf.

Wray:2001:CBA

[159] K. Brad Wray. Collective belief and acceptance. Synthese, 129(3):319– 333, December 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1013148515033.pdf.

Cross:2001:TCS

[160] Charles B. Cross. A theorem concerning syntactical treatments of nonidealized belief. Synthese, 129(3):335–341, December 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1013155912920.pdf.

Bossert:2001:CCR

[161] Walter Bossert. Choices, consequences, and rationality. Synthese, 129(3): 343–369, December 2001. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1017333227870.pdf.

Lewis:2001:WPI

[162] Peter J. Lewis. Why the pessimistic induction is a fallacy. Synthese, 129(3):371–380, December 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1013139410613.pdf.

Jacobson:2001:CGD

[163] Stephen Jacobson. Contextualism and global doubts about the world. Synthese, 129(3):381–404, December 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1013168627459.pdf.

Mozersky:2001:STT

[164] Joshua M. Mozersky. Smith on times and tokens. Synthese, 129(3):405– 411, December 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1013135211562.pdf.

Ganeri:2001:OPC

[165] Jonardon Ganeri. Objectivity and proof in a classical Indian theory of number. Synthese, 129(3):413–437, December 2001. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1013144411940.pdf.

Weintraub:2001:LPR

[166] R. Weintraub. The lottery: A paradox regained and resolved. Synthese, 129(3):439–449, December 2001. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1013192211031.pdf.

Moulines:2002:ISP

[167] C. Ulises Moulines. Introduction: Structuralism as a program for modelling theoretical science. Synthese, 130(1):1–11, January 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1013892808077.pdf.

Diez:2002:PIS

[168] Jose A. Diez. A program for the individuation of scientific concepts. Synthese, 130(1):13–47, January 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1013887521262.pdf.

Balzer:2002:MPS

[169] Wolfgang Balzer. Methodological patterns in a structuralist setting. Synthese, 130(1):49–68, January 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1013871008985.pdf.

Gahde:2002:HUD

[170] Ulrich Gähde. Holism, underdetermination, and the dynamics of empirical theories. Synthese, 130(1):69–90, January 2002. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1013875125824.pdf.

Bartelborth:2002:EU

[171] Thomas Bartelborth. Explanatory unification. Synthese, 130(1):91– 108, January 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1013827209894.pdf.

Forge:2002:RSS

[172] John Forge. Reflections on structuralism and scientific explanation. Synthese, 130(1):109–121, January 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1013879326732.pdf.

Bickle:2002:CST

[173] John Bickle. Concepts structured through reduction: a structuralist resource illuminates the consolidation–long-term potentiation (LTP) link. Synthese, 130(1):123–133, January 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1013831410802.pdf.

Niebergall:2002:SMT

[174] Karl-Georg Niebergall. Structuralism, model theory and reduction. Synthese, 130(1):135–162, January 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1013883427641.pdf.

Lauth:2002:TSD

[175] Bernhard Lauth. Transtheoretical structures and deterministic models. Synthese, 130(1):163–172, January 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1013835511711.pdf.

Anonymous:2002:PPa

[176] Anonymous. Problems of philosophy. Synthese, 130(1):173, January 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017408419153.pdf.

Gochet:2002:DTT

[177] Paul Gochet. The dynamic turn in Twentieth century logic. Synthese, 130(2):175–184, February 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1014435213120.pdf.

Nyiri:2002:PLS

[178] J. C. Nyiri. Post-literacy as a source of Twentieth–century philosophy. Synthese, 130(2):185–199, February 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1014438529050.pdf.

Motzkin:2002:R

[179] Gabriel Motzkin. Representation. Synthese, 130(2):201–212, February 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1014487229958.pdf.

Chateaubriand:2002:DFR

[180] Oswaldo Chateaubriand. Descriptions: Frege and Russell combined. Synthese, 130(2):213–226, February 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1014439314029.pdf.

Bourgeois:2002:XSP

[181] Bernard Bourgeois. Le XX^{e} Siécle philosophant: Post-Hégélien? Synthese, 130(2):227–233, February 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1014491305603.pdf.

Smith:2002:VLC

[182] Sheldon Smith. Violated laws, ceteris paribus clauses, and capacities. Synthese, 130(2):235–264, February 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1014422011637.pdf.

Monton:2002:WFO

[183] Bradley Monton. Wave function ontology. Synthese, 130(2):265–277, February 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1014493527177.pdf.

Pritchard:2002:MPR

[184] Duncan Pritchard. McKinsey paradoxes, radical scepticism, and the transmission of knowledge across known entailments. Synthese, 130(2): 279–302, February 2002. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1014421800473.pdf.

Everett:2002:ASS

[185] Theodore J. Everett. Analyticity without synonymy in simple comparative logic. Synthese, 130(2):303–315, February 2002. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1014402110621.pdf.

Thalos:2002:EGE

[186] Mariam Thalos. Explanation is a genus: An essay on the varieties of scientific explanation. Synthese, 130(3):317–354, March 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1014841517312.pdf.

Douven:2002:TIB

[187] Igor Douven. Testing inference to the best explanation. Synthese, 130(3): 355–377, March 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1014859910339.pdf.

Borsboom:2002:FTE

[188] Denny Borsboom, Gideon J. Mellenbergh, and Jaap van Heerden. Functional thought experiments. Synthese, 130(3):379–387, March 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1014840616403.pdf.

Gundersen:2002:DCA

[189] Lars Gundersen. In defence of the conditional account of dispositions. Synthese, 130(3):389–411, March 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1014845625688.pdf.

Dresner:2002:MTS

[190] Eli Dresner. Measurement theoretic semantics and the semantics of necessity. Synthese, 130(3):413–440, March 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1014815703989.pdf.

Corazza:2002:TIT

[191] Eros Corazza. Temporal indexicals and temporal terms. Synthese, 130(3): 441–460, March 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1014891411735.pdf.

Anonymous:2002:PPb

[192] Anonymous. Problems of philosophy. Synthese, 130(3):461, March 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017405103223.pdf.

Anonymous:2002:CVa

[193] Anonymous. Contents of volume 130. Synthese, 130(3):463–464, March 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017446930583.pdf.

Pincock:2002:RIC

[194] Christopher Pincock. Russell’s influence on Carnap’s Aufbau. Synthese, 131(1):1–37, April 2002. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1015066427566.pdf.

Spohn:2002:BCP

[195] Wolfgang Spohn. A brief comparison of Pollock’s defeasible reasoning and ranking functions. Synthese, 131(1):39–56, April 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1015004212541.pdf.

Buchanan:2002:NDW

[196] Reid Buchanan. Natural doubts: Williams’s diagnosis of scepticism. Synthese, 131(1):57–80, April 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1015082729457.pdf.

Wielenberg:2002:HAR

[197] Erik J. Wielenberg. How to be an alethically rational naturalist. Synthese, 131(1):81–98, April 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1015069614963.pdf.

Thalos:2002:RCP

[198] Mariam Thalos. The reduction of causal processes. Synthese, 131(1): 99–128, April 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1015015024056.pdf.

Persson:2002:CEF

[199] Johannes Persson. Cause, effect, and fake causation. Synthese, 131(1): 129–143, April 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1015055126092.pdf.

Weber:2002:UE

[200] Erik Weber and Maarten Van Dyck. Unification and explanation. Synthese, 131(1):145–154, April 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1015005529380.pdf.

Anonymous:2002:PPc

[201] Anonymous. Problems of philosophy. Synthese, 131(1):155, April 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017457104132.pdf.

Laraudogoitia:2002:DAB

[202] Jon Pérez Laraudogoitia. On the dynamics of Alper and Bridger. Synthese, 131(2):157–171, May 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1015705510482.pdf. See [52].

Laraudogoitia:2002:TWL

[203] Jon Pérez Laraudogoitia, Mark Bridger, and Joseph S. Alper. Two ways of looking at a Newtonian supertask. Synthese, 131(2):173–189, May 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015775628393.pdf.

Okasha:2002:DMS

[204] Samir Okasha. Darwinian metaphysics: Species and the question of essentialism. Synthese, 131(2):191–213, May 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015731831011.pdf.

Pages:2002:SUF

[205] Joan Pagés. Structural universals and formal relations. Synthese, 131(2): 215–221, May 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1015765414172.pdf.

Burger:2002:MIC

[206] Isabella Cornelia Burger and Johannes Heidema. Merging inference and conjecture by information. Synthese, 131(2):223–258, May 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1015737827562.pdf.

MacIntyre:2002:CD

[207] I. D. A. MacIntyre. —n— cheers for democracy. Synthese, 131(2): 259–274, May 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1015749211633.pdf.

Webster:2002:CMB

[208] W. R. Webster. A case of Mind/ brain identity: One small bridge for the explanatory gap. Synthese, 131(2):275–287, May 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1015726204005.pdf.

Anonymous:2002:PPd

[209] Anonymous. Problems of philosophy. Synthese, 131(2):289, May 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015701305041.pdf.

Kuipers:2002:BRT

[210] Theo A. F. Kuipers. Beauty, a road to the truth. Synthese, 131(3): 291–328, June 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1016188509393.pdf.

Dyke:2002:TDT

[211] Heather Dyke. Tokens, dates and tenseless truth conditions. Synthese, 131(3):329–351, June 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1016119503891.pdf.

Faulkner:2002:ROR

[212] P. Faulkner. On the rationality of our response to testimony. Synthese, 131(3):353–370, June 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1016116728471.pdf.

Eklund:2002:HLF

[213] Matti Eklund and Daniel Kolak. Is Hintikka’s logic first-order? Synthese, 131(3):371–388, June 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1016184410627.pdf.

Hintikka:2002:CEK

[214] Jaakko Hintikka. Comment on Eklund and Kolak. Synthese, 131(3): 389–393, June 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1016117912210.pdf.

Leng:2002:WWI

[215] Mary Leng. What’s wrong with indispensability? Synthese, 131(3): 395–417, June 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1016141509719.pdf.

Crowell:2002:TPR

[216] Steven Crowell. Is there a phenomenological research program? Synthese, 131(3):419–444, June 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1016117513804.pdf.

Anonymous:2002:PPe

[217] Anonymous. Problems of philosophy. Synthese, 131(3):445, June 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1016153421879.pdf.

Fetzer:2002:MWC

[218] James H. Fetzer. In memoriam: Wesley C. Salmon (1925–2001). Synthese, 132(1–2):1–3, July 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1019636429609.pdf.

Fetzer:2002:I

[219] James H. Fetzer. Introduction. Synthese, 132(1–2):5–8, July 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1019654314588.pdf.

Kyburg:2002:DTU

[220] Henry E. Kyburg, Jr. Don’t take unnecessary chances! Synthese, 132 (1–2):9–26, July 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1019610615497.pdf.

Fetzer:2002:PFI

[221] James H. Fetzer. Propensities and frequencies: Inference to the best explanation. Synthese, 132(1–2):27–61, July 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1019614716405.pdf.

Gillies:2002:CPB

[222] Donald Gillies. Causality, propensity, and Bayesian networks. Synthese, 132(1–2):63–88, July 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1019618817314.pdf.

Oksanen:2002:PCB

[223] Mika Oksanen. Probabilistic causation in branching time. Synthese, 132 (1–2):89–117, July 2002. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1019622918223.pdf.

Eells:2002:PTP

[224] Ellery Eells. Propensity trajectories, preemption, and the identity of events. Synthese, 132(1–2):119–141, July 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1019627019131.pdf.

Pollock:2002:CP

[225] John L. Pollock. Causal probability. Synthese, 132(1–2):143–185, July 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1019631104110.pdf.

Parikh:2002:SS

[226] Rohit Parikh. Social software. Synthese, 132(3):187–211, September 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1020391420768.pdf.

Dubucs:2002:FL

[227] Jacques Dubucs. Feasibility in logic. Synthese, 132(3):213–237, September 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1020332703930.pdf.

Livingston:2002:HSL

[228] Paul Livingston. Husserl and Schlick on the logical form of experience. Synthese, 132(3):239–272, September 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1020378731241.pdf.

Feist:2002:WAH

[229] Richard Feist. Weyl’s appropriation of Husserl’s and Poincaré”s thought. Synthese, 132(3):273–301, September 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1020370823738.pdf.

Webster:2002:WTC

[230] W. R. Webster. Wavelength theory of colour strikes back: The return of the physical. Synthese, 132(3):303–334, September 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1020345513372.pdf.

Anonymous:2002:PPf

[231] Anonymous. Problems of philosophy. Synthese, 132(3):335, September 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1020359832301.pdf.

Anonymous:2002:VC

[232] Anonymous. Volume contents. Synthese, 132(3):337–338, September 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1020376830100.pdf.

Anonymous:2002:P

[233] Anonymous. Preface. Synthese, 133(1–2):1–4, October 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1020866718668.pdf.

Lowe:2002:FST

[234] Benedikt Löwe. The formal sciences: Their scope, their foundations, and their unity. Synthese, 133(1–2):5–11, October 2002. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1020887832028.pdf.

Weiermann:2002:SVF

[235] Andreas Weiermann. Slow versus fast growing. Synthese, 133(1–2):13– 29, October 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1020899506400.pdf.

Geschke:2002:AES

[236] Stefan Geschke. Applications of elementary submodels in general topology. Synthese, 133(1–2):31–41, October 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1020819407308.pdf.

Stolz:2002:HAM

[237] Michael Stolz. The history of applied mathematics and the history of society. Synthese, 133(1–2):43–57, October 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1020823608217.pdf.

Knobloch:2002:LRF

[238] Eberhard Knobloch. Leibniz’s rigorous foundation of infinitesimal geometry by means of Riemannian sums. Synthese, 133(1–2):59–73, October 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1020859101830.pdf.

Mathias:2002:TL

[239] A. R. D. Mathias. A term of length 4 523 659 424 929. Synthese, 133(1–2): 75–86, October 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1020827725055.pdf.

Promel:2002:LNK

[240] Hans Jürgen Prömel. Large numbers, Knuth’s arrow notation, and Ramsey theory. Synthese, 133(1–2):87–105, October 2002. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1020879709125.pdf.

Matthes:2002:TFP

[241] Ralph Matthes. Tarski’s Fixed-Point Theorem and Lambda calculi with monotone inductive types. Synthese, 133(1–2):107–129, October 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1020831825964.pdf.

Jurjens:2002:GSP

[242] Jan Jürjens. Games in the semantics of programming languagesan elementary introduction. Synthese, 133(1–2):131–158, October 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1020883810034.pdf.

Christensen:2002:IQ

[243] Antje Christensen. The Incan quipus. Synthese, 133(1–2):159–172, October 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1020835926872.pdf.

Merz:2002:MCT

[244] Stephan Merz. Model checking techniques for the analysis of reactive systems. Synthese, 133(1–2):173–201, October 2002. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1020887910943.pdf.

Benzmuller:2002:CAR

[245] Christoph Benzmüller. Comparing approaches to resolution based higher-order theorem proving. Synthese, 133(1–2):203–235, October 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1020840027781.pdf.

Kahle:2002:MPT

[246] Reinhard Kahle. Mathematical proof theory in the light of ordinal analysis. Synthese, 133(1–2):237–255, October 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1020892011851.pdf.

Burr:2002:CAF

[247] Wolfgang Burr. Concepts and aims of functional interpretations: towards a functional interpretation of constructive set theory. Synthese, 133(1– 2):257–274, October 2002. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1020844128690.pdf.

Koepke:2002:CIM

[248] Peter Koepke. The category of inner models. Synthese, 133(1–2):275– 303, October 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1020818827349.pdf.

Moore:2002:CAA

[249] Matthew E. Moore. A Cantorian argument against infinitesimals. Synthese, 133(3):305–330, December 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1021204522829.pdf.

Pambuccian:2002:AHG

[250] Victor Pambuccian. Axiomatizations of hyperbolic geometry: A comparison based on language and quantifier type complexity. Synthese, 133(3): 331–341, December 2002. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1021294808742.pdf. See correction [511].

Steinhart:2002:WNS

[251] Eric Steinhart. Why numbers are sets. Synthese, 133(3):343–361, December 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1021266519848.pdf.

Tieszen:2002:GIC

[252] Richard Tieszen. Gödel and the intuition of concepts. Synthese, 133(3): 363–391, December 2002. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1021247624209.pdf.

Gordon:2002:MBS

[253] Bruce L. Gordon. Maxwell–Boltzmann statistics and the metaphysics of modality. Synthese, 133(3):393–417, December 2002. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1021360805193.pdf.

Webster:2002:RTC

[254] W. R. Webster. Revelation and transparency in colour vision refuted: A case of mind/brain identity and another bridge over the explanatory gap. Synthese, 133(3):419–439, December 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1021294209237.pdf.

Hill:2002:BRW

[255] Claire Hill. Book review: W. Demopoulos (ed.), Frege’s Philosophy of Mathematics, and W. W. Tait (ed.), Early Analytic Philosophy, Frege, Russell, Wittgenstein, Essays in Honor of Leonard Linsky. Synthese, 133(3):441–452, December 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1021209217761.pdf.

Anonymous:2002:PPg

[256] Anonymous. Problems of philosophy. Synthese, 133(3):453, December 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1021248431188.pdf.

Anonymous:2002:BR

[257] Anonymous. Books received 2002. Synthese, 133(3):455, December 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1021244326161.pdf.

Anonymous:2002:IA

[258] Anonymous. Instructions for authors. Synthese, 133(3):457–461, December 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1021269201875.pdf.

Anonymous:2002:CVb

[259] Anonymous. Contents of volume 133. Synthese, 133(3):463–464, December 2002. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1021273717804.pdf.

Anonymous:2003:I

[260] Anonymous. Introduction. Synthese, 134(1–2):1–2, January 2003. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/content/pdf/10.1023/A%3A1022155911458.
pdf.

Gardies:2003:DMC

[261] Jean Louis Gardies. Do mathematical constructions escape logic? Synthese, 134(1–2):3–24, January 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1022148016910.pdf.

Aliseda:2003:MRV

[262] Atocha Aliseda. Mathematical reasoning vs. abductive reasoning: A structural approach. Synthese, 134(1–2):25–44, January 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1022127429205.pdf.

Fallis:2003:IGM

[263] Don Fallis. Intentional gaps in mathematical proofs. Synthese, 134(1–2): 45–69, January 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1022131513275.pdf.

Womach:2003:RPR

[264] Catherine Womach and Matrin Farach. Randomization, persuasiveness and rigor in proofs. Synthese, 134(1–2):71–84, January 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1022183530114.pdf.

Longo:2003:PP

[265] Giuseppe Longo. Proofs and programs. Synthese, 134(1–2):85–117, January 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1022135614184.pdf.

Panza:2003:MP

[266] Marco Panza. Mathematical proofs. Synthese, 134(1–2):119–158, January 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1022187631022.pdf.

Muntersbjorn:2003:RIM

[267] Madeline M. Muntersbjorn. Representational innovation and mathematical ontology. Synthese, 134(1–2):159–180, January 2003. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1022139715092.pdf.

Otte:2003:DMO

[268] M. Otte. Does mathematics have objects? in what sense? Synthese, 134(1–2):181–216, January 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1022191731931.pdf.

Finocchiaro:2003:PMR

[269] Maurice A. Finocchiaro. Physical-mathematical reasoning: Galileo on the extruding power of terrestrial rotation. Synthese, 134(1–2):217– 244, January 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1022143816001.pdf.

Arboleda:2003:FLC

[270] Luis Carlos Arboleda and Luis Cornelio Recalde. Fréchet and the logic of the constitution of abstract spaces from concrete reality. Synthese, 134(1–2):245–272, January 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1022195800071.pdf.

Lacki:2003:SPT

[271] Jan Lacki. Styles of physical thinking versus mathematical ones. Synthese, 134(1–2):273–288, January 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1022175328296.pdf.

Alvarez:2003:TWR

[272] Carlos Alvarez. Two ways of reasoning and two ways of arguing in geometry. Some remarks concerning the application of figures in Euclidean geometry. Synthese, 134(1–2):289–323, January 2003. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1022100100980.pdf.

Lutz:2003:MIT

[273] Robert Lutz and Luis Gonzaga Luis Gonzaga. Modern infinitesimals as a tool to match intuitive and formal reasoning in analysis. Synthese, 134(1–2):325–351, January 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1022152117818.pdf.

Cogburn:2003:MIN

[274] Jon Cogburn. Manifest invalidity: Neil Tennant’s new argument for intuitionism. Synthese, 134(3):353–362, March 2003. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1022921622763.pdf.

Paoli:2003:RFA

[275] Francesco Paoli. A really fuzzy approach to the sorites paradox. Synthese, 134(3):363–387, March 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1022995202767.pdf.

Wisniewski:2003:ESS

[276] Andrzej Wiśniewski. Erotetic search scenarios. Synthese, 134(3):389– 427, March 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1022983325118.pdf.

Martin:2003:ABS

[277] John N. Martin. All brutes are subhuman: Aristotle and Ockham on private negation. Synthese, 134(3):429–461, March 2003. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1022934709521.pdf.

Drai:2003:EI

[278] Dalia Drai. Externalism and identity. Synthese, 134(3):463–475, March 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1022915720037.pdf.

Kriegel:2003:CHO

[279] Uriah Kriegel. Consciousness, higher-order content, and the individuation of vehicles. Synthese, 134(3):477–504, March 2003. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1022913321854.pdf.

Antony:2003:DAM

[280] Michael V. Antony. Davidson’s argument for monism. Synthese, 135 (1):1–12, April 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1022986117864.pdf.

Glanzberg:2003:MP

[281] Michael Glanzberg. Minimalism and paradoxes. Synthese, 135(1):13–36, April 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1022999315312.pdf.

Raatikainen:2003:MPT

[282] Panu Raatikainen. More on Putnam and Tarski. Synthese, 135(1):37–47, April 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1022910022940.pdf.

Fischer:2003:BMA

[283] Eugen Fischer. Bogus mystery about linguistic competence. Synthese, 135(1):49–75, April 2003. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1022945916989.pdf.

Moltmann:2003:PAP

[284] Friederike Moltmann. Propositional attitudes without propositions. Synthese, 135(1):77–118, April 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1022945009188.pdf.

Olsson:2003:AEH

[285] Erik J. Olsson. Avoiding epistemic Hell: Levi on pragmatism and inconsistency. Synthese, 135(1):119–140, April 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1022997026026.pdf.

Levi:2003:CEH

[286] Isaac Levi. Contracting from epistemic Hell is routine. Synthese, 135(1): 141–164, April 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1022915525935.pdf.

Dacey:2003:P

[287] Raymond Dacey. Preface. Synthese, 135(2):165–169, May 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1023435022719.pdf.

Neilson:2003:PTS

[288] William S. Neilson. Probability transformations in the study of behavior toward risk. Synthese, 135(2):171–192, May 2003. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1023408906789.pdf.

Mintz:2003:FEI

[289] Alex Mintz and Steven B. Redd. Framing effects in international relations. Synthese, 135(2):193–213, May 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1023460923628.pdf.

Levy:2003:APT

[290] Jack S. Levy. Applications of prospect theory to political science. Synthese, 135(2):215–241, May 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1023413007698.pdf.

Dacey:2003:SUF

[291] R. Dacey. The s-shaped utility function. Synthese, 135(2):243–272, May 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023465024536.pdf.

Baker:2003:FB

[292] G. P. Baker and P. M. S. Hacker. Functions in begriffsschrift. Synthese, 135(3):273–297, June 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1023509611120.pdf.

Anselmi:2003:NPP

[293] D. Anselmi. A new perspective on the philosophical implications of quantum field theory. Synthese, 135(3):299–328, June 2003. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1023578026046.pdf.

Deveaux:2003:DEC

[294] Sherry Deveaux. The divine essence and the conception of God in Spinoza. Synthese, 135(3):329–338, June 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1023594617953.pdf.

Laraudogoitia:2003:ISG

[295] J. P. Laraudogoitia. An infinite system with gravitation. Synthese, 135 (3):339–346, June 2003. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1023521907716.pdf.

Risse:2003:BGA

[296] M. Risse. Bayesian group agents and two modes of aggregation. Synthese, 135(3):347–377, June 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1023531909371.pdf.

Paseau:2003:OES

[297] A. Paseau. The open-endedness of the set concept and the semantics of set theory. Synthese, 135(3):379–399, June 2003. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1023542621122.pdf.

Bellotti:2003:TLN

[298] Luca Bellotti. Tarski on logical notions. Synthese, 135(3):401–413, June 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023590504284.pdf.

Sober:2003:ECT

[299] E. Sober. An empirical critique of two versions of the Doomsday argument- Gott’s line and Leslie’s wedge. Synthese, 135(3):415–430, June 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023545820214.pdf.

Botham:2003:PFM

[300] T. M. Botham. Plantinga and favorable mini-environments. Synthese, 135(3):431–441, June 2003. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1023551725580.pdf.

Anonymous:2003:CVa

[301] Anonymous. Contents of volume 135. Synthese, 135(3):443–444, June 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023555425712.pdf.

Anonymous:2003:IAa

[302] Anonymous. Instructions for authors. Synthese, 135(3):445–449, June 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023537226621.pdf.

Anonymous:2003:E

[303] Anonymous. Editorial. Synthese, 136(1):1–2, July 2003. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1024170328023.pdf.

Cao:2003:SRI

[304] Tian Yu Cao. Structural realism and the interpretation of quantum field theory. Synthese, 136(1):3–24, July 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1024139713910.pdf.

Cao:2003:AOR

[305] Tian Yu Cao. Appendix: Ontological relativity and fundamentality — is QFT the fundamental theory? Synthese, 136(1):25–30, July 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1024199931657.pdf.

French:2003:RSR

[306] Steven French and James Ladyman. Remodelling structural realism: Quantum physics and the metaphysics of structure. Synthese, 136(1): 31–56, July 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1024156116636.pdf.

Cao:2003:CWD

[307] Tian Yu Cao. Can we dissolve physical entities into mathematical structures? Synthese, 136(1):57–71, July 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1024112417545.pdf.

French:2003:DOB

[308] Steven French and James Ladyman. The dissolution of objects: Between Platonism and phenomenalism. Synthese, 136(1):73–77, July 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1024116502524.pdf.

Saunders:2003:CNT

[309] Simon Saunders. Critical notice: Tian Yu Cao’s “The Conceptual Development of 20th Century Field Theories”. Synthese, 136(1):79–105, July 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1024172603432.pdf. See reply [310].

Cao:2003:WOS

[310] Tian Yu Cao. What is ontological synthesis? — a reply to Simon Saunders. Synthese, 136(1):107–126, July 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1024176704341.pdf. See [309].

Saunders:2003:SRA

[311] Simon Saunders. Structural realism, again. Synthese, 136(1):127–133, July 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1024180822088.pdf.

Jiborn:2003:RFR

[312] Magnus Jiborn and Wlodek Rabinowicz. Reconsidering the Foole’s rejoinder: Backward induction in indefinitely iterated Prisoner’s dilemmas. Synthese, 136(2):135–157, August 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1024731815957.pdf.

Collins:2003:CPS

[313] John Collins. Cowie on the poverty of stimulus. Synthese, 136(2):159– 190, August 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1024738522031.pdf.

Miller:2003:SSR

[314] Alexander Miller. The significance of semantic realism. Synthese, 136(2): 191–217, August 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1024742007683.pdf.

Hansson:2003:WS

[315] S. O. Hansson and G. Helgesson. What is stability? Synthese, 136(2): 219–235, August 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1024733424521.pdf.

Wunderlich:2003:VRN

[316] Mark E. Wunderlich. Vector reliability: A new approach to epistemic justification. Synthese, 136(2):237–262, August 2003. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1024791403375.pdf.

Sansom:2003:WER

[317] R. Sansom. Why evolution is really indeterministic. Synthese, 136(2): 263–279, August 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1024739731042.pdf.

Boos:2003:PMU

[318] William Boos. Parfaits miroirs de l’univers”: A ‘virtual’ interpretation of Leibnizian metaphysics. Synthese, 136(2):281–304, August 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1024745012471.pdf.

Ismael:2003:CCL

[319] J. Ismael. Closed causal loops and the bilking argument. Synthese, 136(3):305–320, September 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1025170026539.pdf.

Guenin:2003:DCN

[320] Louis M. Guenin. Dialogue concerning natural appropriation. Synthese, 136(3):321–336, September 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1025165320946.pdf.

McDonough:2003:RMO

[321] Jeffrey K. McDonough. A Rosa multiflora by any other name: Taxonomic incommensurability and scientific kinds. Synthese, 136(3):337– 358, September 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1025116814353.pdf.

Ehring:2003:PWP

[322] Douglas Ehring. Part-whole physicalism and mental causation. Synthese, 136(3):359–388, September 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1025143104108.pdf.

Roth:2003:M

[323] Paul A. Roth. “Mistakes”. Synthese, 136(3):389–408, September 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025187611264.pdf.

Fitzgerald:2003:MRR

[324] Michael J. Fitzgerald. The medieval roots of reliabilist epistemology: Albert of Saxony’s view of immediate apprehension. Synthese, 136(3):409– 434, September 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1025117406833.pdf.

Boos:2003:VM

[325] William Boos. Virtual modality. Synthese, 136(3):435–492, September 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025101329309.pdf.

Kennedy:2003:I

[326] Juliette Kennedy and Gabriel Sandu. Introduction. Synthese, 137(1–2): 1, November 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1026261615022.pdf.

vanAtten:2003:BNR

[327] Mark van Atten. Brouwer, as never read by Husserl. Synthese, 137(1–2): 3–19, November 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1026270631861.pdf.

Crocco:2003:GCF

[328] Gabriella Crocco. Gödel, Carnap and the Fregean heritage. Synthese, 137(1–2):21–41, November 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1026222715931.pdf.

Fichot:2003:TPF

[329] Jean Fichot. Truth, proofs and functions. Synthese, 137(1–2):43–58, November 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1026274716840.pdf.

Mancosu:2003:RIH

[330] Paolo Mancosu. The Russellian influence on Hilbert and his school. Synthese, 137(1–2):59–101, November 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1026278800910.pdf.

Marion:2003:WB

[331] Mathieu Marion. Wittgenstein and Brouwer. Synthese, 137(1–2):103– 127, November 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1026230917748.pdf.

Moriconi:2003:MHC

[332] Enrico Moriconi. On the meaning of Hilbert’s consistency problem (paris, 1900). Synthese, 137(1–2):129–139, November 2003. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1026282901818.pdf.

Peckhaus:2003:PHP

[333] Volker Peckhaus. The pragmatism of Hilbert’s programme. Synthese, 137(1–2):141–156, November 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1026235118657.pdf.

Raatikainen:2003:HPR

[334] Panu Raatikainen. Hilbert’s program revisited. Synthese, 137(1–2):157– 177, November 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1026287119565.pdf.

Wolenski:2003:PM

[335] Jan Woleński. Psychologism and metalogic. Synthese, 137(1–2):179– 193, November 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1026291203636.pdf.

VonPlato:2003:RG

[336] Jan Von Plato. Rereading Gentzen. Synthese, 137(1–2):195–209, November 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1026243304544.pdf.

Zach:2003:PFE

[337] Richard Zach. The practice of finitism: Epsilon calculus and consistency proofs in Hilbert’s program. Synthese, 137(1–2):211–259, November 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1026247421383.pdf.

Yunez-Naude:2003:WCC

[338] Norma Yunez-Naude. What is Carnap’s conventionalism after all? Synthese, 137(1–2):261–272, November 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1026299405453.pdf.

Hajek:2003:WCP

[339] Alan Hájek. What conditional probability could not be. Synthese, 137(3): 273–323, December 2003. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000004904.91112.16.pdf.

Jones:2003:STC

[340] Ward E. Jones. Is scientific theory-commitment doxastic or practical? Synthese, 137(3):325–344, December 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/B%3ASYNT.0000004901.93310.03.pdf.

Huemer:2003:HHC

[341] Wolfgang Huemer. Husserl and Haugeland on constitution. Synthese, 137 (3):345–368, December 2003. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000004902.07390.18.pdf.

Reck:2003:DSI

[342] Erich H. Reck. Dedekind’s structuralism: An interpretation and partial defense. Synthese, 137(3):369–419, December 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/B%3ASYNT.0000004903.11236.
91.pdf.

Patterson:2003:WCT

[343] D. Patterson. What is a correspondence theory of truth? Synthese, 137(3):421–444, December 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3ASYNT.0000004905.68653.b3.pdf.

Chignell:2003:ATB

[344] Andrew Chignell. Accidentally true belief and warrant. Synthese, 137(3): 445–458, December 2003. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000004906.81978.84.pdf.

Anonymous:2003:CVb

[345] Anonymous. Contents of volume 137. Synthese, 137(3):459–460, December 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000004910.75728.9d.pdf.

Anonymous:2003:IAb

[346] Anonymous. Instructions for authors. Synthese, 137(3):461–465, December 2003. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000005023.11773.9f.pdf.

Symons:2004:EN

[347] John Symons. Editorial note. Synthese, 138(1):v–vii, January 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000012340.55768.ee.pdf.

Gillies:2004:NFE

[348] Anthony S. Gillies. New foundations for epistemic change^{∗}. Synthese, 138(1):1–48, January 2004. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000012202.66263.b2.pdf.

Pelczar:2004:IF

[349] Michael W. Pelczar. The indispensability of Farbung. Synthese, 138(1): 49–78, January 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000012203.59944.c5.pdf.

Campbell:2004:RJI

[350] Scott Campbell and James Franklin. Randomness and the justification of induction. Synthese, 138(1):79–99, January 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/B%3ASYNT.0000012206.01154.
c7.pdf.

Cling:2004:TI

[351] Andrew D. Cling. The trouble with infinitism. Synthese, 138(1):101– 123, January 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000012132.60082.0e.pdf.

Nimtz:2004:TDN

[352] Christian Nimtz. Two-dimensional and natural kind terms. Synthese, 138 (1):125–148, January 2004. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000012205.86593.46.pdf.

Dowell:2004:MSN

[353] J. L. Dowell. From metaphysical to substantive naturalism: A case study. Synthese, 138(2):149–173, January 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3ASYNT.0000013235.88476.e3.pdf.

Ray:2004:WMA

[354] Greg Ray. Williamson’s master argument on vagueness. Synthese, 138 (2):175–206, January 2004. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000013178.23890.4c.pdf.

Hill:2004:RP

[355] Claire Ortiz Hill. Reference and paradox. Synthese, 138(2):207–232, January 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000013241.59493.1d.pdf.

Weintraub:2004:SBV

[356] R. Weintraub. On sharp boundaries for vague terms. Synthese, 138(2): 233–245, January 2004. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000013242.20414.43.pdf.

Tersman:2004:WLG

[357] Folke Tersman. Are we lovers of the good? Synthese, 138(2):247– 260, January 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000013240.14920.f7.pdf.

Burgess:2004:NPU

[358] Simon Burgess. The Newcomb problem: An unqualified resolution. Synthese, 138(2):261–287, January 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3ASYNT.0000013243.57433.e7.pdf.

Corazza:2004:AAN

[359] Eros Corazza. On the alleged ambiguity of ‘now’ and ‘here’. Synthese, 138(2):289–313, January 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3ASYNT.0000013244.65400.a0.pdf.

Heikes:2004:BPW

[360] Deborah K. Heikes. The bias paradox: Why it’s not just for feminists anymore. Synthese, 138(3):315–335, February 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/B%3ASYNT.0000016424.47883.
b9.pdf.

Walsh:2004:BMU

[361] D. M. Walsh. Bookkeeping or metaphysics? The units of selection debate. Synthese, 138(3):337–361, February 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3ASYNT.0000016426.73707.92.pdf.

Liu:2004:LMT

[362] Chuang Liu. Laws and models in a theory of idealization. Synthese, 138 (3):363–385, February 2004. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000016425.36070.37.pdf.

Goldberg:2004:RIU

[363] S. C. Goldberg. Radical interpretation, understanding, and the testimonial transmission of knowledge. Synthese, 138(3):387–416, February 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000016423.39916.ad.pdf.

Muller:2004:IDS

[364] F. A. Muller. The implicit definition of the set-concept. Synthese, 138 (3):417–451, February 2004. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000016439.37687.78.pdf.

Douglas:2004:ICO

[365] Heather Douglas. The irreducible complexity of objectivity. Synthese, 138 (3):453–473, February 2004. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000016451.18182.91.pdf.

Anonymous:2004:CVa

[366] Anonymous. Contents of volume 138. Synthese, 138(3):475–476, February 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000016685.77692.3c.pdf.

Anonymous:2004:AIa

[367] Anonymous. Author index. Synthese, 138(3):477, February 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3ASYNT.
0000016690.01937.ee.pdf.

Anonymous:2004:IAa

[368] Anonymous. Instructions for authors. Synthese, 138(3):479–483, February 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000016686.43901.b6.pdf.

Rueger:2004:RAC

[369] Alexander Rueger. Reduction, autonomy, and causal exclusion among physical properties. Synthese, 139(1):1–21, March 2004. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/B%3ASYNT.0000021305.
91409.d3.pdf.

Banks:2004:PRE

[370] E. C. Banks. The philosophical roots of Ernst Mach’s economy of thought. Synthese, 139(1):23–53, March 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/B%3ASYNT.0000021306.66850.a3.pdf.

Robbins:2004:SS

[371] Philip Robbins. To structure, or not to structure? Synthese, 139(1): 55–80, March 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000021308.22619.83.pdf.

Cheung:2004:SAT

[372] Leo K. C. Cheung. Showing, analysis and the truth–functionality of logical necessity in Wittgenstein’s Tractatus. Synthese, 139(1):81–105, March 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000021309.60256.8e.pdf.

Costantini:2004:EFM

[373] D. Costantini and U. Garibaldi. The Ehrenfest fleas: From model to theory. Synthese, 139(1):107–142, March 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/B%3ASYNT.0000021307.64103.b8.pdf.

Maunu:2004:LSR

[374] Ari Maunu. Leibnizian soft reduction of extrinsic denominations and relations. Synthese, 139(1):143–164, March 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/B%3ASYNT.0000021310.53938.ad.pdf.

vanderHoek:2004:E

[375] Wiebe van der Hoek. Editorial. Synthese, 139(2):5–8, March 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000024924.26800.e3.pdf.

Baltag:2004:LEP

[376] Alexandru Baltag and Lawrence S. Moss. Logics for epistemic programs. Synthese, 139(2):165–224, March 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3ASYNT.0000024912.56773.5e.pdf.

Rott:2004:CSF

[377] Hans Rott. A counterexample to six fundamental principles of belief formation. Synthese, 139(2):225–240, March 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/B%3ASYNT.0000024914.61925.5b.pdf.

Goranko:2004:CSL

[378] Valentin Goranko and Wojciech Jamroga. Comparing semantics of logics for multi-agent systems. Synthese, 139(2):241–280, March 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000024915.66183.d1.pdf.

Bonanno:2004:CNG

[379] Giacomo Bonanno. A characterization of von Neumann games in terms of memory. Synthese, 139(2):281–295, March 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/B%3ASYNT.0000024905.25386.
3d.pdf.

Tuyls:2004:EGT

[380] Karl Tuyls, Ann Nowe, Tom Lenaerts, and Bernard Manderick. An evolutionary game theoretic perspective on learning in multi-agent systems. Synthese, 139(2):297–330, March 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3ASYNT.0000024908.89191.f1.pdf.

VanRooy:2004:ECM

[381] Robert Van Rooy. Evolution of conventional meaning and conversational principles. Synthese, 139(2):331–366, March 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/B%3ASYNT.0000024904.37199.
6c.pdf.

Brogaard:2004:CSG

[382] B. Brogaard. Contextualism, skepticism, and the gettier problem. Synthese, 139(3):367–386, April 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3ASYNT.0000024886.84207.6b.pdf.

Peterson:2004:TDR

[383] Martin Peterson. Transformative decision rules, permutability, and nonsequential framing of decision problems. Synthese, 139(3):387–403, April 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000024885.09894.ba.pdf.

Hitchcock:2004:BB

[384] Christopher Hitchcock. Beauty and the bets. Synthese, 139(3):405– 420, April 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000024889.29125.c0.pdf.

King:2004:TDT

[385] Daniel King. Two-dimensional time: MacBeath’s “Time’s Square” and special relativity. Synthese, 139(3):421–428, April 2004. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/B%3ASYNT.0000024890.
87641.60.pdf.

Boi:2004:TST

[386] Luciano Boi. Theories of space–time in modern physics. Synthese, 139(3): 429–489, April 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000024888.19304.0f.pdf.

Armour-Garb:2004:MGP

[387] Bradley Armour-Garb. Minimalism, the generalization problem and the liar. Synthese, 139(3):491–512, April 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/B%3ASYNT.0000024891.87160.c3.pdf.

Anonymous:2004:VC

[388] Anonymous. Volume contents. Synthese, 139(3):513–514, April 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000024922.94107.f9.pdf.

Anonymous:2004:AIb

[389] Anonymous. Author index. Synthese, 139(3):515, April 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/B%3ASYNT.0000024916.
97501.12.pdf.

Anonymous:2004:IAb

[390] Anonymous. Instructions for authors. Synthese, 139(3):516–520, April 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000024798.84130.7e.pdf.

Bilgrami:2004:I

[391] Akeel Bilgrami. Introduction. Synthese, 140(1–2):1, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/B%3ASYNT.0000029935.
79805.bd.pdf.

Schick:2004:DW

[392] Frederic Schick. A dilemma for whom? Synthese, 140(1–2):3–16, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000029936.77539.27.pdf.

Levi:2004:Sa

[393] Isaac Levi. Schick. Synthese, 140(1–2):17–24, May 2004. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/B%3ASYNT.0000029937.
49242.4a.pdf.

Hintikka:2004:FF

[394] Jaakko Hintikka. A fallacious fallacy? Synthese, 140(1–2):25–35, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000029938.17953.10.pdf.

Levi:2004:JH

[395] Isaac Levi. Jaakko hintikka. Synthese, 140(1–2):37–41, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3ASYNT.
0000029939.13900.04.pdf.

Sen:2004:IRC

[396] Amartya Sen. Incompleteness and reasoned choice. Synthese, 140(1–2): 43–59, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000029940.51537.b3.pdf.

Levi:2004:AS

[397] Isaac Levi. Amartya Sen. Synthese, 140(1–2):61–67, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/B%3ASYNT.0000029941.
44120.87.pdf.

Seidenfeld:2004:CBT

[398] T. Seidenfeld. A contrast between two decision rules for use with (convex) sets of probabilities: Γ-maximin versus E-admissibilty. Synthese, 140(1– 2):69–88, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000029942.11359.8d.pdf.

Levi:2004:Sb

[399] Isaac Levi. Seidenfeld. Synthese, 140(1–2):89–96, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/B%3ASYNT.0000029943.
05040.95.pdf.

Gaifman:2004:RLR

[400] Haim Gaifman. Reasoning with limited resources and assigning probabilities to arithmetical statements. Synthese, 140(1–2):97–119, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000029944.99888.a7.pdf.

Levi:2004:G

[401] Isaac Levi. Gaifman. Synthese, 140(1–2):121–134, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/B%3ASYNT.0000029945.
01194.28.pdf.

Stein:2004:EUE

[402] Howard Stein. The enterprise of understanding and the enterprise of knowledge. Synthese, 140(1–2):135–176, May 2004. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/B%3ASYNT.0000029946.38831.
c9.pdf.

Levi:2004:HS

[403] Isaac Levi. Howard Stein. Synthese, 140(1–2):177–180, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3ASYNT.
0000029947.12801.4d.pdf.

Rovane:2004:WA

[404] Carol Rovane. What is an agent? Synthese, 140(1–2):181–198, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000029948.05384.00.pdf. See comment [405].

Levi:2004:CR

[405] Isaac Levi. Carol Rovane. Synthese, 140(1–2):199–206, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3ASYNT.
0000029949.50644.51.pdf. See [404].

List:2004:ASJ

[406] Christian List and Philip Pettit. Aggregating sets of judgments: Two impossibility results compared. Synthese, 140(1–2):207–235, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000029950.50517.59.pdf. See comment [407].

Levi:2004:LP

[407] Isaac Levi. List and Pettit. Synthese, 140(1–2):237–242, May 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000029951.47770.10.pdf. See [406].

Massimi:2004:WDI

[408] Michela Massimi. What demonstrative induction can do against the threat of underdetermination: Bohr, Heisenberg, and Pauli on spectroscopic anomalies (1921–24). Synthese, 140(3):243–277, June 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000031319.64615.49.pdf.

Leutge:2004:EPS

[409] Christoph Leutge. Economics in philosophy of science: A dismal contribution? Synthese, 140(3):279–305, June 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/B%3ASYNT.0000031318.21858.dd.pdf.

Beebe:2004:RTN

[410] J. R. Beebe. Reliabilism, truetemp and new perceptual faculties. Synthese, 140(3):307–329, June 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3ASYNT.0000031324.35563.1d.pdf.

Alspector-Kelly:2004:SUV

[411] Marc Alspector-Kelly. Seeing the unobservable: Van Fraassen and the limits of experience. Synthese, 140(3):331–353, June 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/B%3ASYNT.0000031323.
19904.45.pdf.

Badino:2004:AIT

[412] Massimiliano Badino. An application of information theory to the problem of the scientific experiment. Synthese, 140(3):355–389, June 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000031325.69601.af.pdf.

Anonymous:2004:CVb

[413] Anonymous. Contents of volume 140. Synthese, 140(3):391–392, June 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000031506.88628.27.pdf.

Anonymous:2004:AIc

[414] Anonymous. Author index. Synthese, 140(3):393, June 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/B%3ASYNT.0000031521.
24323.68.pdf.

Anonymous:2004:IAc

[415] Anonymous. Instructions for authors. Synthese, 140(3):394–398, June 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000031507.22667.03.pdf.

Kuhn:2004:REG

[416] Steven T. Kuhn. Reflections on ethics and game theory. Synthese, 141 (1):1–44, July 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000035846.91195.cb.pdf.

Humberstone:2004:AFI

[417] Lloyd Humberstone. Archetypal forms of inference. Synthese, 141(1): 45–76, July 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000035850.89516.e1.pdf.

Britton:2004:PVC

[418] T. Britton. The problem of verisimilitude and counting partially identical properties. Synthese, 141(1):77–95, July 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/B%3ASYNT.0000035848.61535.f8.pdf.

Vahid:2004:VEC

[419] Hamid Vahid. Varieties of epistemic conservatism. Synthese, 141(1): 97–122, July 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000035849.62840.e8.pdf.

Hanley:2004:NES

[420] Richard Hanley. No end in sight: Causal loops in philosophy, physics and fiction. Synthese, 141(1):123–152, July 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/B%3ASYNT.0000035847.28833.4f.pdf.

Bickle:2004:EI

[421] John Bickle. Editor’s introduction. Synthese, 141(2):153–154, August 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000042950.39098.f6.pdf.

Schutter:2004:ITM

[422] Dennis J. L. G. Schutter, Jack Van Honk, and Jaak Panksepp. Introducing transcranial magnetic stimulation (TMS) and its property of causal inference in investigating brain-function relationships. Synthese, 141(2): 155–173, August 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000042951.25087.16.pdf.

Piccinini:2004:FCT

[423] Gualtiero Piccinini. The first computational theory of mind and brain: A close look at Mcculloch and Pitts’s “logical calculus of ideas immanent in nervous activity”. Synthese, 141(2):175–215, August 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000043018.52445.3e.pdf.

Overgaard:2004:CFC

[424] Morten Overgaard. Confounding factors in contrastive analysis. Synthese, 141(2):217–231, August 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3ASYNT.0000043019.64052.e0.pdf.

Bauer:2004:SNS

[425] Roman Bauer. In search of a neuronal signature of consciousnessfacts, hypotheses and proposals. Synthese, 141(2):233–245, August 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000043020.26100.ba.pdf.

Tibbetts:2004:CVM

[426] Paul Tibbetts. The concept of voluntary motor control in the recent neuroscientific literature. Synthese, 141(2):247–276, August 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000043021.33695.99.pdf.

Ryder:2004:REM

[427] Dan Ryder. Review essay: Meditations on first neuroscience: Critical notice of mark Changizi’s the brain from 25,000 feet. Synthese, 141(2): 277–285, August 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000043022.71332.5d.pdf.

Healey:2004:CCH

[428] Richard Healey. Change without change, and how to observe it in general relativity. Synthese, 141(3):1–35, September 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/B%3ASYNT.0000045127.04908.
1a.pdf.

Cummins:2004:ESR

[429] Robert Cummins, Pierre Poirier, and Martin Roth. Epistemological strata and the rules of right reason. Synthese, 141(3):287–331, September 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000044992.91717.aa.pdf.

Romeyn:2004:HIP

[430] J. W. Romeyn. Hypotheses and inductive predictions. Synthese, 141(3): 333–364, September 2004. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000044993.82886.9e.pdf.

Strevens:2004:BCT

[431] Michael Strevens. Bayesian confirmation theory: Inductive logic, or mere inductive framework? Synthese, 141(3):365–379, September 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000044991.73791.f7.pdf.

Rivadulla:2004:NLR

[432] Ardnés Rivadulla. The Newtonian limit of relativity theory and the rationality of theory change. Synthese, 141(3):417–429, September 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000044994.31650.45.pdf.

Slater:2004:RT

[433] B. H. Slater. Ramsey’s tests. Synthese, 141(3):431–444, September 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000044995.88380.6a.pdf.

Lagerspet:2004:PGK

[434] E. Lagerspet. Predictability and the growth of knowledge. Synthese, 141(3):445–459, September 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3ASYNT.0000044953.98134.4d.pdf.

Bigaj:2004:CSE

[435] Tomasz Bigaj. Counterfactuals and spatiotemporal events. Synthese, 142(1):1–19, October 2004. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000047707.96275.08.pdf.

Olsson:2004:CWT

[436] E. J. Olsson and T. Shogenji. Can we trust our memories? C. I. Lewis’s coherence argument. Synthese, 142(1):21–41, October 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/B%
3ASYNT.0000047708.33913.2b.pdf.

Toader:2004:DRC

[437] Iulian D. Toader. A diagrammatic reconstruction of Carnap’s quasianalysis. Synthese, 142(1):43–59, October 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/B%3ASYNT.0000047711.23748.ca.pdf.

Floridi:2004:LUG

[438] L. Floridi. On the logical unsolvability of the Gettier problem. Synthese, 142(1):61–79, October 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3ASYNT.0000047709.27594.c4.pdf.

Morganti:2004:PES

[439] Matteo Morganti. On the preferability of epistemic structural realism. Synthese, 142(1):81–107, October 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3ASYNT.0000047712.39407.c3.pdf.

LePoidevin:2004:PCT

[440] Robin Le Poidevin. A puzzle concerning time perception. Synthese, 142 (1):109–142, October 2004. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000047710.71824.b1.pdf.

Blutner:2004:NIN

[441] Reinhard Blutner. Nonmonotonic inferences and neural networks. Synthese, 142(2):143–174, September 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-1929-y;   http://link.springer.com/
content/pdf/10.1007/s11229-004-1929-y.pdf.

Dietrich:2004:MJD

[442] Franz Dietrich and Christian List. A model of jury decisions where all jurors have the same evidence. Synthese, 142(2):175–202, September 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-1276-z;   http://link.springer.com/content/pdf/10.1007/
s11229-004-1276-z.pdf.

Kacprzak:2004:SBA

[443] M. Kacprzak and W. Penczek. A SAT-based approach to unbounded model checking for alternating-time temporal epistemic logic. Synthese, 142(2):203–227, September 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-2446-8;   http://link.springer.com/
content/pdf/10.1007/s11229-004-2446-8.pdf.

Hommersom:2004:USS

[444] Arjen Hommersom, John-Jules Meyer, and Erik De Vink. Update semantics of security protocols. Synthese, 142(2):229–267, September 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-2247-0;   http://link.springer.com/content/pdf/10.1007/
s11229-004-2247-0.pdf.

Anonymous:2005:I

[445] Anonymous. Introduction. Synthese, 142(3):269–271, February 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-3715-x;   http://link.springer.com/content/pdf/10.1007/
s11229-005-3715-x.pdf.

Azzouni:2005:TQT

[446] Jody Azzouni. Tarski, Quine, and the transcendence of the vernacular “true”. Synthese, 142(3):273–288, February 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-3717-8;
 http://link.springer.com/content/pdf/10.1007/s11229-005-3717-
8.pdf.

Glanzberg:2005:TRH

[447] Michael Glanzberg. Truth, reflection, and hierarchies. Synthese, 142(3): 289–315, February 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-3718-7;   http://link.springer.com/content/
pdf/10.1007/s11229-005-3718-7.pdf.

Heck:2005:TD

[448] Richard G. Heck, Jr. Truth and disquotation. Synthese, 142(3):317– 352, February 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-3719-6;   http://link.springer.com/content/
pdf/10.1007/s11229-005-3719-6.pdf.

Hintikka:2005:TA

[449] Jaakko Hintikka. On Tarski’s assumptions. Synthese, 142(3):353– 369, February 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-3720-0;   http://link.springer.com/content/
pdf/10.1007/s11229-005-3720-0.pdf.

Mcgee:2005:TSE

[450] V. Mcgee. Tarski’s staggering existential assumptions. Synthese, 142(3): 371–387, February 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-3721-z;   http://link.springer.com/content/
pdf/10.1007/s11229-005-3721-z.pdf.

Halonen:2005:TTP

[451] Ilpo Halonen and Jaakko Hintikka. Toward a theory of the process of explanation. Synthese, 143(1–2):5–61, January 2005. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-3113-4;
 http://link.springer.com/content/pdf/10.1007/s11229-005-3113-
4.pdf.

Batens:2005:TPE

[452] Diderik Batens. The theory of the process of explanation generalized to include the inconsistent case. Synthese, 143(1–2):63–88, January 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-3114-3;   http://link.springer.com/content/pdf/10.1007/
s11229-005-3114-3.pdf.

Bird:2005:EM

[453] Alexander Bird. Explanation and metaphysics. Synthese, 143(1–2): 89–107, January 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-3115-2;   http://link.springer.com/content/
pdf/10.1007/s11229-005-3115-2.pdf.

Hitchcock:2005:LTE

[454] Christopher Hitchcock. … And away from a theory of explanation itself. Synthese, 143(1–2):109–124, January 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-3116-1;
 http://link.springer.com/content/pdf/10.1007/s11229-005-3116-
1.pdf.

Keranen:2005:ECV

[455] Jukka Keränen and Wesley Salmon. Explanatoriness: Cause versus Craig. Synthese, 143(1–2):125–147, January 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-3117-0;
 http://link.springer.com/content/pdf/10.1007/s11229-005-3117-
0.pdf.

Schurz:2005:ESL

[456] Gerhard Schurz. Explanations in science and the logic of whyquestions: Discussion of the Halonen–Hintikka-approach and alternative proposal. Synthese, 143(1–2):149–178, January 2005. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-3118-z;
 http://link.springer.com/content/pdf/10.1007/s11229-005-3118-
z.pdf.

Sintonen:2005:SEC

[457] Matti Sintonen. Scientific explanation: Conclusiveness conditions on explanation-seeking questions. Synthese, 143(1–2):179–205, January 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-3119-y;   http://link.springer.com/content/pdf/10.1007/
s11229-005-3119-y.pdf.

Hintikka:2005:ERR

[458] Jaakko Hintikka and Ilpo Halonen. Explanation: Retrospective reflections. Synthese, 143(1–2):207–222, January 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-3120-5;
 http://link.springer.com/content/pdf/10.1007/s11229-005-3120-
5.pdf.

Muller:2005:DS

[459] F. A. Muller. Deflating skolem. Synthese, 143(3):223–253, February 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-0800-0;   http://link.springer.com/content/pdf/10.1007/
s11229-005-0800-0.pdf.

Wechsler:2005:INI

[460] Sergio Wechsler, L. G. Esteves, A. Simonis, and C. Peixoto. Indifference, neutrality and informativeness: generalizing the three prisoners paradox. Synthese, 143(3):255–272, February 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-7016-1;
 http://link.springer.com/content/pdf/10.1007/s11229-005-7016-
1.pdf.

Wong:2005:SPI

[461] Wai-Hung Wong. The skeptical paradox and the indispensability of knowledge-beliefs. Synthese, 143(3):273–290, February 2005. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-7017-0;
 http://link.springer.com/content/pdf/10.1007/s11229-005-7017-
0.pdf.

Carter:2005:IOP

[462] Jessica Carter. Individuation of objects- a problem for structuralism? Synthese, 143(3):291–307, February 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-0848-x;
 http://link.springer.com/content/pdf/10.1007/s11229-005-0848-
x.pdf.

Brown:2005:BAP

[463] Campbell Brown and Yujin Nagasawa. The best of all possible worlds. Synthese, 143(3):309–320, February 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-7043-y;

 http://link.springer.com/content/pdf/10.1007/s11229-005-7043-
y.pdf.

Marconi:2005:TDS

[464] Diego Marconi. Two-dimensional semantics and the articulation problem. Synthese, 143(3):321–349, February 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-7044-x;
 http://link.springer.com/content/pdf/10.1007/s11229-005-7044-
x.pdf.

Hull:2005:BDH

[465] Gerald Hull. Bipolar disorder: Horgan on vagueness and incoherence. Synthese, 143(3):351–369, February 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-5069-9;
 http://link.springer.com/content/pdf/10.1007/s11229-005-5069-
9.pdf.

Deutsch:2005:II

[466] Max Deutsch. Intentionalism and intransitivity. Synthese, 144(1): 1–22, March 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-1022-1;   http://link.springer.com/content/
pdf/10.1007/s11229-005-1022-1.pdf.

Lombardi:2005:DST

[467] O. Lombardi. Dretske, Shannon’s theory and the interpretation of information. Synthese, 144(1):23–39, March 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-9127-0;
 http://link.springer.com/content/pdf/10.1007/s11229-005-9127-
0.pdf.

Wright:2005:DDU

[468] Wayne Wright. Distracted drivers and unattended experience. Synthese, 144(1):41–68, March 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-9128-z;   http://link.springer.com/content/
pdf/10.1007/s11229-005-9128-z.pdf.

Bloomfield:2005:LRA

[469] Paul Bloomfield. Let’s be realistic about serious metaphysics. Synthese, 144(1):69–90, March 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-9129-y;   http://link.springer.com/content/
pdf/10.1007/s11229-005-9129-y.pdf.

Bradley:2005:NDA

[470] D. J. Bradley. No Doomsday argument without knowledge of birth rank: a defense of Bostrom. Synthese, 144(1):91–100, March 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-9131-4;   http://link.springer.com/content/pdf/10.1007/
s11229-005-9131-4.pdf. See [126].

Byrne:2005:CMC

[471] Darragh Byrne. Compositionality and the manifestation challenge. Synthese, 144(1):101–136, March 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-0384-8;   http://link.springer.com/
content/pdf/10.1007/s11229-005-0384-8.pdf.

DeRegt:2005:CAS

[472] Henk W. De Regt and Dennis Dieks. A contextual approach to scientific understanding. Synthese, 144(1):137–170, March 2005. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-5000-4;
 http://link.springer.com/content/pdf/10.1007/s11229-005-5000-
4.pdf.

Ledwig:2005:NPA

[473] Marion Ledwig. The no probabilities for acts-principle. Synthese, 144 (2):171–180, March 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-2010-6;   http://link.springer.com/content/
pdf/10.1007/s11229-004-2010-6.pdf.

Jaeger:2005:LIP

[474] Manfred Jaeger. A logic for inductive probabilistic reasoning. Synthese, 144(2):181–248, March 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6153-2;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6153-2.pdf.

Hosni:2005:RC

[475] Hykel Hosni and Jeff Paris. Rationality as conformity. Synthese, 144 (2):249–285, March 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-4684-1;   http://link.springer.com/content/
pdf/10.1007/s11229-004-4684-1.pdf.

Wheeler:2005:SRA

[476] Gregory R. Wheeler. On the structure of rational acceptance: Comments on Hawthorne and bovens. Synthese, 144(2):287–304, March 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-2722-2;   http://link.springer.com/content/pdf/10.1007/
s11229-005-2722-2.pdf.

Mckitrick:2005:I

[477] Jennifer Mckitrick. Introduction. Synthese, 144(3):305–308, April 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-5847-4;   http://link.springer.com/content/pdf/10.1007/
s11229-005-5847-4.pdf.

Armstrong:2005:FDA

[478] David M. Armstrong. Four disputes about properties. Synthese, 144 (3):309–320, April 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-5852-7;   http://link.springer.com/content/
pdf/10.1007/s11229-005-5852-7.pdf.

Cross:2005:WD

[479] Troy Cross. what is a disposition? Synthese, 144(3):321–341, April 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-5857-2;   http://link.springer.com/content/pdf/10.1007/
s11229-005-5857-2.pdf.

Heil:2005:D

[480] John Heil. Dispositions. Synthese, 144(3):343–356, April 2005. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-5864-3;
 http://link.springer.com/content/pdf/10.1007/s11229-005-5864-
3.pdf.

Mckitrick:2005:DCR

[481] Jennifer Mckitrick. Are dispositions causally relevant? Synthese, 144 (3):357–371, April 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-5868-z;   http://link.springer.com/content/
pdf/10.1007/s11229-005-5868-z.pdf.

Anderson:2005:HGG

[482] Erik Anderson. How general is generalized scientific essentialism? Synthese, 144(3):373–379, April 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-5870-5;   http://link.springer.com/
content/pdf/10.1007/s11229-005-5870-5.pdf.

Drewery:2005:ENL

[483] Alice Drewery. Essentialism and the necessity of the laws of nature. Synthese, 144(3):381–396, April 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-5871-4;   http://link.springer.com/
content/pdf/10.1007/s11229-005-5871-4.pdf.

Mumford:2005:LL

[484] Stephen Mumford. Laws and lawlessness. Synthese, 144(3):397–413, April 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-5873-2;   http://link.springer.com/content/pdf/10.1007/
s11229-005-5873-2.pdf.

Lange:2005:LTS

[485] Marc Lange. laws and their stability. Synthese, 144(3):415–432, April 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-5874-1;   http://link.springer.com/content/pdf/10.1007/
s11229-005-5874-1.pdf.

Roberts:2005:MPL

[486] John T. Roberts. Measurability and physical laws. Synthese, 144(3): 433–447, April 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-5875-0;   http://link.springer.com/content/
pdf/10.1007/s11229-005-5875-0.pdf.

Anonymous:2005:CVa

[487] Anonymous. Contents of volume 144. Synthese, 144(3):449–450, April 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-8134-5;   http://link.springer.com/content/pdf/10.1007/
s11229-005-8134-5.pdf.

Anonymous:2005:AIa

[488] Anonymous. Author index. Synthese, 144(3):451, April 2005. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-8135-4;
 http://link.springer.com/content/pdf/10.1007/s11229-005-8135-
4.pdf.

Anonymous:2005:IAa

[489] Anonymous. Instructions for authors. Synthese, 144(3):452–456, April 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-8136-3;   http://link.springer.com/content/pdf/10.1007/
s11229-005-8136-3.pdf.

Wright:2005:FPA

[490] Cory D. Wright. On the functionalization of pluralist approaches to truth. Synthese, 145(1):1–28, May 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-5863-9;   http://link.springer.com/
content/pdf/10.1007/s11229-004-5863-9.pdf.

Lynch:2005:AFO

[491] M. P. Lynch. Alethic functionalism and our folk theory of truth. Synthese, 145(1):29–43, May 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-1771-2;   http://link.springer.com/
content/pdf/10.1007/s11229-004-1771-2.pdf.

Herstein:2005:DIP

[492] G. L. Herstein. Davidson on the impossibility of psychophysical laws. Synthese, 145(1):45–63, May 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-5866-6;   http://link.springer.com/
content/pdf/10.1007/s11229-004-5866-6.pdf.

Wikforss:2005:NNK

[493] Åsa Maria Wikforss. Naming natural kinds. Synthese, 145(1):65–87, May 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-7013-9;   http://link.springer.com/content/pdf/10.1007/
s11229-004-7013-9.pdf.

Albers:2005:TRT

[494] Casper J. Albers, Barteld P. Kooi, and Willem Schaafsma. Trying to resolve the two-envelope problem. Synthese, 145(1):89–109, May 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-7665-5;   http://link.springer.com/content/pdf/10.1007/
s11229-004-7665-5.pdf.

Hung:2005:PEH

[495] Edwin H.-C. Hung. Projective explanation: How theories explain empirical data in spite of theory–data incommensurability. Synthese, 145 (1):111–129, May 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-0034-6;   http://link.springer.com/content/
pdf/10.1007/s11229-004-0034-6.pdf.

Boogerd:2005:EPN

[496] F. C. Boogerd, F. J. Bruggeman, R. C. Richardson, A. Stephan, and H. V. Westerhoff. Emergence and its place in nature: A case study of biochemical networks. Synthese, 145(1):131–164, May 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-4421-9;   http://link.springer.com/content/pdf/10.1007/
s11229-004-4421-9.pdf.

Guenin:2005:I

[497] Louis M. Guenin. Introduction. Synthese, 145(2):165–168, June 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-3744-5;   http://link.springer.com/content/pdf/10.1007/
s11229-005-3744-5.pdf.

Hintikka:2005:ODE

[498] Jaakko Hintikka. Omitting data — ethical or strategic problem? Synthese, 145(2):169–176, June 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-3745-4;   http://link.springer.com/
content/pdf/10.1007/s11229-005-3745-4.pdf.

Guenin:2005:IH

[499] Louis M. Guenin. Intellectual honesty. Synthese, 145(2):177–232, June 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-3746-3;   http://link.springer.com/content/pdf/10.1007/
s11229-005-3746-3.pdf.

Roush:2005:TC

[500] Sherrilyn Roush. Testability and candor. Synthese, 145(2):233–275, June 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-3748-1;   http://link.springer.com/content/pdf/10.1007/
s11229-005-3748-1.pdf.

Holton:2005:CIS

[501] Gerald Holton. Candor and integrity in science. Synthese, 145(2): 277–294, June 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-3749-0;   http://link.springer.com/content/
pdf/10.1007/s11229-005-3749-0.pdf.

Ketland:2005:YPI

[502] Jeffrey Ketland. Yablo’s paradox and ω-inconsistency. Synthese, 145 (3):295–302, July 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-6201-6;   http://link.springer.com/content/
pdf/10.1007/s11229-005-6201-6.pdf.

Asher:2005:FCP

[503] Nicholas Asher and Daniel Bonevac. Free choice permission is strong permission. Synthese, 145(3):303–323, July 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-6196-z;
 http://link.springer.com/content/pdf/10.1007/s11229-005-6196-
z.pdf.

Colyvan:2005:PAF

[504] Mark Colyvan, Jay L. Garfield, and Graham Priest. Problems with the argument from fine tuning. Synthese, 145(3):325–338, July 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-6195-0;   http://link.springer.com/content/pdf/10.1007/
s11229-005-6195-0.pdf.

Walton:2005:ASE

[505] D. Walton and C. A. Reed. Argumentation schemes and enthymemes. Synthese, 145(3):339–370, July 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-6198-x;   http://link.springer.com/
content/pdf/10.1007/s11229-005-6198-x.pdf.

Novaes:2005:MOL

[506] C. Dutilh Novaes. Medieval obligationes as logical games of consistency maintenance. Synthese, 145(3):371–395, July 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-6197-y;
 http://link.springer.com/content/pdf/10.1007/s11229-005-6197-
y.pdf.

Narens:2005:TBS

[507] Louis Narens. A theory of belief for scientific refutations. Synthese, 145(3):397–423, July 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-6199-9;   http://link.springer.com/content/
pdf/10.1007/s11229-005-6199-9.pdf.

Levine:2005:CME

[508] Alex Levine. Conjoining mathematical empiricism with mathematical realism: Maddy’s account of set perception revisited. Synthese, 145 (3):425–448, July 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-6202-5;   http://link.springer.com/content/
pdf/10.1007/s11229-005-6202-5.pdf.

Fahrbach:2005:UBF

[509] Ludwig Fahrbach. Understanding brute facts. Synthese, 145(3):449–466, July 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-6200-7;   http://link.springer.com/content/pdf/10.1007/
s11229-005-6200-7.pdf.

Lawlor:2005:RPR

[510] Krista Lawlor. Reason and the past: The role of rationality in diachronic self-knowledge. Synthese, 145(3):467–495, July 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-6220-3;   http://link.springer.com/content/pdf/10.1007/
s11229-005-6220-3.pdf.

Pambuccian:2005:CAH

[511] Victor Pambuccian. Correction to “Axiomatizations of Hyperbolic Geometry”. Synthese, 145(3):497, July 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-5405-5;   http://link.springer.com/
content/pdf/10.1007/s11229-004-5405-5.pdf. See [250].

Anonymous:2005:CVb

[512] Anonymous. Contents of volume 145. Synthese, 145(3):499–500, July 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-8687-8;   http://link.springer.com/content/pdf/10.1007/
s11229-004-8687-8.pdf.

Anonymous:2005:AIb

[513] Anonymous. Author index. Synthese, 145(3):501, July 2005. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-8688-7;
 http://link.springer.com/content/pdf/10.1007/s11229-004-8688-
7.pdf.

Anonymous:2005:IAb

[514] Anonymous. Instructions for authors. Synthese, 145(3):502–506, July 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-8689-6;   http://link.springer.com/content/pdf/10.1007/
s11229-004-8689-6.pdf.

Schurz:2005:I

[515] Gerhard Schurz and Hannes Leitgeb. Introduction. Synthese, 146(1– 2):1–5, August 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-9061-1;   http://link.springer.com/content/
pdf/10.1007/s11229-005-9061-1.pdf.

Pelletier:2005:CPD

[516] Francis Jeffry Pelletier and Renée Elio. The case for psychologism in default and inheritance reasoning. Synthese, 146(1–2):7–35, August 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-9063-z;   http://link.springer.com/content/pdf/10.1007/
s11229-005-9063-z.pdf.

Schurz:2005:NMR

[517] Gerhard Schurz. Non-monotonic reasoning from an evolution-theoretic perspective: Ontic, logical and cognitive foundations. Synthese, 146 (1–2):37–51, August 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-9067-8;   http://link.springer.com/content/
pdf/10.1007/s11229-005-9067-8.pdf.

Benferhat:2005:OPH

[518] Salem Benferhat, Jean F. Bonnefon, and Rui da Silva Neves. An overview of possibilistic handling of default reasoning, with experimental studies. Synthese, 146(1–2):53–70, August 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-9069-6;
 http://link.springer.com/content/pdf/10.1007/s11229-005-9069-
6.pdf.

Ford:2005:HNR

[519] Marilyn Ford. Human nonmonotonic reasoning: the importance of seeing the logical strength of arguments. Synthese, 146(1–2):71–92, August 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-9071-z;   http://link.springer.com/content/pdf/10.1007/
s11229-005-9071-z.pdf.

Pfeifer:2005:CNH

[520] Niki Pfeifer and Gernot D. Kleiter. Coherence and nonmonotonicity in human reasoning. Synthese, 146(1–2):93–109, August 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-9073-x;   http://link.springer.com/content/pdf/10.1007/
s11229-005-9073-x.pdf.

Horsten:2005:QSI

[521] Leon Horsten. On the quantitative scalar or-implicature. Synthese, 146 (1–2):111–127, August 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-9076-7;   http://link.springer.com/content/
pdf/10.1007/s11229-005-9076-7.pdf.

Adams:2005:PAS

[522] Ernest W. Adams. On a proportionality analysis of syllogistic private reasoning. Synthese, 146(1–2):129–138, August 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-9078-5;
 http://link.springer.com/content/pdf/10.1007/s11229-005-9078-
5.pdf.

Gilio:2005:PLU

[523] Angelo Gilio. Probabilistic logic under coherence, conditional interpretations, and default reasoning. Synthese, 146(1–2):139–152, August 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-9080-y;   http://link.springer.com/content/pdf/10.1007/
s11229-005-9080-y.pdf.

Lukasiewicz:2005:NPR

[524] Thomas Lukasiewicz. Nonmonotonic probabilistic reasoning under variable-strength inheritance with overriding. Synthese, 146(1–2):153– 169, August 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-9082-9;   http://link.springer.com/content/
pdf/10.1007/s11229-005-9082-9.pdf.

Brewka:2005:ASQ

[525] Gerhard Brewka. Answer sets and qualitative decision making. Synthese, 146(1–2):171–187, August 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-9084-7;   http://link.springer.com/
content/pdf/10.1007/s11229-005-9084-7.pdf.

Leitgeb:2005:IDS

[526] Hannes Leitgeb. Interpreted dynamical systems and qualitative laws: from neural networks to evolutionary systems. Synthese, 146(1–2): 189–202, August 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-9086-5;   http://link.springer.com/content/
pdf/10.1007/s11229-005-9086-5.pdf.

Werning:2005:TDT

[527] Markus Werning. The temporal dimension of thought. Synthese, 146(1– 2):203–224, August 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-9089-2;   http://link.springer.com/content/
pdf/10.1007/s11229-005-9089-2.pdf.

Dilworth:2005:DCP

[528] John Dilworth. The double content of perception. Synthese, 146(3): 225–243, September 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6209-3;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6209-3.pdf.

Ben-Menahem:2005:BWG

[529] Yemima Ben-Menahem. Black, white and gray: Quine on convention. Synthese, 146(3):245–282, September 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6211-9;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6211-
9.pdf.

Iacona:2005:RB

[530] A. Iacona. Rethinking bivalence. Synthese, 146(3):283–302, September 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-6237-7;   http://link.springer.com/content/pdf/10.1007/
s11229-005-6237-7.pdf.

Williams:2005:SDD

[531] Neil Edward Williams. Static and dynamic dispositions. Synthese, 146(3):303–324, September 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-6212-8;   http://link.springer.com/
content/pdf/10.1007/s11229-004-6212-8.pdf.

Balaguer:2005:IPR

[532] Mark Balaguer. Indexical propositions and De Re belief ascriptions. Synthese, 146(3):325–355, September 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-6217-3;   http://link.springer.com/
content/pdf/10.1007/s11229-004-6217-3.pdf.

Hofmann:2005:EME

[533] Frank Hofmann. Epistemic means and ends: In defense of some Sartwellian insights. Synthese, 146(3):357–369, September 2005. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6210-x;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6210-
x.pdf.

DeClercq:2005:C

[534] Rafael De Clercq and Leon Horsten. Closer. Synthese, 146(3):371– 393, September 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6219-1;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6219-1.pdf.

Comesana:2005:UK

[535] Juan Comesaña. Unsafe knowledge. Synthese, 146(3):395–404, September 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6213-7;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6213-7.pdf.

Vision:2005:TJB

[536] G. Vision. Truly justified belief. Synthese, 146(3):405–446, September 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6228-0;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6228-0.pdf.

Cat:2005:MCC

[537] Jordi Cat. Modeling cracks and cracking models: Structures, mechanisms, boundary conditions, constraints, inconsistencies and the proper domains of natural laws. Synthese, 146(3):447–487, September 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6259-6;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6259-6.pdf.

Anonymous:2005:CVc

[538] Anonymous. Contents of volume 146. Synthese, 146(3):489–490, September 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-1328-z;   http://link.springer.com/content/pdf/10.1007/
s11229-005-1328-z.pdf.

Anonymous:2005:AIc

[539] Anonymous. Author index. Synthese, 146(3):491, September 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-1329-y;   http://link.springer.com/content/pdf/10.1007/
s11229-005-1329-y.pdf.

Buldt:2005:RFH

[540] Bernd Buldt, Volker Halbach, and Reinhard Kahle. Reflections on Frege and Hilbert. Synthese, 147(1):1–2, October 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6203-9;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6203-
9.pdf.

Ferreira:2005:AFG

[541] Fernando Ferreira. Amending Frege’s Grundgesetze der Arithmetik. Synthese, 147(1):3–19, October 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-6204-8;   http://link.springer.com/
content/pdf/10.1007/s11229-004-6204-8.pdf.

Hale:2005:RNS

[542] Bob Hale. Real numbers and set theory — extending the neo-Fregean programme beyond arithmetic. Synthese, 147(1):21–41, October 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6205-7;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6205-7.pdf.

Wehmeier:2005:FPA

[543] Kai Frederick Wehmeier and Peter Schroeder-Heister. Frege’s permutation argument revisited. Synthese, 147(1):43–61, October 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

004-6206-6;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6206-6.pdf.

Carty:2005:PRH

[544] D. C. Mc. Carty. Problems and riddles: Hilbert and the du Bois-Reymonds. Synthese, 147(1):63–79, October 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6207-5;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6207-
5.pdf.

Rathjen:2005:CHP

[545] Michael Rathjen. The constructive Hilbert program and the limits of Martin-Löf type theory. Synthese, 147(1):81–120, October 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6208-4;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6208-4.pdf.

Sieg:2005:DAN

[546] Wilfried Sieg and Dirk Schlimm. Dedekind’s analysis of number: Systems and axioms. Synthese, 147(1):121–170, October 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6300-9;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6300-9.pdf.

Schlimm:2005:AAI

[547] Dirk Schlimm. Against against intuitionism. Synthese, 147(1):171– 188, October 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6299-y;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6299-y.pdf.

Bonanno:2005:LFT

[548] Giacomo Bonanno. Logic and the foundations of the theory of games and decisions: Introduction^{∗}. Synthese, 147(2):189–192, November 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-1346-x;   http://link.springer.com/content/pdf/10.1007/
s11229-005-1346-x.pdf.

Bonanno:2005:SML

[549] Giacomo Bonanno. A simple modal logic for belief revision. Synthese, 147(2):193–228, November 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-1348-8;   http://link.springer.com/
content/pdf/10.1007/s11229-005-1348-8.pdf.

VanDitmarsch:2005:PDL

[550] Hans P. Van Ditmarsch. Prolegomena to dynamic logic for belief revision. Synthese, 147(2):229–275, November 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-1349-7;
 http://link.springer.com/content/pdf/10.1007/s11229-005-1349-
7.pdf.

Laverny:2005:KBP

[551] Noël Laverny and Jérôme Lang. From knowledge-based programs to graded belief-based programs, part i: On-line reasoning^{∗}. Synthese, 147(2):277–321, November 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-1350-1;   http://link.springer.com/
content/pdf/10.1007/s11229-005-1350-1.pdf.

Peterson:2005:OIT

[552] Martin Peterson and Sven Ove Hansson. Order-independent transformative decision rules. Synthese, 147(2):323–342, November 2005. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-1351-0;
 http://link.springer.com/content/pdf/10.1007/s11229-005-1351-
0.pdf.

Schulz:2005:PSP

[553] Katrin Schulz. A pragmatic solution for the paradox of free choice permission. Synthese, 147(2):343–377, November 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-1353-y;
 http://link.springer.com/content/pdf/10.1007/s11229-005-1353-
y.pdf.

Sillari:2005:LFC

[554] Giacomo Sillari. A logical framework for convention. Synthese, 147(2): 379–400, November 2005. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-1352-z;   http://link.springer.com/content/
pdf/10.1007/s11229-005-1352-z.pdf.

Bickle:2005:I

[555] J. Bickle. Introduction. Synthese, 147(3):401–402, December 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-8362-8;   http://link.springer.com/content/pdf/10.1007/
s11229-005-8362-8.pdf.

Chemero:2005:OEP

[556] Anthony Chemero and Charles Heyser. Object exploration and a problem with reductionism. Synthese, 147(3):403–423, December 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-8363-7;   http://link.springer.com/content/pdf/10.1007/
s11229-005-8363-7.pdf.

Musacchio:2005:WDQ

[557] José M. Musacchio. Why do qualia and the mind seem nonphysical? Synthese, 147(3):425–460, December 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-8364-6;
 http://link.springer.com/content/pdf/10.1007/s11229-005-8364-
6.pdf.

Gray:2005:CS

[558] R. Gray. On the concept of a sense. Synthese, 147(3):461–475, December 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-1334-1;   http://link.springer.com/content/pdf/10.1007/
s11229-005-1334-1.pdf.

Horst:2005:MLE

[559] Steven Horst. Modeling, localization and the explanation of phenomenal properties: Philosophy and the cognitive sciences at the beginning of the Millennium. Synthese, 147(3):477–513, December 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-8365-5;   http://link.springer.com/content/pdf/10.1007/
s11229-005-8365-5.pdf.

Vacariu:2005:MBE

[560] Gabriel Vacariu. Mind, brain, and epistemologically different worlds. Synthese, 147(3):515–548, December 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-005-8366-4;   http://link.springer.
com/content/pdf/10.1007/s11229-005-8366-4.pdf.

Anonymous:2005:CVd

[561] Anonymous. Contents of volume 147. Synthese, 147(3):549–550, December 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-4774-8;   http://link.springer.com/content/pdf/10.1007/
s11229-005-4774-8.pdf.

Anonymous:2005:AId

[562] Anonymous. Author index. Synthese, 147(3):551, December 2005. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-4773-9;
 http://link.springer.com/content/pdf/10.1007/s11229-005-4773-
9.pdf.

Anonymous:2005:IAc

[563] Anonymous. Instructions for authors. Synthese, 147(3):553–558, December 2005. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-5462-4;   http://link.springer.com/content/pdf/10.1007/
s11229-005-5462-4.pdf.

vanBenthem:2006:E

[564] Johan van Benthem, Vincent F. Hendricks, and John Symons. Editorial. Synthese, 148(1):1–3, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-5892-z;   http://link.springer.com/
content/pdf/10.1007/s11229-005-5892-z.pdf.

Predelli:2006:PTR

[565] Stefano Predelli. The problem with token-reflexivity. Synthese, 148 (1):5–29, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-003-6194-y;   http://link.springer.com/content/
pdf/10.1007/s11229-003-6194-y.pdf.

Tomassi:2006:TWS

[566] Paul Tomassi. Truth, warrant and superassertibility. Synthese, 148 (1):31–56, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-003-6227-6;   http://link.springer.com/content/
pdf/10.1007/s11229-003-6227-6.pdf.

Orilia:2006:QMS

[567] Francesco Orilia. Quantum-mechanical statistics and the inclusivist approach to the nature of particulars. Synthese, 148(1):57–77, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6216-4;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6216-4.pdf.

DePinedo:2006:AMO

[568] M. De Pinedo. Anomalous monism: Oscillating between dogmas. Synthese, 148(1):79–97, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-6218-2;   http://link.springer.com/
content/pdf/10.1007/s11229-004-6218-2.pdf.

Bassler:2006:SMP

[569] O. Bradley Bassler. The surveyability of mathematical proof: A historical perspective. Synthese, 148(1):99–133, January 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6221-7;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6221-
7.pdf.

Zalabardo:2006:BER

[570] José L. Zalabardo. Bonjour, externalism and the regress problem*. Synthese, 148(1):135–169, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-2245-2;   http://link.springer.com/
content/pdf/10.1007/s11229-004-2245-2.pdf.

Schirn:2006:HPA

[571] Matthias Schirn. Hume’s principle and axiom v reconsidered: Critical reflections on Frege and his interpreters. Synthese, 148(1):171–227, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

004-2829-x;   http://link.springer.com/content/pdf/10.1007/
s11229-004-2829-x.pdf.

Nathan:2006:PD

[572] Amos Nathan. Probability dynamics. Synthese, 148(1):229–256, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-0197-9;   http://link.springer.com/content/pdf/10.1007/
s11229-005-0197-9.pdf.

Anonymous:2006:Ea

[573] Anonymous. Erratum. Synthese, 148(1):257, January 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-7147-z;
 http://link.springer.com/content/pdf/10.1007/s11229-006-7147-
z.pdf.

Bandyopadhyay:2006:AES

[574] Prasanta S. Bandyopadhyay and Gordon G. Brittan, Jr. Acceptibility, evidence, and severity. Synthese, 148(2):259–293, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6222-6;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6222-6.pdf.

Magnus:2006:WNA

[575] P. D. Magnus. What’s new about the new induction? Synthese, 148 (2):295–301, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6223-5;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6223-5.pdf.

Reed:2006:SCH

[576] Baron Reed. Shelter for the cognitively homeless. Synthese, 148(2): 303–308, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6224-4;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6224-4.pdf.

Zahn:2006:NMC

[577] Peter Zahn. A normative model of classical reasoning in higher order languages. Synthese, 148(2):309–343, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6225-3;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6225-
3.pdf.

Glock:2006:TT

[578] Hans Johann Glock. Truth in the Tractatus. Synthese, 148(2):345– 368, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6226-2;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6226-2.pdf.

Choi:2006:SVR

[579] Sungho Choi. The simple vs. reformed conditional analysis of dispositions. Synthese, 148(2):369–379, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6229-z;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6229-
z.pdf.

Weinstein:2006:SSR

[580] Steven Weinstein. Superluminal signaling and relativity. Synthese, 148 (2):381–399, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6231-5;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6231-5.pdf.

Moyer:2006:SLS

[581] Mark Moyer. Statues and lumps: A strange coincidence? Synthese, 148 (2):401–423, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6235-1;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6235-1.pdf.

Hansson:2006:CSV

[582] Sven Ove Hansson. Category-specified value statements. Synthese, 148 (2):425–432, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6236-0;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6236-0.pdf.

Laraudogoitia:2006:LSR

[583] J. P. Laraudogoitia. A look at the Staccato run. Synthese, 148(2): 433–441, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6238-y;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6238-y.pdf.

Earman:2006:TCR

[584] J. Earman. Two challenges to the requirement of substantive general covariance. Synthese, 148(2):443–468, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6239-x;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6239-
x.pdf.

Fischer:2006:PP

[585] Eugen Fischer. Philosophical pictures. Synthese, 148(2):469–501, January 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-1007-0;   http://link.springer.com/content/pdf/10.1007/
s11229-005-1007-0.pdf.

Kahle:2006:IPT

[586] Reinhard Kahle and Peter Schroeder-Heister. Introduction: Prooftheoretic semantics. Synthese, 148(3):503–506, February 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6292-5;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6292-
5.pdf.

Prawitz:2006:MAP

[587] Dag Prawitz. Meaning approached via proofs. Synthese, 148(3):507– 524, February 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6295-2;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6295-2.pdf.

Schroeder-Heister:2006:VCP

[588] Peter Schroeder-Heister. Validity concepts in proof-theoretic semantics. Synthese, 148(3):525–571, February 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-6296-1;   http://link.springer.com/
content/pdf/10.1007/s11229-004-6296-1.pdf.

Contu:2006:JLL

[589] Patrizio Contu. The justification of the logical laws revisited. Synthese, 148(3):573–588, February 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-6270-y;   http://link.springer.com/
content/pdf/10.1007/s11229-004-6270-y.pdf.

Hallnas:2006:PTF

[590] Lars Hallnäs. On the proof-theoretic foundation of general definition theory. Synthese, 148(3):589–602, February 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6291-6;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6291-
6.pdf.

Tait:2006:PTS

[591] William W. Tait. Proof-theoretic semantics for classical mathematics. Synthese, 148(3):603–622, February 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-6271-x;   http://link.springer.com/
content/pdf/10.1007/s11229-004-6271-x.pdf.

Sundholm:2006:SVN

[592] Göran Sundholm. Semantic values for natural deduction derivations. Synthese, 148(3):623–638, February 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-6298-z;   http://link.springer.com/
content/pdf/10.1007/s11229-004-6298-z.pdf.

Dosen:2006:MD

[593] Kosta Dosen. Models of deduction*. Synthese, 148(3):639–657, February 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6290-7;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6290-7.pdf.

Kahle:2006:PTV

[594] Reinhard Kahle. A proof-theoretic view of necessity. Synthese, 148(3): 659–673, February 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6293-4;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6293-4.pdf.

Usberti:2006:TSB

[595] Gabriele Usberti. Towards a semantics based on the notion of justification. Synthese, 148(3):675–699, February 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6272-9;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6272-
9.pdf.

Mints:2006:NCN

[596] Grigori Mints. Notes on constructive negation. Synthese, 148(3): 701–717, February 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6294-3;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6294-3.pdf.

Rathjen:2006:TOP

[597] Michael Rathjen. Theories and ordinals in proof theory. Synthese, 148 (3):719–743, February 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6297-0;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6297-0.pdf.

Anonymous:2006:CVa

[598] Anonymous. Contents of volume 148. Synthese, 148(3):745–747, February 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-8040-5;   http://link.springer.com/content/pdf/10.1007/
s11229-006-8040-5.pdf.

Anonymous:2006:AIa

[599] Anonymous. Author index. Synthese, 148(3):749, February 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-8041-4;
 http://link.springer.com/content/pdf/10.1007/s11229-006-8041-
4.pdf.

Anonymous:2006:IAa

[600] Anonymous. Instructions for authors. Synthese, 148(3):751–756, February 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-8042-3;   http://link.springer.com/content/pdf/10.1007/
s11229-006-8042-3.pdf.

Weiner:2006:HCP

[601] Matthew Weiner and Nuel Belnap. How causal probabilities might fit into our objectively indeterministic world. Synthese, 149(1):1–36, March 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6240-4;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6240-4.pdf.

Lynch:2006:ZCP

[602] M. P. Lynch. Zombies and the case of the phenomenal pickpocket. Synthese, 149(1):37–58, March 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-6241-3;   http://link.springer.com/
content/pdf/10.1007/s11229-004-6241-3.pdf.

Zimmerman:2006:SVC

[603] Aaron Z. Zimmerman. Self-verification and the content of thought. Synthese, 149(1):59–75, March 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-6247-x;   http://link.springer.com/
content/pdf/10.1007/s11229-004-6247-x.pdf.

Prosser:2006:TMZ

[604] Simon Prosser. Temporal metaphysics in z-land. Synthese, 149(1): 77–96, March 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6249-8;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6249-8.pdf.

Douven:2006:BCM

[605] Igor Douven and Wouter Meijs. Bootstrap confirmation made quantitative. Synthese, 149(1):97–132, March 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6250-2;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6250-
2.pdf.

Thalos:2006:NP

[606] Mariam Thalos. Nonreductive physics. Synthese, 149(1):133–178, March 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6251-1;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6251-1.pdf.

Moltmann:2006:PQD

[607] Friederike Moltmann. Presuppositions and quantifier domains. Synthese, 149(1):179–224, March 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6254-y;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6254-y.pdf.

Williams:2006:WMA

[608] John N. Williams. Wittgenstein, Moorean absurdity and its disappearance from speech. Synthese, 149(1):225–254, March 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6252-0;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6252-
0.pdf.

VanOtterloo:2006:F

[609] Sieuwert Van Otterloo, Michael Wooldridge, and Peter Mcburney. Foreword. Synthese, 149(2):255–256, March 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-4332-4;
 http://link.springer.com/content/pdf/10.1007/s11229-005-4332-
4.pdf.

Sevenster:2006:CCI

[610] Merlijn Sevenster. On the computational consequences of independence in propositional logic. Synthese, 149(2):257–283, March 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-3878-5;   http://link.springer.com/content/pdf/10.1007/
s11229-005-3878-5.pdf.

Dechesne:2006:TTI

[611] Francien Dechesne. Thompson transformations for if-logic. Synthese, 149 (2):285–309, March 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-3876-7;   http://link.springer.com/content/
pdf/10.1007/s11229-005-3876-7.pdf.

Pacuit:2006:LKB

[612] Eric Pacuit, Rohit Parikh, and Eva Cogan. The logic of knowledge based obligation. Synthese, 149(2):311–341, March 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-3877-6;
 http://link.springer.com/content/pdf/10.1007/s11229-005-3877-
6.pdf.

VanDitmarsch:2006:LP

[613] Hans P. Van Ditmarsch. The logic of pit. Synthese, 149(2):343– 374, March 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-4331-5;   http://link.springer.com/content/
pdf/10.1007/s11229-005-4331-5.pdf.

Aagotnes:2006:AKA

[614] Thomas Ågotnes. Action and knowledge in alternating-time temporal logic. Synthese, 149(2):375–407, March 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-3875-8;
 http://link.springer.com/content/pdf/10.1007/s11229-005-3875-
8.pdf.

Teepe:2006:PPA

[615] Wouter Teepe. Proving possession of arbitrary secrets while not giving them away: New protocols and a proof in GNY logic. Synthese, 149 (2):409–443, March 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-3879-4;   http://link.springer.com/content/
pdf/10.1007/s11229-005-3879-4.pdf.

Bird:2006:I

[616] Alexander Bird and Johannes Persson. Introduction. Synthese, 149 (3):445–450, April 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-0568-2;   http://link.springer.com/content/
pdf/10.1007/s11229-005-0568-2.pdf.

Hawley:2006:SGM

[617] Katherine Hawley. Science as a guide to metaphysics? Synthese, 149 (3):451–470, April 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-0569-1;   http://link.springer.com/content/
pdf/10.1007/s11229-005-0569-1.pdf.

Mumford:2006:UA

[618] Stephen Mumford. The ungrounded argument. Synthese, 149(3): 471–489, April 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-0570-8;   http://link.springer.com/content/
pdf/10.1007/s11229-005-0570-8.pdf.

Bird:2006:PM

[619] Alexander Bird. Potency and modality. Synthese, 149(3):491–508, April 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-0574-4;   http://link.springer.com/content/pdf/10.1007/
s11229-005-0574-4.pdf.

Beebee:2006:DAH

[620] Helen Beebee. Does anything hold the universe together? Synthese, 149(3):509–533, April 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-0576-2;   http://link.springer.com/content/
pdf/10.1007/s11229-005-0576-2.pdf.

Persson:2006:CC

[621] Johannes Persson. Compartment causation. Synthese, 149(3):535–550, April 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-0577-1;   http://link.springer.com/content/pdf/10.1007/
s11229-005-0577-1.pdf.

Noordhof:2006:EDC

[622] Paul Noordhof. Environment-dependent content and the virtues of causal explanation. Synthese, 149(3):551–575, April 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-0579-z;
 http://link.springer.com/content/pdf/10.1007/s11229-005-0579-
z.pdf.

Anonymous:2006:CVb

[623] Anonymous. Contents of volume 149. Synthese, 149(3):577–578, April 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-8085-5;   http://link.springer.com/content/pdf/10.1007/
s11229-006-8085-5.pdf.

Anonymous:2006:AIb

[624] Anonymous. Author index. Synthese, 149(3):579, April 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-8086-4;
 http://link.springer.com/content/pdf/10.1007/s11229-006-8086-
4.pdf.

Anonymous:2006:IAb

[625] Anonymous. Instructions for authors. Synthese, 149(3):581–586, April 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-8087-3;   http://link.springer.com/content/pdf/10.1007/
s11229-006-8087-3.pdf.

Kastner:2006:CTI

[626] Ruth E. Kastner. Cramer’s transactional interpretation and causal loop problems. Synthese, 150(1):1–14, May 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-004-6264-9;   http://link.springer.
com/content/pdf/10.1007/s11229-004-6264-9.pdf.

Heitner:2006:PLP

[627] Reese M. Heitner. From a phono-logical point of view: Neutralizing Quine’s argument against analyticity. Synthese, 150(1):15–39, May 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6253-z;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6253-z.pdf.

Hetherington:2006:KBP

[628] Stephen Hetherington. Knowledge’s boundary problem. Synthese, 150 (1):41–56, May 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6255-x;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6255-x.pdf.

Wasserman:2006:FSO

[629] Ryan Wasserman. The future similarity objection revisited. Synthese, 150(1):57–67, May 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6256-9;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6256-9.pdf.

Delancey:2006:OTD

[630] Craig S. Delancey. Ontology and teleofunctions: A defense and revision of the systematic account of teleological explanation. Synthese, 150(1):69–98, May 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6257-8;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6257-8.pdf.

Marcus:2006:ESM

[631] Eric Marcus. Events, sortals, and the mind–body problem. Synthese, 150(1):99–129, May 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6258-7;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6258-7.pdf.

Bovens:2006:DAC

[632] Luc Bovens and Wlodek Rabinowicz. Democratic answers to complex questions- an epistemic perspective. Synthese, 150(1):131–153, May 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-0005-1;   http://link.springer.com/content/pdf/10.1007/
s11229-006-0005-1.pdf.

Haukioja:2006:PR

[633] Jussi Haukioja. Proto-rigidity*. Synthese, 150(2):155–169, May 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6263-x;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6263-x.pdf.

Gert:2006:PMT

[634] Joshua Gert. Problems for moral twin Earth arguments*. Synthese, 150(2):171–183, May 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-6262-y;   http://link.springer.com/content/
pdf/10.1007/s11229-004-6262-y.pdf.

Norton:2006:HFE

[635] John D. Norton. How the formal equivalence of grue and green defeats what is new in the new riddle of induction. Synthese, 150(2):185–207, May 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

004-6261-z;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6261-z.pdf.

Green:2006:PSF

[636] Karen Green. A pinch of salt for Frege. Synthese, 150(2):209–228, May 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6260-0;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6260-0.pdf.

Dubucs:2006:BAE

[637] Jacques Dubucs and Sandra Lapointe. On Bolzano’s alleged explicativism. Synthese, 150(2):229–246, May 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-004-6265-8;   http://link.springer.
com/content/pdf/10.1007/s11229-004-6265-8.pdf.

Neta:2006:EFN

[638] Ram Neta. Epistemology factualized: New contractarian foundations for epistemology. Synthese, 150(2):247–280, May 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-004-6266-7;
 http://link.springer.com/content/pdf/10.1007/s11229-004-6266-
7.pdf.

Georgalis:2006:RFP

[639] Nicholas Georgalis. Representation and the first-person perspective. Synthese, 150(2):281–325, May 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-6268-5;   http://link.springer.com/
content/pdf/10.1007/s11229-004-6268-5.pdf.

Brauner:2006:P

[640] Torben Braüner, Per Hasle, and Peter Øhrstrøm. Preface. Synthese, 150(3):327–328, June 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-5510-0;   http://link.springer.com/content/
pdf/10.1007/s11229-005-5510-0.pdf.

Blackburn:2006:APH

[641] Patrick Blackburn. Arthur Prior and hybrid logic. Synthese, 150(3): 329–372, June 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-5512-y;   http://link.springer.com/content/
pdf/10.1007/s11229-005-5512-y.pdf.

Copeland:2006:MPH

[642] B. Jack Copeland. Meredith, Prior, and the history of possible worlds semantics. Synthese, 150(3):373–397, June 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-5514-9;
 http://link.springer.com/content/pdf/10.1007/s11229-005-5514-
9.pdf.

Fine:2006:RT

[643] Kit Fine. The reality of tense. Synthese, 150(3):399–414, June 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-5515-8;   http://link.springer.com/content/pdf/10.1007/
s11229-005-5515-8.pdf.

Galton:2006:OVA

[644] Antony Galton. Operators vs. arguments: The ins and outs of reification. Synthese, 150(3):415–441, June 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-5516-7;
 http://link.springer.com/content/pdf/10.1007/s11229-005-5516-
7.pdf.

Simons:2006:LL

[645] Peter Simons. The logic of location. Synthese, 150(3):443–458, June 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-5517-6;   http://link.springer.com/content/pdf/10.1007/
s11229-005-5517-6.pdf.

Wansing:2006:LCC

[646] Heinrich Wansing. Logical connectives for constructive modal logic. Synthese, 150(3):459–482, June 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-5518-5;   http://link.springer.com/
content/pdf/10.1007/s11229-005-5518-5.pdf.

Zanardo:2006:MHD

[647] Alberto Zanardo. Moment/history duality in Prior’s logics of branching-time. Synthese, 150(3):483–507, June 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-5519-4;
 http://link.springer.com/content/pdf/10.1007/s11229-005-5519-
4.pdf.

Anonymous:2006:CVc

[648] Anonymous. Contents of volume 150. Synthese, 150(3):509–510, June 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-8345-4;   http://link.springer.com/content/pdf/10.1007/
s11229-006-8345-4.pdf.

Anonymous:2006:AIc

[649] Anonymous. Author index. Synthese, 150(3):511, June 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-8346-3;
 http://link.springer.com/content/pdf/10.1007/s11229-006-8346-
3.pdf.

Anonymous:2006:IAc

[650] Anonymous. Instructions for authors. Synthese, 150(3):513–518, June 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-8347-2;   http://link.springer.com/content/pdf/10.1007/
s11229-006-8347-2.pdf.

Patterson:2006:TNR

[651] Douglas Eden Patterson. Tarski on the necessity reading of convention T. Synthese, 151(1):1–32, July 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-6230-6;   http://link.springer.com/
content/pdf/10.1007/s11229-004-6230-6.pdf.

Wallace:2006:DNC

[652] David Wallace. In defence of Naiveté: The conceptual status of Lagrangian quantum field theory. Synthese, 151(1):33–80, July 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-6248-9;   http://link.springer.com/content/pdf/10.1007/
s11229-004-6248-9.pdf.

Marton:2006:VVR

[653] Peter Marton. Verificationists versus realists: The battle over knowability. Synthese, 151(1):81–98, July 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-6269-4;   http://link.springer.com/
content/pdf/10.1007/s11229-004-6269-4.pdf.

Stevens:2006:RRP

[654] Graham Stevens. Russell’s repsychologising of the proposition. Synthese, 151(1):99–124, July 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-2246-1;   http://link.springer.com/content/
pdf/10.1007/s11229-004-2246-1.pdf.

Novaes:2006:RSO

[655] Catarina Dutilh Novaes. Roger Swyneshed’s obligationes: A logical game of inference recognition? Synthese, 151(1):125–153, July 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-2248-z;   http://link.springer.com/content/pdf/10.1007/
s11229-004-2248-z.pdf.

Anonymous:2006:Eb

[656] Anonymous. Erratum. Synthese, 151(1):155–156, July 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-0013-1;
 http://link.springer.com/content/pdf/10.1007/s11229-006-0013-
1.pdf.

Wrigley:2006:AP

[657] Anthony Wrigley. Abstracting propositions. Synthese, 151(2):157–176, July 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-2249-y;   http://link.springer.com/content/pdf/10.1007/
s11229-004-2249-y.pdf.

Bontly:2006:WEA

[658] Thomas D. Bontly. What is an empirical analysis of causation? Synthese, 151(2):177–200, July 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-2470-8;   http://link.springer.com/content/
pdf/10.1007/s11229-004-2470-8.pdf.

VanDitmarsch:2006:SMSa

[659] Hans Van Ditmarsch and Barteld Kooi. The secret of my success. Synthese, 151(2):201–232, July 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-3384-9;   http://link.springer.com/
content/pdf/10.1007/s11229-005-3384-9.pdf.

Mcarthur:2006:CCS

[660] Dan Mcarthur. Contra Cartwright: Structural realism, ontological pluralism and fundamentalism about laws. Synthese, 151(2):233–255, July 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-5406-4;   http://link.springer.com/content/pdf/10.1007/
s11229-004-5406-4.pdf.

Coffman:2006:DKC

[661] E. J. Coffman. Defending Klein on closure and skepticism. Synthese, 151(2):257–272, July 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-7324-x;   http://link.springer.com/content/
pdf/10.1007/s11229-004-7324-x.pdf.

Kallestrup:2006:PCS

[662] Jesper Kallestrup. Physicalism, conceivability and strong necessities. Synthese, 151(2):273–295, July 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-7325-9;   http://link.springer.com/
content/pdf/10.1007/s11229-004-7325-9.pdf.

Webster:2006:HZM

[663] William Robert Webster. Human zombies are metaphysically impossible. Synthese, 151(2):297–310, July 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-8006-4;   http://link.springer.com/
content/pdf/10.1007/s11229-004-8006-4.pdf.

Kistler:2006:NPR

[664] Max Kistler. New perspectives on reduction and emergence in physics, biology and psychology. Synthese, 151(3):311–312, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9014-3;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9014-3.pdf.

Moulines:2006:ORE

[665] C. Ulises Moulines. Ontology, reduction, emergence: A general frame. Synthese, 151(3):313–323, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9011-6;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9011-

6.pdf. See reply [666].

Ruphy:2006:ORR

[666] Stéphanie Ruphy. Ontology relativized: Reply to Moulines. Synthese, 151 (3):325–330, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9012-5;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9012-5.pdf. See [665] and reply [667].

Moulines:2006:RR

[667] C. Ulises Moulines. Reply to Ruphy. Synthese, 151(3):331–334, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9013-4;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9013-4.pdf. See [666].

Rueger:2006:FRE

[668] Alexander Rueger. Functional reduction and emergence in the physical sciences. Synthese, 151(3):335–346, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9027-y;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9027-

y.pdf. See reply [669].

Kistler:2006:REP

[669] Max Kistler. Reduction and emergence in the physical sciences: Reply to Rueger. Synthese, 151(3):347–354, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9028-x;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9028-

x.pdf. See [668].

Morange:2006:PGB

[670] Michel Morange. Post-genomics, between reduction and emergence. Synthese, 151(3):355–360, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9029-9;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9029-9.pdf.

Soto:2006:EDV

[671] Ana M. Soto and Carlos Sonnenschein. Emergentism by default: A view from the bench. Synthese, 151(3):361–376, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9030-3;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9030-3.pdf.

Schaffner:2006:RCC

[672] Kenneth F. Schaffner. Reduction: the Cheshire cat problem and a return to roots. Synthese, 151(3):377–402, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9031-2;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9031-

2.pdf. See reply [673].

Faucher:2006:WBS

[673] Luc Faucher. What’s behind a smile? The return of mechanism: Reply to Schaffner. Synthese, 151(3):403–409, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9032-1;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9032-1.pdf. See [672].

Bickle:2006:RMM

[674] John Bickle. Reducing mind to molecular pathways: explicating the reductionism implicit in current cellular and molecular neuroscience. Synthese, 151(3):411–434, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9015-2;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9015-
2.pdf.

deJong:2006:EPW

[675] Huib Looren de Jong. Explicating pluralism: Where the mind to molecule pathway gets off the track-reply to bickle. Synthese, 151(3):435–443, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9016-1;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9016-1.pdf.

Wimsatt:2006:RHM

[676] William C. Wimsatt. Reductionism and its heuristics: Making methodological reductionism honest. Synthese, 151(3):445–475, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9017-0;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9017-0.pdf. See reply [677].

Poirier:2006:FPE

[677] Pierre Poirier. Finding a place for elimination in inter-level reductionist activities: Reply to Wimsatt. Synthese, 151(3):477–483, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9018-z;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9018-z.pdf. See [676].

Stephan:2006:DRE

[678] Achim Stephan. The dual role of ‘emergence’ in the philosophy of mind and in cognitive science. Synthese, 151(3):485–498, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9019-y;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9019-y.pdf.

Harre:2006:RER

[679] Rom Harré. Resolving the emergence-reduction debate. Synthese, 151 (3):499–509, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9020-5;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9020-5.pdf.

Massin:2006:CCR

[680] Olivier Massin. Complementarity cannot resolve the emergencereduction debate: Reply to Harré. Synthese, 151(3):511–517, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9021-4;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9021-4.pdf.

Andler:2006:FSC

[681] Daniel Andler. Federalism in science- complementarity vs perspectivism: Reply to Harré. Synthese, 151(3):519–522, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9022-3;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9022-3.pdf.

Kirk:2006:PSI

[682] Robert Kirk. Physicalism and strict implication. Synthese, 151(3): 523–536, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9023-2;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9023-2.pdf.

Schroder:2006:PSI

[683] Jürgen Schröder. Physicalism and strict implication. Synthese, 151 (3):537–545, August 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9024-1;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9024-1.pdf.

Kim:2006:ECI

[684] Jaegwon Kim. Emergence: Core ideas and issues. Synthese, 151(3):547– 559, August 2006. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-006-9025-0;   https://link.springer.com/content/pdf/
10.1007/s11229-006-9025-0.pdf. See reply [685].

Marras:2006:ERR

[685] Ausonio Marras. Emergence and reduction: Reply to Kim. Synthese, 151(3):561–569, August 2006. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9026-z;   https://link.springer.
com/content/pdf/10.1007/s11229-006-9026-z.pdf. See [684].

Winsberg:2006:MSV

[686] Eric Winsberg. Models of success versus the success of models: Reliability without truth. Synthese, 152(1):1–19, September 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
004-5404-6;   http://link.springer.com/content/pdf/10.1007/
s11229-004-5404-6.pdf.

Schnieder:2006:TMT

[687] Benjamin Schnieder. Truth-making without truth-makers. Synthese, 152 (1):21–46, September 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-004-7905-8;   http://link.springer.com/content/
pdf/10.1007/s11229-004-7905-8.pdf.

Brogaard:2006:GEA

[688] B. Brogaard. The ‘Gray’s Elegy’ argument, and the prospects for the theory of denoting concepts. Synthese, 152(1):47–79, September 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-0547-7;   http://link.springer.com/content/pdf/10.1007/
s11229-005-0547-7.pdf.

Otte:2006:CEP

[689] R. Otte. Counterfactuals and epistemic probability. Synthese, 152(1): 81–93, September 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-1332-3;   http://link.springer.com/content/
pdf/10.1007/s11229-005-1332-3.pdf.

Briscoe:2006:IEI

[690] Robert Eamon Briscoe. Individualism, externalism and idiolectical meaning. Synthese, 152(1):95–128, September 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-1333-2;
 http://link.springer.com/content/pdf/10.1007/s11229-005-1333-
2.pdf.

Arlo-Costa:2006:CDT

[691] Horacio Arló-Costa and Isaac Levi. Contraction: On the decisiontheoretical origins of minimal change and entrenchment. Synthese, 152(1):129–154, September 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-0351-4;   http://link.springer.com/
content/pdf/10.1007/s11229-005-0351-4.pdf.

Lapointe:2006:E

[692] Sandra Lapointe. Erratum. Synthese, 152(1):155, September 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

006-8465-x;   http://link.springer.com/content/pdf/10.1007/
s11229-006-8465-x.pdf.

Atkinson:2006:CRP

[693] Katie Atkinson, Trevor Bench-Capon, and Peter McBurney. Computational representation of practical argument. Synthese, 152(2):157– 206, September 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-3488-2;   http://link.springer.com/content/
pdf/10.1007/s11229-005-3488-2.pdf.

Hahn:2006:BAI

[694] Ulrike Hahn and Mike Oaksford. A Bayesian approach to informal argument fallacies. Synthese, 152(2):207–236, September 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-5233-2;   http://link.springer.com/content/pdf/10.1007/
s11229-005-5233-2.pdf.

Walton:2006:EDM

[695] Douglas Walton. Epistemic and dialectical models of begging the question. Synthese, 152(2):237–284, September 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-3984-4;
 http://link.springer.com/content/pdf/10.1007/s11229-005-3984-
4.pdf.

Pigozzi:2006:BMD

[696] Gabriella Pigozzi. Belief merging and the discursive dilemma: an argument-based account to paradoxes of judgment aggregation. Synthese, 152(2):285–298, September 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9063-7;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9063-7.pdf.

Bardon:2006:I

[697] Adrian Bardon. Introduction. Synthese, 152(3):299–300, October 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9006-3;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9006-3.pdf.

Garrett:2006:HNT

[698] Don Garrett. Hume’s naturalistic theory of representation. Synthese, 152 (3):301–319, October 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9007-2;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9007-2.pdf.

Loeb:2006:PES

[699] Louis E. Loeb. Psychology, epistemology, and skepticism in Hume’s argument about induction. Synthese, 152(3):321–338, October 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9008-1;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9008-1.pdf.

Stroud:2006:CHN

[700] Barry Stroud. The constraints of Hume’s naturalism. Synthese, 152 (3):339–351, October 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9009-0;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9009-0.pdf.

Radcliffe:2006:MIM

[701] Elizabeth S. Radcliffe. Moral internalism and moral cognitivism in Hume’s metaethics. Synthese, 152(3):353–370, October 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9003-6;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9003-
6.pdf.

Postema:2006:WAH

[702] Gerald J. Postema. Whence avidity? Hume’s psychology and the origins of justice. Synthese, 152(3):371–391, October 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9004-5;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9004-
5.pdf.

Mcintyre:2006:SMP

[703] Jane L. Mcintyre. Strength of mind: Prospects and problems for a Humean account. Synthese, 152(3):393–401, October 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9005-4;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9005-
4.pdf.

Dresner:2006:MTA

[704] Eli Dresner. A measurement theoretic account of propositions. Synthese, 153(1):1–22, November 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-004-3383-2;   http://link.springer.com/
content/pdf/10.1007/s11229-004-3383-2.pdf.

Voltolini:2006:FBI

[705] Alberto Voltolini. Fiction as a base of interpretation contexts. Synthese, 153(1):23–47, November 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-0001-5;   http://link.springer.com/
content/pdf/10.1007/s11229-006-0001-5.pdf.

Krasner:2006:SI

[706] Daniel Asher Krasner. Smith on indexicals. Synthese, 153(1):49– 67, November 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-2724-0;   http://link.springer.com/content/
pdf/10.1007/s11229-005-2724-0.pdf.

Collins:2006:PLN

[707] J. Collins. Proxytypes and linguistic nativism. Synthese, 153(1):69– 104, November 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-4063-6;   http://link.springer.com/content/
pdf/10.1007/s11229-005-4063-6.pdf.

Avigad:2006:MMP

[708] Jeremy Avigad. Mathematical method and proof. Synthese, 153(1): 105–159, November 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-4064-5;   http://link.springer.com/content/
pdf/10.1007/s11229-005-4064-5.pdf.

Luper:2006:RRS

[709] S. Luper. Restorative rigging and the safe indication account. Synthese, 153(1):161–170, November 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-6399-3;   http://link.springer.com/
content/pdf/10.1007/s11229-005-6399-3.pdf.

Manor:2006:SH

[710] Ruth Manor. Solving the heap. Synthese, 153(2):171–186, November 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-2535-3;   http://link.springer.com/content/pdf/10.1007/
s11229-005-2535-3.pdf.

Atkinson:2006:PAW

[711] David Atkinson and Jeanne Peijnenburg. Probability all the way up. Synthese, 153(2):187–197, November 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-2723-1;   http://link.springer.com/
content/pdf/10.1007/s11229-005-2723-1.pdf.

Moltmann:2006:UAP

[712] Friederike Moltmann. Unbound anaphoric pronouns: E-type, dynamic, and structured-propositions approaches. Synthese, 153(2):199– 260, November 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-5469-x;   http://link.springer.com/content/
pdf/10.1007/s11229-005-5469-x.pdf.

Gillett:2006:SAE

[713] Carl Gillett. Samuel Alexander’s emergentism: Or, higher causation for physicalists. Synthese, 153(2):261–296, November 2006. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-5470-4;
 http://link.springer.com/content/pdf/10.1007/s11229-005-5470-
4.pdf.

Holton:2006:PFH

[714] Gerald Holton. Philipp Frank at Harvard University: His work and his influence. Synthese, 153(2):297–311, November 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-5471-3;
 http://link.springer.com/content/pdf/10.1007/s11229-005-5471-
3.pdf.

Ellis:2006:MON

[715] Stephen Ellis. Multiple objectives: A neglected problem in the theory of human action. Synthese, 153(2):313–338, November 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-5751-y;   http://link.springer.com/content/pdf/10.1007/
s11229-005-5751-y.pdf.

VanDitmarsch:2006:SMSb

[716] Hans Van Ditmarsch and Barteld Kooi. The secret of my success. Synthese, 153(2):339, November 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-8493-6;   http://link.springer.com/
content/pdf/10.1007/s11229-006-8493-6.pdf.

Bickle:2006:EI

[717] John Bickle. Editor’s introduction. Synthese, 153(3):341–342, December 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9094-0;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9094-0.pdf.

Piccinini:2006:CEN

[718] Gualtiero Piccinini. Computational explanation in neuroscience. Synthese, 153(3):343–353, December 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9096-y;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9096-y.pdf.

Craver:2006:WMM

[719] Carl F. Craver. When mechanistic models explain. Synthese, 153(3): 355–376, December 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9097-x;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9097-x.pdf.

Egan:2006:DCN

[720] Frances Egan and Robert J. Matthews. Doing cognitive neuroscience: a third way. Synthese, 153(3):377–391, December 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9098-9;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9098-
9.pdf.

Shagrir:2006:WWV

[721] Oron Shagrir. Why we view the brain as a computer. Synthese, 153(3): 393–416, December 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9099-8;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9099-8.pdf.

Grush:2006:HHB

[722] Rick Grush. How to, and how not to, bridge computational cognitive neuroscience and Husserlian phenomenology of time consciousness. Synthese, 153(3):417–450, December 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9100-6;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9100-
6.pdf.

Bentwich:2006:DPI

[723] Jonathan Bentwich. The duality principle: irreducibility of sub-threshold psychophysical computation to neuronal brain activation. Synthese, 153 (3):451–455, December 2006. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9101-5;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9101-5.pdf.

Davelaar:2006:CWL

[724] G. G. Davelaar and L. Abelmann. Comment on Wang, Liu, and Wang (2003). Synthese, 153(3):457–458, December 2006. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9095-z;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9095-
z.pdf.

vanBenthem:2007:E

[725] Johan van Benthem, Vincent F. Hendricks, and John Symons. Editorial. Synthese, 154(1):1–3, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-0014-0;   http://link.springer.com/
content/pdf/10.1007/s11229-006-0014-0.pdf.

Bartha:2007:TSI

[726] P. Bartha. Taking stock of infinite value: Pascal’s wager and relative utilities. Synthese, 154(1):5–52, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-8006-z;
 http://link.springer.com/content/pdf/10.1007/s11229-005-8006-
z.pdf.

Black:2007:DS

[727] Tim Black and Peter Murphy. In defense of sensitivity. Synthese, 154 (1):53–71, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-8487-9;   http://link.springer.com/content/
pdf/10.1007/s11229-005-8487-9.pdf.

Moretti:2007:PMC

[728] Luca Moretti and Ken Akiba. Probabilistic measures of coherence and the problem of belief individuation. Synthese, 154(1):73–95, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-0193-0;   http://link.springer.com/content/pdf/10.1007/
s11229-005-0193-0.pdf.

Piccinini:2007:CCT

[729] Gualtiero Piccinini. Computationalism, the Church–Turing thesis, and the Church–Turing fallacy. Synthese, 154(1):97–120, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-0194-z;   http://link.springer.com/content/pdf/10.1007/
s11229-005-0194-z.pdf.

Hanks:2007:HWD

[730] Peter W. Hanks. How Wittgenstein defeated Russell’s multiple relation theory of judgment. Synthese, 154(1):121–146, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-0195-y;   http://link.springer.com/content/pdf/10.1007/
s11229-005-0195-y.pdf.

Welch:2007:VIM

[731] J. R. Welch. Vagueness and inductive molding. Synthese, 154(1): 147–172, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-0196-x;   http://link.springer.com/content/
pdf/10.1007/s11229-005-0196-x.pdf.

Placek:2007:CHP

[732] Tomasz Placek and Thomas Müller. Counterfactuals and historical possibility. Synthese, 154(2):173–197, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-0617-x;
 http://link.springer.com/content/pdf/10.1007/s11229-005-0617-
x.pdf.

Bellotti:2007:FSS

[733] Luca Bellotti. Formalization, syntax and the standard model of arithmetic. Synthese, 154(2):199–229, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-0900-x;
 http://link.springer.com/content/pdf/10.1007/s11229-005-0900-
x.pdf.

Delancey:2007:MNM

[734] Craig Stephen Delancey. Meaning naturalism, meaning irrealism, and the work of language. Synthese, 154(2):231–257, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-1006-1;   http://link.springer.com/content/pdf/10.1007/
s11229-005-1006-1.pdf.

Marcum:2007:ESJ

[735] James A. Marcum. Experimental series and the justification of Temin’s DNA provirus hypothesis. Synthese, 154(2):259–292, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-1793-4;   http://link.springer.com/content/pdf/10.1007/
s11229-005-1793-4.pdf.

Gemes:2007:VC

[736] Ken Gemes. Verisimilitude and content. Synthese, 154(2):293–306, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-2565-x;   http://link.springer.com/content/pdf/10.1007/
s11229-005-2565-x.pdf.

VanLaar:2007:OSA

[737] Jan Albert Van Laar. One-sided arguments. Synthese, 154(2):307– 327, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-2882-0;   http://link.springer.com/content/
pdf/10.1007/s11229-005-2882-0.pdf.

Schupbach:2007:MSR

[738] Jonah N. Schupbach. Must the scientific realist be a rationalist? Synthese, 154(2):329–334, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-3490-8;   http://link.springer.com/
content/pdf/10.1007/s11229-005-3490-8.pdf.

Caplan:2007:NDM

[739] Ben Caplan. A new defence of the modal existence requirement. Synthese, 154(2):335–343, January 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-3491-7;   http://link.springer.com/
content/pdf/10.1007/s11229-005-3491-7.pdf.

Beziau:2007:NTF

[740] Jean-Yves Béziau and Décio Krause. New trends in the foundations of science. Synthese, 154(3):345–347, February 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9140-y;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9140-
y.pdf.

deBarros:2007:RET

[741] J. Acacio de Barros, J. P. R. F. de Mendonça, and N. Pinto-Neto. Realism in energy transition processes: an example from Bohmian quantum mechanics. Synthese, 154(3):349–370, February 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9123-z;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9123-z.pdf.

Beziau:2007:SPI

[742] Jean-Yves Béziau. Sentence, proposition and identity. Synthese, 154(3): 371–382, February 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9124-y;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9124-y.pdf.

Bueno:2007:QTP

[743] Otávio Bueno and Newton C. A. da Costa. Quasi-truth, paraconsistency, and the foundations of science. Synthese, 154(3):383–399, February 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9125-x;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9125-x.pdf.

Doria:2007:IVF

[744] Francisco Antonio Doria. Informal versus formal mathematics. Synthese, 154(3):401–415, February 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9126-9;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9126-9.pdf.

Krause:2007:QSP

[745] Décio Krause and Steven French. Quantum sortal predicates. Synthese, 154(3):417–430, February 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9127-8;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9127-8.pdf.

Buchsbaum:2007:LER

[746] Arthur Buchsbaum, Tarcisio Pequeno, and Marcelino Pequeno. A logical expression of reasoning. Synthese, 154(3):431–466, February 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9128-7;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9128-7.pdf.

SantAnna:2007:SPC

[747] Adonai S. Sant’Anna and Gabriel Guerrer. Some problems concerning language and physics. Synthese, 154(3):467–484, February 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9129-6;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9129-6.pdf.

Suppes:2007:SCP

[748] Patrick Suppes. Statistical concepts in philosophy of science. Synthese, 154(3):485–496, February 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9122-0;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9122-0.pdf.

Hudson:2007:WRI

[749] Robert G. Hudson. What’s really at issue with novel predictions? Synthese, 155(1):1–20, March 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-6267-1;   http://link.springer.com/
content/pdf/10.1007/s11229-005-6267-1.pdf.

Gerbrandy:2007:SED

[750] J. Gerbrandy. The surprise examination in dynamic epistemic logic. Synthese, 155(1):21–33, March 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-2211-7;   http://link.springer.com/
content/pdf/10.1007/s11229-005-2211-7.pdf.

Levin:2007:BHK

[751] Michael E. Levin. Bundling Hume with Kripkenstein. Synthese, 155 (1):35–64, March 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-2420-0;   http://link.springer.com/content/
pdf/10.1007/s11229-005-2420-0.pdf.

Aizawa:2007:BMC

[752] K. Aizawa. The biochemistry of memory consolidation: A model system for the philosophy of mind. Synthese, 155(1):65–98, March 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-2566-9;   http://link.springer.com/content/pdf/10.1007/
s11229-005-2566-9.pdf.

Montminy:2007:ECS

[753] Martin Montminy. Epistemic contextualism and the semantics– pragmatics distinction. Synthese, 155(1):99–125, March 2007. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-2881-1;
 http://link.springer.com/content/pdf/10.1007/s11229-005-2881-
1.pdf.

Levy:2007:DR

[754] Neil Levy. Doxastic responsibility. Synthese, 155(1):127–155, March 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-3983-5;   http://link.springer.com/content/pdf/10.1007/
s11229-005-3983-5.pdf.

Mcdaniel:2007:DDS

[755] K. Mcdaniel. Distance and discrete space. Synthese, 155(1):157– 162, March 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-5034-7;   http://link.springer.com/content/
pdf/10.1007/s11229-005-5034-7.pdf.

Paglieri:2007:CMR

[756] Fabio Paglieri. Changing minds: the role of beliefs in cognitive dynamics. Synthese, 155(2):163–166, March 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9146-5;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9146-
5.pdf.

Leitgeb:2007:DDL

[757] Hannes Leitgeb and Krister Segerberg. Dynamic doxastic logic: why, how, and where to? Synthese, 155(2):167–190, March 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9143-8;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9143-8.pdf.

vanDitmarsch:2007:MBA

[758] Hans van Ditmarsch and Willem Labuschagne. My beliefs about your beliefs: a case study in theory of mind and epistemic logic. Synthese, 155 (2):191–209, March 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9144-7;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9144-7.pdf.

Jansen:2007:RTB

[759] Brenda R. J. Jansen, Maartje E. J. Raijmakers, and Ingmar Visser. Rule transition on the balance scale task: a case study in belief change. Synthese, 155(2):211–236, March 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9142-9;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9142-
9.pdf.

Castelfranchi:2007:RBG

[760] Cristiano Castelfranchi and Fabio Paglieri. The role of beliefs in goal dynamics: prolegomena to a constructive theory of intentions. Synthese, 155(2):237–263, March 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9156-3;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9156-3.pdf.

vanderHoek:2007:TTI

[761] Wiebe van der Hoek, Wojciech Jamroga, and Michael Wooldridge. Towards a theory of intention revision. Synthese, 155(2):265–290, March 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9145-6;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9145-6.pdf.

Mclntyre:2007:PCT

[762] Lee Mclntyre. The philosophy of chemistry: ten years later. Synthese, 155(3):291–292, April 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9114-0;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9114-0.pdf.

Bhushan:2007:WCP

[763] Nalini Bhushan. What is a chemical property? Synthese, 155(3):293–305, April 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9115-z;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9115-z.pdf.

Ramsey:2007:CCM

[764] Jeffry L. Ramsey. Calibrating and constructing models of protein folding. Synthese, 155(3):307–320, April 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9113-1;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9113-
1.pdf.

Hoffmann:2007:WMP

[765] Roald Hoffmann. What might philosophy of science look like if chemists built it? Synthese, 155(3):321–336, April 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9118-9;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9118-
9.pdf.

McIntyre:2007:ERC

[766] Lee McIntyre. Emergence and reduction in chemistry: ontological or epistemological concepts? Synthese, 155(3):337–343, April 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9111-3;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9111-3.pdf.

vanderHoek:2007:SLA

[767] Wiebe van der Hoek, Mark Roberts, and Michael Wooldridge. Social laws in alternating time: effectiveness, feasibility, and synthesis. Synthese, 156(1):1–19, May 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9072-6;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9072-6.pdf.

Laraudogoitia:2007:AIM

[768] J. P. Laraudogoitia. Avoiding infinite masses. Synthese, 156(1):21–31, May 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-2212-6;   http://link.springer.com/content/pdf/10.1007/
s11229-005-2212-6.pdf.

Pruss:2007:CDL

[769] Alexander R. Pruss. Conjunctions, disjunctions and Lewisian semantics for counterfactuals. Synthese, 156(1):33–52, May 2007. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-3487-3;
 http://link.springer.com/content/pdf/10.1007/s11229-005-3487-
3.pdf.

Steel:2007:BCT

[770] Daniel Steel. Bayesian confirmation theory and the likelihood principle. Synthese, 156(1):53–77, May 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-3492-6;   http://link.springer.com/
content/pdf/10.1007/s11229-005-3492-6.pdf.

Daly:2007:ART

[771] Chris John Daly. Acquaintance and de re thought. Synthese, 156 (1):79–96, May 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-3985-3;   http://link.springer.com/content/
pdf/10.1007/s11229-005-3985-3.pdf.

Gross:2007:TSV

[772] Steven Gross. Trivalent semantics and the vaguely vague. Synthese, 156(1):97–117, May 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-4542-9;   http://link.springer.com/content/
pdf/10.1007/s11229-005-4542-9.pdf.

Huemer:2007:EP

[773] Michael Huemer. Epistemic possibility. Synthese, 156(1):119–142, May 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
005-4782-8;   http://link.springer.com/content/pdf/10.1007/
s11229-005-4782-8.pdf.

Cameron:2007:LRM

[774] Ross P. Cameron. Lewisian realism: Methodology, epistemology, and circularity. Synthese, 156(1):143–159, May 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-005-2003-0;
 http://link.springer.com/content/pdf/10.1007/s11229-005-2003-
0.pdf.

Weiskopf:2007:CNC

[775] Daniel A. Weiskopf. Compound nominals, context, and compositionality. Synthese, 156(1):161–204, May 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-005-3489-1;   http://link.springer.com/
content/pdf/10.1007/s11229-005-3489-1.pdf.

Kulvicki:2007:WWL

[776] John Kulvicki. What is what it’s like? Introducing perceptual modes of presentation. Synthese, 156(2):205–229, May 2007. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-0003-3;
 http://link.springer.com/content/pdf/10.1007/s11229-006-0003-
3.pdf.

Johnston:2007:UTF

[777] Colin Johnston. The unity of a tractarian fact. Synthese, 156(2):231–251, May 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-0002-4;   http://link.springer.com/content/pdf/10.1007/
s11229-006-0002-4.pdf.

Johnson:2007:TAU

[778] Kent Johnson. Tacit and accessible understanding of language. Synthese, 156(2):253–279, May 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-0006-0;   http://link.springer.com/content/
pdf/10.1007/s11229-006-0006-0.pdf.

Hartimo:2007:TCH

[779] Mirja Helena Hartimo. Towards completeness: Husserl on theories of manifolds 1890–1901. Synthese, 156(2):281–310, May 2007. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-0008-y;
 http://link.springer.com/content/pdf/10.1007/s11229-006-0008-
y.pdf.

DeRosa:2007:TAC

[780] Raffaella De Rosa. A teleological account of Cartesian sensations? Synthese, 156(2):311–336, May 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-0010-4;   http://link.springer.com/
content/pdf/10.1007/s11229-006-0010-4.pdf.

Shapiro:2007:OM

[781] Stewart Shapiro. The objectivity of mathematics. Synthese, 156(2): 337–381, May 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-005-5298-y;   http://link.springer.com/content/
pdf/10.1007/s11229-005-5298-y.pdf.

Dodd:2007:NTT

[782] Julian Dodd. Negative truths and truthmaker principles. Synthese, 156(2):383–401, May 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-0007-z;   http://link.springer.com/content/
pdf/10.1007/s11229-006-0007-z.pdf.

Bovens:2007:SIS

[783] Luc Bovens and Stephan Hartmann. Special issue of synthese on Bayesian epistemology. Synthese, 156(3):403, June 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9130-0;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9130-
0.pdf.

Douven:2007:MC

[784] Igor Douven and Wouter Meijs. Measuring coherence. Synthese, 156 (3):405–425, June 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9131-z;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9131-z.pdf.

Dieks:2007:RAF

[785] Dennis Dieks. Reasoning about the future: Doom and beauty. Synthese, 156(3):427–439, June 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9132-y;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9132-y.pdf.

Suppes:2007:WDB

[786] Patrick Suppes. Where do Bayesian priors come from? Synthese, 156 (3):441–471, June 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9133-x;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9133-x.pdf.

Fitelson:2007:LBR

[787] Branden Fitelson. Likelihoodism, Bayesianism, and relational confirmation. Synthese, 156(3):473–489, June 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-006-9134-9;   http://link.springer.
com/content/pdf/10.1007/s11229-006-9134-9.pdf.

Howson:2007:LN

[788] Colin Howson. Logic with numbers. Synthese, 156(3):491–512, June 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9135-8;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9135-8.pdf.

Bradley:2007:KBD

[789] Richard Bradley. The kinematics of belief and desire. Synthese, 156 (3):513–535, June 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9136-7;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9136-7.pdf.

Joyce:2007:NPR

[790] James M. Joyce. Are Newcomb problems really decisions? Synthese, 156(3):537–562, June 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9137-6;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9137-6.pdf.

Hajek:2007:RCP

[791] Alan Hájek. The reference class problem is your problem too. Synthese, 156(3):563–585, June 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9138-5;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9138-5.pdf.

Albert:2007:PTD

[792] M. Albert. The propensity theory: a decision-theoretic restatement. Synthese, 156(3):587–603, June 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9139-4;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9139-4.pdf.

McGrath:2007:MEC

[793] Matthew McGrath. Memory and epistemic conservatism. Synthese, 157(1):1–24, July 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-0011-3;   http://link.springer.com/content/
pdf/10.1007/s11229-006-0011-3.pdf.

Jenkins:2007:ER

[794] C. S. Jenkins. Entitlement and rationality. Synthese, 157(1):25–45, July 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-0012-2;   http://link.springer.com/content/pdf/10.1007/
s11229-006-0012-2.pdf.

Byrd:2007:PMP

[795] Jeremy Allen Byrd. The perfect murder: A philosophical whodunit. Synthese, 157(1):47–58, July 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-0004-2;   http://link.springer.com/
content/pdf/10.1007/s11229-006-0004-2.pdf.

Bostrom:2007:SBS

[796] Nick Bostrom. Sleeping beauty and self-location: A hybrid model. Synthese, 157(1):59–78, July 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9010-7;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9010-7.pdf.

Batitsky:2007:WGT

[797] Vadim Batitsky and Zoltan Domotor. When good theories make bad predictions. Synthese, 157(1):79–103, July 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9033-0;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9033-
0.pdf.

Dieveney:2007:DIA

[798] Patrick S. Dieveney. Dispensability in the indispensability argument. Synthese, 157(1):105–128, July 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9035-y;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9035-y.pdf.

Smith:2007:SES

[799] Deborah C. Smith. Superassertibility and the equivalence schema: A dilemma for Wright’s antirealist. Synthese, 157(1):129–139, July 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9037-9;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9037-9.pdf.

Bardsley:2007:CIC

[800] Nicholas Bardsley. On collective intentions: collective action in economics and philosophy. Synthese, 157(2):141–159, July 2007. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9034-z;

 http://link.springer.com/content/pdf/10.1007/s11229-006-9034-
z.pdf.

Gemes:2007:ISB

[801] Ken Gemes. Irrelevance: Strengthening the Bayesian requirements. Synthese, 157(2):161–166, July 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-0009-x;   http://link.springer.com/
content/pdf/10.1007/s11229-006-0009-x.pdf.

Vanderschraaf:2007:CR

[802] Peter Vanderschraaf. Covenants and reputations. Synthese, 157(2): 167–195, July 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9147-4;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9147-4.pdf.

Walton:2007:EPR

[803] Douglas Walton. Evaluating practical reasoning. Synthese, 157(2): 197–240, July 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9157-x;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9157-x.pdf.

Huffer:2007:AOT

[804] Beth Huffer. Actions and outcomes: two aspects of agency. Synthese, 157(2):241–265, July 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9107-z;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9107-z.pdf.

Olsson:2007:GEI

[805] Erik J. Olsson. Guest editor’s introduction. Synthese, 157(3):267– 274, August 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9061-9;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9061-9.pdf.

Glass:2007:CMI

[806] David H. Glass. Coherence measures and inference to the best explanation. Synthese, 157(3):275–296, August 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9055-7;

 http://link.springer.com/content/pdf/10.1007/s11229-006-9055-
7.pdf.

Olsson:2007:RCM

[807] Erik J. Olsson and Stefan Schubert. Reliability conducive measures of coherence. Synthese, 157(3):297–308, August 2007. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9056-6;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9056-
6.pdf.

Moretti:2007:WWC

[808] Luca Moretti. Ways in which coherence is confirmation conducive. Synthese, 157(3):309–319, August 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9057-5;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9057-5.pdf.

Angere:2007:DNC

[809] Staffan Angere. The defeasible nature of coherentist justification. Synthese, 157(3):321–335, August 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9058-4;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9058-4.pdf.

Huemer:2007:WBC

[810] Michael Huemer. Weak Bayesian coherentism. Synthese, 157(3): 337–346, August 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9059-3;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9059-3.pdf.

Meijs:2007:AIC

[811] Wouter Meijs and Igor Douven. On the alleged impossibility of coherence. Synthese, 157(3):347–360, August 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9060-x;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9060-
x.pdf.

Shogenji:2007:WDC

[812] Tomoji Shogenji. Why does coherence appear truth-conducive? Synthese, 157(3):361–372, August 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9062-8;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9062-8.pdf.

Landry:2007:SSN

[813] Elaine Landry. Shared structure need not be shared set-structure. Synthese, 158(1):1–17, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9047-7;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9047-7.pdf.

Graham:2007:TDS

[814] Peter J. Graham. The theoretical diagnosis of skepticism. Synthese, 158 (1):19–39, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9048-6;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9048-6.pdf.

Horsten:2007:UPA

[815] Leon Horsten and Philip Welch. The undecidability of propositional adaptive logic. Synthese, 158(1):41–60, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9049-5;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9049-
5.pdf.

Wray:2007:KRR

[816] K. Brad Wray. Kuhnian revolutions revisited. Synthese, 158(1):61– 73, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9050-z;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9050-z.pdf.

Zwart:2007:ITV

[817] Sjoerd D. Zwart and Maarten Franssen. An impossibility theorem for verisimilitude. Synthese, 158(1):75–92, September 2007. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9051-y;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9051-
y.pdf.

Adam:2007:TNS

[818] Matthias Adam. Two notions of scientific justification. Synthese, 158(1): 93–108, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9052-x;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9052-x.pdf.

Keijzer:2007:ECM

[819] Fred Keijzer and Maurice Schouten. Embedded cognition and mental causation: Setting empirical bounds on metaphysics. Synthese, 158(1): 109–125, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9053-9;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9053-9.pdf.

Schlenker:2007:HES

[820] Philippe Schlenker. How to eliminate self-reference: a précis. Synthese, 158(1):127–138, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9054-8;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9054-8.pdf.

Levy:2007:BMH

[821] Ken Levy. Baumann on the Monty Hall problem and single-case probabilities. Synthese, 158(1):139–151, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9065-5;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9065-

5.pdf. See comments [927].

Anonymous:2007:IUH

[822] Anonymous. International union of history and philosophy of science, division of logic, methodology and philosophy of science. Synthese, 158(1): 153–163, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9183-8;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9183-8.pdf.

Plaza:2007:LPC

[823] Jan Plaza. Logics of public communications. Synthese, 158(2):165– 179, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9168-7;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9168-7.pdf.

vanDitmarsch:2007:CLP

[824] Hans P. van Ditmarsch. Comments to ‘logics of public communications’. Synthese, 158(2):181–187, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9167-8;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9167-
8.pdf.

Steele:2007:DIB

[825] Katie Steele. Distinguishing indeterminate belief from “risk-averse” preferences. Synthese, 158(2):189–205, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9119-8;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9119-
8.pdf.

Hughes:2007:SME

[826] Jesse Hughes, Peter Kroes, and Sjoerd Zwart. A semantics for meansend relations. Synthese, 158(2):207–231, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9036-x;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9036-
x.pdf.

Pauly:2007:ACJ

[827] Marc Pauly. Axiomatizing collective judgment sets in a minimal logical language. Synthese, 158(2):233–250, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9079-z;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9079-
z.pdf.

Perea:2007:OPD

[828] Andrés Perea. A one-person doxastic characterization of Nash strategies. Synthese, 158(2):251–271, September 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9217-2;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9217-
2.pdf.

Pritchard:2007:I

[829] Duncan Pritchard. Introduction. Synthese, 158(3):273–275, October 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9038-8;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9038-8.pdf.

Pritchard:2007:ALE

[830] Duncan Pritchard. Anti-luck epistemology. Synthese, 158(3):277– 297, October 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9039-7;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9039-7.pdf.

Greco:2007:WAP

[831] John Greco. Worries about Pritchard’s safety. Synthese, 158(3): 299–302, October 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9040-1;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9040-1.pdf.

Hiller:2007:SEL

[832] Avram Hiller and Ram Neta. Safety and epistemic luck. Synthese, 158 (3):303–313, October 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9041-0;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9041-0.pdf.

Goldberg:2007:HLC

[833] Sanford Goldberg. How lucky can you get? Synthese, 158(3):315– 327, October 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9042-z;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9042-z.pdf.

Riggs:2007:WET

[834] Wayne Riggs. Why epistemologists are so down on their luck. Synthese, 158(3):329–344, October 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9043-y;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9043-y.pdf.

Lackey:2007:WWD

[835] Jennifer Lackey. Why we don’t deserve credit for everything we know. Synthese, 158(3):345–361, October 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9044-x;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9044-
x.pdf.

Axtell:2007:TSA

[836] Guy Axtell. Two for the show: Anti-luck and virtue epistemologies in consonance. Synthese, 158(3):363–383, October 2007. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9045-9;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9045-
9.pdf.

Coffman:2007:TAL

[837] E. J. Coffman. Thinking about luck. Synthese, 158(3):385–398, October 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9046-8;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9046-8.pdf.

Leeds:2007:CTS

[838] Stephen Leeds. Correspondence truth and scientific realism. Synthese, 159(1):1–21, November 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9064-6;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9064-6.pdf.

Awodey:2007:CDG

[839] S. Awodey and A. W. Carus. Carnap’s dream: Gödel, Wittgenstein, and logical, syntax. Synthese, 159(1):23–45, November 2007. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9066-4;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9066-
4.pdf.

Ellis:2007:CEP

[840] Jonathan Ellis. Content externalism and phenomenal character: a new worry about privileged access. Synthese, 159(1):47–60, November 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9067-3;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9067-3.pdf.

vanderSchaar:2007:ACM

[841] Maria van der Schaar. The assertion-candidate and the meaning of mood. Synthese, 159(1):61–82, November 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9068-2;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9068-
2.pdf.

Pollard:2007:MDT

[842] Stephen Pollard. Mathematical determinacy and the transferability of aboutness. Synthese, 159(1):83–98, November 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9069-1;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9069-
1.pdf.

Jager:2007:LSP

[843] Gerhard Jäger and Robert van Rooij. Language structure: psychological and social constraints. Synthese, 159(1):99–130, November 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9073-5;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9073-5.pdf.

Guarini:2007:CNB

[844] Marcello Guarini. Critical notice: BonJour and Sosa on epistemic justification. Synthese, 159(1):131–148, November 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9229-y;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9229-

y.pdf. See [8].

Bogdan:2007:I

[845] Radu Bogdan. Introduction. Synthese, 159(2):149–150, November 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9210-9;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9210-9.pdf.

Gordon:2007:ARP

[846] Robert M. Gordon. Ascent routines for propositional attitudes. Synthese, 159(2):151–165, November 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9202-9;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9202-9.pdf.

Russell:2007:CCK

[847] James Russell. Controlling core knowledge: conditions for the ascription of intentional states to self and others by children. Synthese, 159(2): 167–196, November 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9203-8;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9203-8.pdf.

Carruthers:2007:ICW

[848] Peter Carruthers. The illusion of conscious will. Synthese, 159(2): 197–213, November 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9204-7;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9204-7.pdf.

Nichols:2007:IIT

[849] Shaun Nichols. Imagination and immortality: thinking of me. Synthese, 159(2):215–233, November 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9205-6;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9205-6.pdf.

Bogdan:2007:ILD

[850] Radu J. Bogdan. Inside loops: developmental premises of selfascriptions. Synthese, 159(2):235–251, November 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9206-5;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9206-
5.pdf.

Perner:2007:IR

[851] Josef Perner, Daniela Kloo, and Elisabeth Stöttinger. Introspection & remembering. Synthese, 159(2):253–270, November 2007. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9207-4;

 http://link.springer.com/content/pdf/10.1007/s11229-007-9207-
4.pdf.

Proust:2007:MMS

[852] Joëlle Proust. Metacognition and metarepresentation: is a self-directed theory of mind a precondition for metacognition? Synthese, 159(2): 271–295, November 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9208-3;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9208-3.pdf.

Olson:2007:SAI

[853] David R. Olson. Self-ascription of intention: responsibility, obligation and self-control. Synthese, 159(2):297–314, November 2007. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9209-2;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9209-
2.pdf.

Hohwy:2007:FIM

[854] Jakob Hohwy. Functional integration and the mind. Synthese, 159(3): 315–328, December 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9240-3;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9240-3.pdf.

Anderson:2007:MRE

[855] Michael L. Anderson. Massive redeployment, exaptation, and the functional integration of cognitive operations. Synthese, 159(3):329– 345, December 2007. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9233-2;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9233-2.pdf.

McGeer:2007:WNM

[856] Victoria McGeer. Why neuroscience matters to cognitive neuropsychology. Synthese, 159(3):347–371, December 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9234-1;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9234-
1.pdf.

Eliasmith:2007:HBB

[857] Chris Eliasmith. How to build a brain: from function to implementation. Synthese, 159(3):373–388, December 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9235-0;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9235-
0.pdf.

Grush:2007:STV

[858] Rick Grush. Skill theory v2.0: dispositions, emulation, and spatial perception. Synthese, 159(3):389–416, December 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9236-z;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9236-
z.pdf.

Friston:2007:FEB

[859] Karl J. Friston and Klaas E. Stephan. Free-energy and the brain. Synthese, 159(3):417–458, December 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9237-y;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9237-y.pdf.

Gerrans:2007:MTT

[860] Philip Gerrans. Mental time travel, somatic markers and “myopia for the future”. Synthese, 159(3):459–474, December 2007. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9238-x;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9238-
x.pdf.

Bayne:2007:NCA

[861] Tim Bayne and Elisabeth Pacherie. Narrators and comparators: the architecture of agentive self-awareness. Synthese, 159(3):475–491, December 2007. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9239-9;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9239-9.pdf.

vanBenthem:2008:E

[862] Johan van Benthem, Vincent F. Hendricks, and John Symons. Editorial. Synthese, 160(1):1–4, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9274-6;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9274-6.pdf.

Atkinson:2008:RZE

[863] David Atkinson. A relativistic Zeno effect. Synthese, 160(1):5–12, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9071-7;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9071-7.pdf.

Bangu:2008:IBE

[864] Sorin Ioan Bangu. Inference to the best explanation and mathematical realism. Synthese, 160(1):13–20, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9070-8;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9070-
8.pdf.

Skyrms:2008:TRS

[865] Brian Skyrms. Trust, risk, and the social contract. Synthese, 160(1): 21–25, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9075-3;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9075-3.pdf.

Esfeld:2008:MSR

[866] Michael Esfeld and Vincent Lam. Moderate structural realism about space-time. Synthese, 160(1):27–46, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9076-2;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9076-
2.pdf.

Hawley:2008:WJ

[867] Patrick Hawley. What justifies that? Synthese, 160(1):47–61, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9078-0;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9078-0.pdf.

Woodbridge:2008:PV

[868] James A. Woodbridge and Bradley Armour-Garb. The pathology of validity. Synthese, 160(1):63–74, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9077-1;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9077-
1.pdf.

Crawford:2008:QPA

[869] Sean Crawford. Quantifiers and propositional attitudes: Quine revisited. Synthese, 160(1):75–96, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9080-6;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9080-6.pdf.

Pust:2008:HSB

[870] Joel Pust. Horgan on Sleeping Beauty. Synthese, 160(1):97–101, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9102-4;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9102-4.pdf.

Fernandez:2008:MPS

[871] Jordi Fernández. Memory, past and self. Synthese, 160(1):103–121, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9104-2;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9104-2.pdf.

Wallis:2008:CCK

[872] Charles Wallis. Consciousness, context, and know-how. Synthese, 160 (1):123–153, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9103-3;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9103-3.pdf.

Horgan:2008:SBU

[873] Terry Horgan. Synchronic Bayesian updating and the sleeping beauty problem: reply to pust. Synthese, 160(2):155–159, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9121-1;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9121-1.pdf.

Weisberg:2008:SOS

[874] Josh Weisberg. Same old, same old: the same-order representation theory of consciousness and the division of phenomenal labor. Synthese, 160 (2):161–181, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9106-0;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9106-0.pdf.

Abad:2008:ISS

[875] Jordi Valor Abad. The inclosure scheme and the solution to the paradoxes of self-reference. Synthese, 160(2):183–202, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9109-x;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9109-x.pdf.

Thompson:2008:RCI

[876] Brad Thompson. Representationalism and the conceivability of inverted spectra. Synthese, 160(2):203–213, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9110-4;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9110-
4.pdf.

Tramel:2008:HFF

[877] Peter Tramel. Haack’s foundherentism is a foundationalism. Synthese, 160(2):215–228, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9108-y;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9108-y.pdf.

Bishop:2008:DCF

[878] Robert C. Bishop. Downward causation in fluid convection. Synthese, 160 (2):229–248, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9112-2;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9112-2.pdf.

Day:2008:CIS

[879] Mark Day and George S. Botterill. Contrast, inference and scientific realism. Synthese, 160(2):249–267, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9117-x;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9117-
x.pdf.

Farkas:2008:TTT

[880] Katalin Farkas. Time, tense, truth. Synthese, 160(2):269–284, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9116-y;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9116-y.pdf.

Sprevak:2008:KPC

[881] Mark D. Sprevak. Kripke’s paradox and the Church–Turing thesis. Synthese, 160(2):285–295, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9120-2;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9120-2.pdf.

Handfield:2008:UD

[882] Toby Handfield. Unfinkable dispositions. Synthese, 160(2):297–308, January 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9148-3;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9148-3.pdf.

Zahavi:2008:MWW

[883] Dan Zahavi. The mind without, the world within. Synthese, 160(3): 309–311, February 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9081-5;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9081-5.pdf.

Smith:2008:HE

[884] A. David Smith. Husserl and externalism. Synthese, 160(3):313– 333, February 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9082-4;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9082-4.pdf.

Crowell:2008:PIN

[885] Steven Crowell. Phenomenological immanence, normativity, and semantic externalism. Synthese, 160(3):335–354, February 2008. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9083-3;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9083-
3.pdf.

Zahavi:2008:IET

[886] Dan Zahavi. Internalism, externalism, and transcendental idealism. Synthese, 160(3):355–374, February 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-006-9084-2;   http://link.springer.com/
content/pdf/10.1007/s11229-006-9084-2.pdf.

OMurchadha:2008:REI

[887] Felix O’Murchadha. Reduction, externalism and immanence in Husserl and Heidegger. Synthese, 160(3):375–395, February 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9085-1;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9085-
1.pdf.

Thompson:2008:RPM

[888] Evan Thompson. Representationalism and the phenomenology of mental imagery. Synthese, 160(3):397–415, February 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9086-0;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9086-
0.pdf.

Zinck:2008:CED

[889] Alexandra Zinck and Albert Newen. Classifying emotion: a developmental account. Synthese, 161(1):1–25, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9149-2;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9149-
2.pdf.

Cameron:2008:TNC

[890] Ross Paul Cameron. Truthmakers and necessary connections. Synthese, 161(1):27–45, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9152-7;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9152-7.pdf.

Chien:2008:SIC

[891] Arnold Chien. Scalar implicature and contrastive explanation. Synthese, 161(1):47–66, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9153-6;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9153-6.pdf.

Eagle:2008:MCA

[892] Antony Eagle. Mathematics and conceptual analysis. Synthese, 161 (1):67–88, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9151-8;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9151-8.pdf.

Huber:2008:ATB

[893] Franz Huber. Assessing theories, Bayes style. Synthese, 161(1): 89–118, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9141-x;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9141-x.pdf.

Press:2008:SUR

[894] Joel Kenton Press. The scientific use of ‘representation’ and ‘function’: avoiding explanatory vacuity. Synthese, 161(1):119–139, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9154-5;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9154-5.pdf.

Severo:2008:PII

[895] Rogério Passos Severo. “Plausible insofar as it is intelligible”: Quine on underdetermination. Synthese, 161(1):141–165, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9150-9;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9150-9.pdf.

Siebel:2008:ETT

[896] Mark Siebel and Werner Wolff. Equivalent testimonies as a touchstone of coherence measures. Synthese, 161(2):167–182, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

006-9155-4;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9155-4.pdf.

Benbaji:2008:CEG

[897] Hagit Benbaji. Constitution and the explanatory gap. Synthese, 161 (2):183–202, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9162-0;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9162-0.pdf.

Edin:2008:ABF

[898] Benoni B. Edin. Assigning biological functions: making sense of causal chains. Synthese, 161(2):203–218, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9160-2;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9160-
2.pdf.

Floridi:2008:DIS

[899] Luciano Floridi. A defence of informational structural realism. Synthese, 161(2):219–253, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9163-z;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9163-z.pdf.

Lawry:2008:AMU

[900] Jonathan Lawry. Appropriateness measures: an uncertainty model for vague concepts. Synthese, 161(2):255–269, March 2008. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9158-9;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9158-
9.pdf.

Pierson:2008:EWS

[901] Robert Pierson and Richard Reiner. Explanatory warrant for scientific realism. Synthese, 161(2):271–282, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9159-8;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9159-
8.pdf.

Thompson:2008:GM

[902] J. Robert Thompson. Grades of meaning. Synthese, 161(2):283– 308, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9161-1;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9161-1.pdf.

Tilley:2008:RRR

[903] John J. Tilley. Reasons, rational requirements, and the putative pseudo-question “why be moral?”. Synthese, 161(2):309–323, March 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9164-y;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9164-y.pdf.

Nottelmann:2008:I

[904] Nikolaj Nottelmann. Introduction. Synthese, 161(3):325–337, April 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9087-z;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9087-z.pdf.

Feldman:2008:MDE

[905] Richard Feldman. Modest deontologism in epistemology. Synthese, 161 (3):339–355, April 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9088-y;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9088-y.pdf.

Hieronymi:2008:RB

[906] Pamela Hieronymi. Responsibility for believing. Synthese, 161(3): 357–373, April 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-006-9089-x;   http://link.springer.com/content/
pdf/10.1007/s11229-006-9089-x.pdf.

Steup:2008:DF

[907] Matthias Steup. Doxastic freedom. Synthese, 161(3):375–392, April 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9090-4;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9090-4.pdf.

Montmarquet:2008:VV

[908] James Montmarquet. Virtue and voluntarism. Synthese, 161(3):393–402, April 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9091-3;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9091-3.pdf.

Audi:2008:EBD

[909] Robert Audi. The ethics of belief: doxastic self-control and intellectual virtue. Synthese, 161(3):403–418, April 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-006-9092-2;
 http://link.springer.com/content/pdf/10.1007/s11229-006-9092-
2.pdf.

Leite:2008:BOR

[910] Adam Leite. Believing one’s reasons are good. Synthese, 161(3):419–441, April 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
006-9093-1;   http://link.springer.com/content/pdf/10.1007/
s11229-006-9093-1.pdf.

Carter:2008:CWM

[911] Jessica Carter. Categories for the working mathematician: making the impossible possible. Synthese, 162(1):1–13, May 2008. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9166-9;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9166-
9.pdf.

Karbasizadeh:2008:RCL

[912] Amir Eshan Karbasizadeh. Revising the concept of lawhood: special sciences and natural kinds. Synthese, 162(1):15–30, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9165-x;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9165-x.pdf.

Comesana:2008:CTE

[913] Juan Comesaña. Could there be exactly two things? Synthese, 162 (1):31–35, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9170-0;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9170-0.pdf.

Daly:2008:MGM

[914] Chris John Daly. The methodology of genuine modal realism. Synthese, 162(1):37–52, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9176-7;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9176-7.pdf.

Fischer:2008:WNC

[915] Eugen Fischer. Wittgenstein’s ‘non-cognitivism’ — explained and vindicated. Synthese, 162(1):53–84, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9172-y;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9172-y.pdf.

Ritchie:2008:SRD

[916] Jack Ritchie. Structural realism and Davidson. Synthese, 162(1):85–100, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9171-z;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9171-z.pdf.

Schnieder:2008:WWC

[917] Benjamin Schnieder. On what we can ensure. Synthese, 162(1):101–115, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9175-8;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9175-8.pdf.

Tallant:2008:WIB

[918] Jonathan Tallant. What is it to “B” a relation? Synthese, 162 (1):117–132, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9173-x;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9173-x.pdf.

Wang:2008:ECC

[919] Linton Wang. Epistemic comparative conditionals. Synthese, 162(1): 133–156, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9174-9;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9174-9.pdf.

Ainsworth:2008:CIP

[920] Peter Mark Ainsworth. Cosmic inflation and the past hypothesis. Synthese, 162(2):157–165, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9179-4;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9179-4.pdf.

Brueckner:2008:RCC

[921] Anthony Brueckner. Reply to Coffman on closure and skepticism. Synthese, 162(2):167–171, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9180-y;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9180-y.pdf.

Coffman:2008:WT

[922] E. J. Coffman. Warrant without truth? Synthese, 162(2):173–194, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9178-5;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9178-5.pdf.

Darrigol:2008:MSP

[923] Olivier Darrigol. The modular structure of physical theories. Synthese, 162(2):195–223, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9181-x;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9181-x.pdf.

Hartimo:2008:GP

[924] Mirja Helena Hartimo. From geometry to phenomenology. Synthese, 162(2):225–233, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9177-6;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9177-6.pdf.

Schlatter:2008:WPL

[925] Mark Schlatter and Ken Aizawa. Walter Pitts and “a logical calculus”. Synthese, 162(2):235–250, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9182-9;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9182-
9.pdf.

LeMorvan:2008:EME

[926] Pierre Le Morvan. Epistemic means and ends: a reply to Hofmann. Synthese, 162(2):251–264, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9184-7;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9184-7.pdf.

Baumann:2008:SCP

[927] Peter Baumann. Single-case probabilities and the case of Monty Hall: Levy’s view. Synthese, 162(2):265–273, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9185-6;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9185-

6.pdf. See [821].

DAgostino:2008:NET

[928] Fred D’Agostino. Naturalizing the essential tension. Synthese, 162 (2):275–308, May 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9192-7;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9192-7.pdf.

Catterson:2008:I

[929] Troy Catterson. Introduction. Synthese, 162(3):309–311, June 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9254-x;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9254-x.pdf.

Shoemaker:2008:PAI

[930] Sydney Shoemaker. Persons, animals, and identity. Synthese, 162 (3):313–324, June 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9253-y;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9253-y.pdf.

Martin:2008:WRM

[931] Raymond Martin. What really matters. Synthese, 162(3):325–340, June 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9251-0;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9251-0.pdf.

Kolak:2008:RVM

[932] Daniel Kolak. Room for a view: on the metaphysical subject of personal identity. Synthese, 162(3):341–372, June 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9252-z;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9252-

z.pdf. See remarks [936].

Thomson:2008:CS

[933] Garrett Thomson. Counting subjects. Synthese, 162(3):373–384, June 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9249-7;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9249-7.pdf.

Catterson:2008:CSS

[934] Troy Catterson. Changing the subject: on the subject of subjectivity. Synthese, 162(3):385–404, June 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9250-1;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9250-1.pdf.

Schechtman:2008:DUP

[935] Marya Schechtman. Diversity in unity: practical unity and personal boundaries. Synthese, 162(3):405–423, June 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9247-9;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9247-
9.pdf.

Zovko:2008:MIC

[936] Jure Zovko. Metaphysics as interpretation of conscious life: some remarks on D. Henrich’s and D. Kolak’s thinking. Synthese, 162(3):425–438, June 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9248-8;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9248-8.pdf. See [932].

Lombard:2008:SCN

[937] Jay Lombard. Synchrnoic consciousness from a neurological point of view: the philosophical foundations for neuroethics. Synthese, 162 (3):439–450, June 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9246-x;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9246-x.pdf.

Cariani:2008:DFJ

[938] Fabrizio Cariani, Marc Pauly, and Josh Snyder. Decision framing in judgment aggregation. Synthese, 163(1):1–24, July 2008. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9306-x;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9306-
x.pdf.

Betz:2008:EDS

[939] Gregor Betz. Evaluating dialectical structures with Bayesian methods. Synthese, 163(1):25–44, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9276-4;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9276-4.pdf.

Lorini:2008:LIA

[940] Emiliano Lorini and Andreas Herzig. A logic of intention and attempt. Synthese, 163(1):45–77, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9309-7;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9309-7.pdf.

deBruin:2008:CKP

[941] Boudewijn de Bruin. Common knowledge of payoff uncertainty in games. Synthese, 163(1):79–97, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9275-5;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9275-5.pdf.

Shope:2008:ACV

[942] Robert K. Shope. Abnormality, cognitive virtues, and knowledge. Synthese, 163(1):99–118, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9313-y;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9313-y.pdf.

Carter:2008:SPM

[943] Jessica Carter. Structuralism as a philosophy of mathematical practice. Synthese, 163(2):119–131, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9169-6;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9169-6.pdf.

Sundstrom:2008:MIP

[944] Pär Sundström. Is the mystery an illusion? Papineau on the problem of consciousness. Synthese, 163(2):133–143, July 2008. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9193-6;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9193-
6.pdf.

Romdenh-Romluc:2008:FPT

[945] Komarine Romdenh-Romluc. First-person thought and the use of ‘i’. Synthese, 163(2):145–156, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9194-5;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9194-5.pdf.

Choi:2008:IDP

[946] Sungho Choi. The incompleteness of dispositional predicates. Synthese, 163(2):157–174, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9195-4;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9195-4.pdf.

Aikin:2008:MEV

[947] Scott F. Aikin. Meta-epistemology and the varieties of epistemic infinitism. Synthese, 163(2):175–185, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9196-3;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9196-
3.pdf.

Slater:2008:HND

[948] Barry Hartley Slater. Harmonising natural deduction. Synthese, 163 (2):187–198, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9197-2;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9197-2.pdf.

Hofer-Szabo:2008:SVC

[949] Gábor Hofer-Szabó. Separate- versus common-common-cause-type derivations of the Bell inequalities. Synthese, 163(2):199–215, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9198-1;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9198-1.pdf.

Johansson:2008:FCS

[950] Ingvar Johansson. Formalizing common sense: an operator-based approach to the Tibbles–Tib problem. Synthese, 163(2):217–225, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9199-0;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9199-0.pdf.

Pauly:2008:RLS

[951] Marc Pauly. On the role of language in social choice theory. Synthese, 163(2):227–243, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9200-y;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9200-y.pdf.

Hillman:2008:ERM

[952] T. Allan Hillman. The early Russell on the metaphysics of substance in Leibniz and Bradley. Synthese, 163(2):245–261, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9201-x;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9201-x.pdf.

deSa:2008:GPP

[953] Dan López de Sa. The over-generalization problem: predicates rigidly signifying the “unnatural”. Synthese, 163(2):263–272, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9211-8;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9211-8.pdf.

Grim:2008:GMG

[954] Patrick Grim, Randy Au, Nancy Louie, Robert Rosenberger, William Braynen, Evan Selinger, and Robb E. Eason. A graphic measure for game-theoretic robustness. Synthese, 163(2):273–297, July 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9212-7;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9212-7.pdf.

Staley:2008:I

[955] Kent W. Staley. Introduction. Synthese, 163(3):299–304, August 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9302-6;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9302-6.pdf.

Mayo:2008:ESP

[956] Deborah Mayo and Jean Miller. The error statistical philosopher as normative naturalist. Synthese, 163(3):305–314, August 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9303-5;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9303-5.pdf.

Wheeler:2008:MNE

[957] Gregory Wheeler and Luís Moniz Pereira. Methodological naturalism and epistemic internalism. Synthese, 163(3):315–328, August 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9300-8;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9300-8.pdf.

deCristofaro:2008:NFP

[958] Rodolfo de Cristofaro. A new formulation of the principle of indifference. Synthese, 163(3):329–339, August 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9301-7;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9301-
7.pdf.

Williamson:2008:OBP

[959] Jon Williamson. Objective Bayesianism with predicate languages. Synthese, 163(3):341–356, August 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9298-y;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9298-y.pdf.

Taper:2008:MSA

[960] Mark L. Taper, David F. Staples, and Bradley B. Shepard. Model structure adequacy analysis: selecting models on the basis of their ability to answer scientific questions. Synthese, 163(3):357–370, August 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9299-x;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9299-x.pdf.

Parker:2008:CST

[961] Wendy S. Parker. Computer simulation through an error-statistical lens. Synthese, 163(3):371–384, August 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9296-0;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9296-0.pdf.

Bartz-Beielstein:2008:HEA

[962] Thomas Bartz-Beielstein. How experimental algorithmics can benefit from Mayo’s extensions to Neyman–Pearson theory of testing. Synthese, 163(3):385–396, August 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9297-z;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9297-z.pdf.

Staley:2008:ESE

[963] Kent Staley. Error-statistical elimination of alternative hypotheses. Synthese, 163(3):397–408, August 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9294-2;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9294-2.pdf.

Zhang:2008:EPI

[964] Jiji Zhang. Error probabilities for inference of causal directions. Synthese, 163(3):409–418, August 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9295-1;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9295-1.pdf.

Ward:2008:ACE

[965] Andrew Ward and Pamela Jo Johnson. Addressing confounding errors when using non-experimental, observational data to make causal claims. Synthese, 163(3):419–432, August 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9292-4;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9292-
4.pdf.

Eberhardt:2008:SCP

[966] Frederick Eberhardt. A sufficient condition for pooling data. Synthese, 163(3):433–442, August 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9293-3;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9293-3.pdf.

Pagin:2008:IAS

[967] Peter Pagin. Indeterminacy and the analytic/synthetic distinctions: a survey. Synthese, 164(1):1–18, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9213-6;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9213-
6.pdf.

Douven:2008:EST

[968] Igor Douven. The evidential support theory of conditionals. Synthese, 164(1):19–44, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9214-5;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9214-5.pdf.

Guevara:2008:RFV

[969] Daniel Guevara. Rebutting formally valid counterexamples to the Humean “is-ought” dictum. Synthese, 164(1):45–60, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9215-4;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9215-4.pdf.

Raftopoulos:2008:PSR

[970] Athanasios Raftopoulos. Perceptual systems and realism. Synthese, 164 (1):61–91, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9216-3;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9216-3.pdf.

Baumgartner:2008:AF

[971] Michael Baumgartner and Timm Lampert. Adequate formalization. Synthese, 164(1):93–115, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9218-1;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9218-1.pdf.

Hobson:2008:FBS

[972] Kenneth Hobson. Foundational beliefs and the structure of justification. Synthese, 164(1):117–139, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9219-0;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9219-
0.pdf.

Muller:2008:FBB

[973] Thomas Müller, Nuel Belnap, and Kohei Kishida. Funny business in branching space-times: infinite modal correlations. Synthese, 164(1): 141–159, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9220-7;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9220-7.pdf.

Reydon:2008:STF

[974] Thomas A. C. Reydon. Species in three and four dimensions. Synthese, 164(2):161–184, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9221-6;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9221-6.pdf.

McCain:2008:VEC

[975] Kevin McCain. The virtues of epistemic conservatism. Synthese, 164(2): 185–200, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9222-5;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9222-5.pdf.

Schurz:2008:PA

[976] G. Schurz. Patterns of abduction. Synthese, 164(2):201–234, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9223-4;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9223-4.pdf.

Williams:2008:GA

[977] John Robert Gareth Williams. Gavagai again. Synthese, 164(2):235– 259, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9224-3;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9224-3.pdf.

Cameron:2008:TM

[978] Ross Paul Cameron. Truthmakers and modality. Synthese, 164(2): 261–280, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9225-2;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9225-2.pdf.

Draper:2008:DDB

[979] Kai Draper and Joel Pust. Diachronic Dutch books and Sleeping Beauty. Synthese, 164(2):281–287, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9226-1;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9226-
1.pdf.

Pelling:2008:EIN

[980] Charles Pelling. Exactness, inexactness, and the non-transitivity of perceptual indiscriminability. Synthese, 164(2):289–312, September 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9227-0;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9227-0.pdf.

Mancosu:2008:IIE

[981] Paolo Mancosu. Introduction: Interpolations — essays in honor of William Craig. Synthese, 164(3):313–319, October 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9350-6;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9350-
6.pdf.

Craig:2008:EPL

[982] William Craig. Elimination problems in logic: a brief history. Synthese, 164(3):321–332, October 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9352-4;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9352-4.pdf.

Craig:2008:RTT

[983] William Craig. The road to two theorems of logic. Synthese, 164(3): 333–339, October 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9353-3;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9353-3.pdf.

Feferman:2008:HLC

[984] Solomon Feferman. Harmonious logic: Craig’s interpolation theorem and its descendants. Synthese, 164(3):341–357, October 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9354-2;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9354-2.pdf.

Demopoulos:2008:SRB

[985] William Demopoulos. Some remarks on the bearing of model theory on the theory of theories. Synthese, 164(3):359–383, October 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9355-1;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9355-1.pdf.

Friedman:2008:WRL

[986] Michael Friedman. Wissenschaftslogik: The role of logic in the philosophy of science. Synthese, 164(3):385–400, October 2008. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9356-0;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9356-
0.pdf.

Vaananen:2008:CIT

[987] Jouko Väänänen. The Craig Interpolation Theorem in abstract model theory. Synthese, 164(3):401–420, October 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9357-z;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9357-
z.pdf.

DAgostino:2008:INC

[988] Giovanna D’Agostino. Interpolation in non-classical logics. Synthese, 164 (3):421–435, October 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9359-x;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9359-x.pdf.

deLavalette:2008:ICS

[989] Gerard R. Renardel de Lavalette. Interpolation in computing science: the semantics of modularization. Synthese, 164(3):437–450, October 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9358-y;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9358-y.pdf.

vanBenthem:2008:MFI

[990] Johan van Benthem. The many faces of interpolation. Synthese, 164 (3):451–460, October 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9351-5;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9351-5.pdf.

Dieveney:2008:OI

[991] Patrick Dieveney. Ontological infidelity. Synthese, 165(1):1–12, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9228-z;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9228-z.pdf.

Andrews:2008:YNP

[992] Kristin Andrews. It’s in your nature: a pluralistic folk psychology. Synthese, 165(1):13–29, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9230-5;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9230-5.pdf.

Martin:2008:LBG

[993] John Neil Martin. The lover of the beautiful and the good: Platonic foundations of aesthetic and moral value. Synthese, 165(1):31–51, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9231-4;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9231-4.pdf.

McGivern:2008:RLM

[994] Patrick McGivern. Reductive levels and multi-scale structure. Synthese, 165(1):53–75, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9232-3;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9232-3.pdf.

Yaqub:2008:TTD

[995] Aladdin M. Yaqub. Two types of deflationism. Synthese, 165(1):77– 106, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9241-2;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9241-2.pdf.

Williams:2008:PKK

[996] John N. Williams. Propositional knowledge and know-how. Synthese, 165(1):107–125, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9242-1;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9242-1.pdf.

Espinoza:2008:SIA

[997] Nicolas Espinoza. The small improvement argument. Synthese, 165(1): 127–139, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9243-0;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9243-0.pdf.

Klein:2008:DIS

[998] Colin Klein. Dispositional implementation solves the superfluous structure problem. Synthese, 165(1):141–153, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9244-z;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9244-
z.pdf.

Liu:2008:E

[999] Fenrong Liu, Frank Veltman, and Minghui Xiong. Editorial. Synthese, 165(2):155–157, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9405-8;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9405-8.pdf.

Alechina:2008:PBB

[1000] Natasha Alechina, Mark Jago, and Brian Logan. Preference-based belief revision for rule-based agents. Synthese, 165(2):159–177, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9364-0;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9364-0.pdf.

Baltag:2008:PDB

[1001] Alexandru Baltag and Sonja Smets. Probabilistic dynamic belief revision. Synthese, 165(2):179–202, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9369-8;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9369-
8.pdf.

Primiero:2008:MMA

[1002] Giuseppe Primiero and Joke Meheus. Majority merging by adaptive counting. Synthese, 165(2):203–223, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9370-2;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9370-
2.pdf.

Renne:2008:PPC

[1003] Bryan Renne. Public and private communication are different: results on relative expressivity. Synthese, 165(2):225–245, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9395-6;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9395-6.pdf.

vanBenthem:2008:MSG

[1004] Johan van Benthem, Sujata Ghosh, and Fenrong Liu. Modelling simultaneous games in dynamic logic. Synthese, 165(2):247–268, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9390-y;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9390-y.pdf.

Aagotnes:2008:QCL

[1005] Thomas Ågotnes, Wiebe van der Hoek, and Michael Wooldridge. Quantified coalition logic. Synthese, 165(2):269–294, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9363-1;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9363-1.pdf.

Yamada:2008:LDS

[1006] Tomoyuki Yamada. Logical dynamics of some speech acts that affect obligations and preferences. Synthese, 165(2):295–315, November 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9368-9;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9368-9.pdf.

Stojanovic:2008:SPD

[1007] Isidora Stojanovic. The semantics/pragmatics distinction. Synthese, 165 (3):317–319, December 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9190-9;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9190-9.pdf.

Carston:2008:LCS

[1008] Robyn Carston. Linguistic communication and the semantics/pragmatics distinction. Synthese, 165(3):321–345, December 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9191-8;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9191-
8.pdf.

Korta:2008:PC

[1009] Kepa Korta and John Perry. The pragmatic circle. Synthese, 165(3): 347–357, December 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9188-3;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9188-3.pdf.

Gauker:2008:ZTP

[1010] Christopher Gauker. Zero tolerance for pragmatics. Synthese, 165(3): 359–371, December 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9189-2;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9189-2.pdf.

Gillon:2008:SPD

[1011] Brendan S. Gillon. On the semantics/pragmatics distinction. Synthese, 165(3):373–384, December 2008. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9186-5;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9186-5.pdf.

Katsos:2008:SPI

[1012] Napoleon Katsos. The semantics/pragmatics interface from an experimental perspective: the case of scalar implicature. Synthese, 165(3): 385–401, December 2008. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9187-4;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9187-4.pdf.

Dokic:2009:MET

[1013] Jérôme Dokic and Paul Égré. Margin for error and the transparency of knowledge. Synthese, 166(1):1–20, January 2009. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9245-y;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9245-
y.pdf.

Abrams:2009:WDB

[1014] Marshall Abrams. What determines biological fitness? The problem of the reference environment. Synthese, 166(1):21–40, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9255-9;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9255-9.pdf.

Link:2009:WL

[1015] Montgomery Link. Wittgenstein and logic. Synthese, 166(1):41–54, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9256-8;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9256-8.pdf.

Schulz:2009:CCB

[1016] Armin W. Schulz. Condorcet and communitarianism: Boghossian’s fallacious inference. Synthese, 166(1):55–68, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9257-7;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9257-
7.pdf.

Krohs:2009:FBC

[1017] Ulrich Krohs. Functions as based on a concept of general design. Synthese, 166(1):69–89, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9258-6;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9258-6.pdf.

Nelson:2009:HHM

[1018] Kevin Nelson. How and how not to make predictions with temporal Copernicanism. Synthese, 166(1):91–111, January 2009. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9259-5;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9259-
5.pdf.

Batens:2009:YFM

[1019] Diderik Batens, Kristof De Clercq, Peter Verdée, and Joke Meheus. Yes fellows, most human reasoning is complex. Synthese, 166(1): 113–131, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9268-4;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9268-4.pdf.

Jaworski:2009:LHQ

[1020] William Jaworski. The logic of how-questions. Synthese, 166(1): 133–155, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9269-3;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9269-3.pdf.

Turri:2009:RAI

[1021] John Turri. On the regress argument for infinitism. Synthese, 166(1): 157–163, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9270-x;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9270-x.pdf.

Nelson:2009:BUT

[1022] Kevin Nelson. On background: using two-argument chance. Synthese, 166(1):165–186, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9271-9;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9271-9.pdf.

Epstein:2009:OIR

[1023] Brian Epstein. Ontological individualism reconsidered. Synthese, 166 (1):187–213, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9272-8;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9272-8.pdf.

Brogaard:2009:IRT

[1024] Berit Brogaard. Introduction to relative truth. Synthese, 166(2): 215–229, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9288-0;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9288-0.pdf.

MacFarlane:2009:NC

[1025] John MacFarlane. Nonindexical contextualism. Synthese, 166(2):231– 250, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9286-2;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9286-2.pdf.

Egan:2009:BBS

[1026] Andy Egan. Billboards, bombs and shotgun weddings. Synthese, 166 (2):251–279, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9284-4;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9284-4.pdf.

Glanzberg:2009:STR

[1027] Michael Glanzberg. Semantics and truth relative to a world. Synthese, 166(2):281–307, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9285-3;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9285-3.pdf.

Moruzzi:2009:TAA

[1028] Sebastiano Moruzzi and Crispin Wright. Trumping assessments and the Aristotelian future. Synthese, 166(2):309–331, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9282-6;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9282-6.pdf.

Weatherson:2009:CIR

[1029] Brian Weatherson. Conditionals and indexical relativism. Synthese, 166 (2):333–357, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9283-5;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9283-5.pdf.

Lasersohn:2009:RTS

[1030] Peter Lasersohn. Relative truth, speaker commitment, and control of implicit arguments. Synthese, 166(2):359–374, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9280-8;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9280-8.pdf.

Kolbel:2009:ER

[1031] Max Kölbel. The evidence for relativism. Synthese, 166(2):375–395, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9281-7;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9281-7.pdf.

Pritchard:2009:DER

[1032] Duncan Pritchard. Defusing epistemic relativism. Synthese, 166(2): 397–412, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9278-2;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9278-2.pdf.

Capps:2009:CMR

[1033] David Capps, Michael P. Lynch, and Daniel Massey. A coherent moral relativism. Synthese, 166(2):413–430, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9279-1;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9279-
1.pdf.

Hales:2009:MRE

[1034] Steven D. Hales. Moral relativism and evolutionary psychology. Synthese, 166(2):431–447, January 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9287-1;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9287-1.pdf.

Bickhard:2009:IIS

[1035] Mark H. Bickhard. Interactivism: introduction to the special issue. Synthese, 166(3):449–451, February 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9371-1;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9371-1.pdf.

Campbell:2009:PBM

[1036] Richard Campbell. A process-based model for an interactive ontology. Synthese, 166(3):453–477, February 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9372-0;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9372-0.pdf.

Seibt:2009:FEI

[1037] Johanna Seibt. Forms of emergent interaction in general process theory. Synthese, 166(3):479–512, February 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9373-z;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9373-
z.pdf.

Hooker:2009:IBC

[1038] C. A. Hooker. Interaction and bio-cognitive order. Synthese, 166(3): 513–546, February 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9374-y;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9374-y.pdf.

Bickhard:2009:IM

[1039] Mark H. Bickhard. The interactivist model. Synthese, 166(3):547– 591, February 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9375-x;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9375-x.pdf.

Levine:2009:PEA

[1040] Alex Levine. Partition epistemology and arguments from analogy. Synthese, 166(3):593–600, February 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9376-9;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9376-9.pdf.

Placek:2009:IEL

[1041] Tomasz Placek and Leszek Wroński. On infinite EPR-like correlations. Synthese, 167(1):1–32, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9273-7;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9273-7.pdf.

Dunlop:2009:WEG

[1042] Katherine Dunlop. Why Euclid’s geometry brooked no doubt: J. H. Lambert on certainty and the existence of models. Synthese, 167 (1):33–65, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9277-3;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9277-3.pdf.

Lupher:2009:PCP

[1043] Tracy Lupher. A physical critique of physical causation. Synthese, 167 (1):67–80, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9289-z;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9289-z.pdf.

Briggs:2009:BBB

[1044] Rachael Briggs. The big bad bug bites anti-realists about chance. Synthese, 167(1):81–92, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9290-6;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9290-6.pdf.

Verdee:2009:ALU

[1045] Peter Verdée. Adaptive logics using the minimal abnormality strategy are Π_{1}^{1}-complex. Synthese, 167(1):93–104, March 2009. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9291-5;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9291-
5.pdf.

Textor:2009:RFT

[1046] Mark Textor. A repair of Frege’s theory of thoughts. Synthese, 167 (1):105–123, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9304-z;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9304-z.pdf.

Weisberg:2009:LIB

[1047] Jonathan Weisberg. Locating IBE in the Bayesian framework. Synthese, 167(1):125–143, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9305-y;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9305-y.pdf.

Johnston:2009:TOL

[1048] Colin Johnston. Tractarian objects and logical categories. Synthese, 167 (1):145–161, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9307-9;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9307-9.pdf.

McKeon:2009:PLO

[1049] Matthew William McKeon. A plea for logical objects. Synthese, 167 (1):163–182, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9308-8;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9308-8.pdf.

Kowalenko:2009:HTA

[1050] Robert Kowalenko. How (not) to think about idealisation and ceteris paribus-laws. Synthese, 167(1):183–201, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9310-1;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9310-
1.pdf.

Floridi:2009:I

[1051] Luciano Floridi and Sebastian Sequoiah-Grayson. Introduction. Synthese, 167(2):203–205, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9462-7;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9462-7.pdf.

Abramsky:2009:IB

[1052] Samson Abramsky and Jouko Väänänen. From IF to BI. Synthese, 167 (2):207–230, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9415-6;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9415-6.pdf.

Allo:2009:RAD

[1053] Patrick Allo. Reasoning about data and information. Synthese, 167 (2):231–249, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9407-6;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9407-6.pdf.

vanBenthem:2009:IIL

[1054] Johan van Benthem. The information in intuitionistic logic. Synthese, 167(2):251–270, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9408-5;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9408-5.pdf.

DAgostino:2009:ESD

[1055] Marcello D’Agostino and Luciano Floridi. The enduring scandal of deduction. Synthese, 167(2):271–315, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9409-4;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9409-
4.pdf.

Floridi:2009:LFI

[1056] Luciano Floridi. Logical fallacies as informational shortcuts. Synthese, 167(2):317–325, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9410-y;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9410-y.pdf.

Jago:2009:LIE

[1057] Mark Jago. Logical information and epistemic space. Synthese, 167 (2):327–341, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9411-x;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9411-x.pdf.

Mares:2009:GIR

[1058] Edwin D. Mares. General information in relevant logic. Synthese, 167 (2):343–362, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9412-9;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9412-9.pdf.

Primiero:2009:ELB

[1059] Giuseppe Primiero. An epistemic logic for becoming informed. Synthese, 167(2):363–389, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9413-8;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9413-8.pdf.

Sadrzadeh:2009:ORR

[1060] Mehrnoosh Sadrzadeh. Ockham’s razor and reasoning about information flow. Synthese, 167(2):391–408, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9414-7;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9414-
7.pdf.

Sequoiah-Grayson:2009:PIL

[1061] Sebastian Sequoiah-Grayson. A positive information logic for inferential information. Synthese, 167(2):409–431, March 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9406-7;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9406-
7.pdf.

Aizawa:2009:EI

[1062] Kenneth Aizawa. Editor’s introduction. Synthese, 167(3):433–438, April 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9384-9;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9384-9.pdf.

Shapiro:2009:MSM

[1063] Lawrence Shapiro. Making sense of mirror neurons. Synthese, 167 (3):439–456, April 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9385-8;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9385-8.pdf.

Polger:2009:EEM

[1064] Thomas W. Polger. Evaluating the evidence for multiple realization. Synthese, 167(3):457–472, April 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9386-7;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9386-7.pdf.

Richardson:2009:MRM

[1065] Robert C. Richardson. Multiple realization and methodological pluralism. Synthese, 167(3):473–492, April 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-008-9387-6;   http://link.springer.
com/content/pdf/10.1007/s11229-008-9387-6.pdf.

Aizawa:2009:NMR

[1066] Ken Aizawa. Neuroscience and multiple realization: a reply to Bechtel and Mundale. Synthese, 167(3):493–510, April 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9388-5;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9388-
5.pdf.

Sullivan:2009:MEP

[1067] Jacqueline A. Sullivan. The multiplicity of experimental protocols: a challenge to reductionist and non-reductionist models of the unity of neuroscience. Synthese, 167(3):511–539, April 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9389-4;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9389-4.pdf.

Balaguer:2009:WTN

[1068] Mark Balaguer. Why there are no good arguments for any interesting version of determinism. Synthese, 168(1):1–21, May 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9459-2;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9459-2.pdf.

Douven:2009:CSD

[1069] Igor Douven. Can the skepticism debate be resolved? Synthese, 168 (1):23–52, May 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9311-0;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9311-0.pdf.

Donnelly:2009:MVE

[1070] Maureen Donnelly. Mereological vagueness and existential vagueness. Synthese, 168(1):53–79, May 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9312-z;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9312-z.pdf.

Barnes:2009:IIC

[1071] Elizabeth Barnes. Indeterminacy, identity and counterparts: Evans reconsidered. Synthese, 168(1):81–96, May 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-008-9314-x;   http://link.springer.
com/content/pdf/10.1007/s11229-008-9314-x.pdf.

Hetherington:2009:SPN

[1072] Stephen Hetherington. Sceptical possibilities? No worries. Synthese, 168(1):97–118, May 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9332-8;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9332-8.pdf.

Ellerman:2009:CDC

[1073] David Ellerman. Counting distinctions: on the conceptual foundations of Shannon’s information theory. Synthese, 168(1):119–149, May 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9333-7;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9333-7.pdf.

Floridi:2009:ADO

[1074] Luciano Floridi. Against digital ontology. Synthese, 168(1):151–178, May 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9334-6;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9334-6.pdf.

Hughes:2009:AUI

[1075] Jesse Hughes. An artifact is to use: an introduction to instrumental functions. Synthese, 168(1):179–199, May 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9335-5;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9335-
5.pdf.

Adam:2009:LFO

[1076] C. Adam, A. Herzig, and D. Longin. A logical formalization of the OCC theory of emotions. Synthese, 168(2):201–248, May 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9460-9;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9460-9.pdf.

Huss:2009:TCT

[1077] Brian Huss. Three challenges (and three replies) to the ethics of belief. Synthese, 168(2):249–271, May 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9394-7;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9394-7.pdf.

Spiekermann:2009:SYN

[1078] Kai P. Spiekermann. Sort out your neighbourhood. Synthese, 168 (2):273–294, May 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9424-5;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9424-5.pdf.

Kim:2009:SBS

[1079] Namjoong Kim. Sleeping beauty and shifted Jeffrey conditionalization. Synthese, 168(2):295–312, May 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9443-2;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9443-2.pdf.

Olsen:2009:TSE

[1080] Jan Kyrre Berg Olsen and Frederik Voetmann Christiansen. Technology and science epistemology, rationality and the empirical turn. Synthese, 168(3):313–318, June 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9452-1;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9452-1.pdf.

Auyang:2009:KSE

[1081] Sunny Y. Auyang. Knowledge in science and engineering. Synthese, 168(3):319–331, June 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9453-0;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9453-0.pdf.

Bucciarelli:2009:EIE

[1082] Louis L. Bucciarelli. The epistemic implications of engineering rhetoric. Synthese, 168(3):333–356, June 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9454-z;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9454-z.pdf.

Misa:2009:FFF

[1083] Thomas J. Misa. Findings follow framings: navigating the empirical turn. Synthese, 168(3):357–375, June 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9447-y;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9447-
y.pdf.

Selinger:2009:TRF

[1084] Evan Selinger. Towards a reflexive framework for development: technology transfer after the empirical turn. Synthese, 168(3):377–403, June 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9450-3;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9450-3.pdf.

Rip:2009:TPO

[1085] Arie Rip. Technology as prospective ontology. Synthese, 168(3):405–422, June 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

008-9449-9;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9449-9.pdf.

Hansson:2009:CJ

[1086] Sven Ove Hansson. From the casino to the jungle. Synthese, 168 (3):423–432, June 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9444-1;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9444-1.pdf.

Tiles:2009:TPG

[1087] Mary Tiles. Technology and the possibility of global environmental science. Synthese, 168(3):433–452, June 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9448-x;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9448-
x.pdf.

Ihde:2009:VCB

[1088] Don Ihde. From da Vinci to CAD and beyond. Synthese, 168(3):453–467, June 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9445-0;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9445-0.pdf.

Pickering:2009:BDC

[1089] Andrew Pickering. Beyond design: cybernetics, biological computers and hylozoism. Synthese, 168(3):469–491, June 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9446-z;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9446-
z.pdf.

Castagnino:2009:GNE

[1090] Mario Castagnino and Olimpia Lombardi. The global non-entropic arrow of time: from global geometrical asymmetry to local energy flow. Synthese, 169(1):1–25, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9495-y;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9495-y.pdf.

Earman:2009:ESA

[1091] John Earman. Essential self-adjointness: implications for determinism and the classical-quantum correspondence. Synthese, 169(1):27–50, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9341-7;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9341-7.pdf.

Baave:2009:DTR

[1092] Arvid Båve. A deflationary theory of reference. Synthese, 169(1):51–73, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9336-4;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9336-4.pdf.

Campbell:2009:KEP

[1093] Neil Campbell and Dwayne Moore. On Kim’s exclusion principle. Synthese, 169(1):75–90, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9337-3;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9337-3.pdf.

Earman:2009:DLP

[1094] John Earman, Christopher Smeenk, and Christian Wüthrich. Do the laws of physics forbid the operation of time machines? Synthese, 169 (1):91–124, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9338-2;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9338-2.pdf.

Carrara:2009:FGM

[1095] Massimiliano Carrara and Pieter E. Vermaas. The fine-grained metaphysics of artifactual and biological functional kinds. Synthese, 169 (1):125–143, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9339-1;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9339-1.pdf.

Weiskopf:2009:PC

[1096] Daniel Aaron Weiskopf. The plurality of concepts. Synthese, 169 (1):145–173, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9340-8;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9340-8.pdf.

Rescorla:2009:PCR

[1097] Michael Rescorla. Predication and cartographic representation. Synthese, 169(1):175–200, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9343-5;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9343-5.pdf.

Riggs:2009:TPE

[1098] Wayne Riggs. Two problems of easy credit. Synthese, 169(1):201–216, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9342-6;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9342-6.pdf.

Horsten:2009:UPA

[1099] Leon Horsten and Philip Welch. The undecidability of propositional adaptive logic. Synthese, 169(1):217–218, July 2009. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9496-x;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9496-
x.pdf.

Aagotnes:2009:LII

[1100] Thomas Ågotnes, Johan van Benthem, and Eric Pacuit. Logic and intelligent interaction. Synthese, 169(2):219–221, July 2009. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9558-0;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9558-
0.pdf.

Kurzen:2009:RAC

[1101] Lena Kurzen. Reasoning about cooperation, actions and preferences. Synthese, 169(2):223–240, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9551-7;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9551-7.pdf.

Sack:2009:EPD

[1102] Joshua Sack. Extending probabilistic dynamic epistemic logic. Synthese, 169(2):241–257, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9555-3;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9555-3.pdf.

Hoshi:2009:DEL

[1103] Tomohiro Hoshi and Audrey Yap. Dynamic epistemic logic with branching temporal structures. Synthese, 169(2):259–281, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9552-6;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9552-6.pdf.

Velazquez-Quesada:2009:IU

[1104] Fernando Raymundo Velázquez-Quesada. Inference and update. Synthese, 169(2):283–300, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9556-2;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9556-2.pdf.

Baltag:2009:KHR

[1105] Alexandru Baltag, Sonja Smets, and Jonathan Alexander Zvesper. Keep ‘hoping’ for rationality: a solution to the backward induction paradox. Synthese, 169(2):301–333, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9559-z;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9559-
z.pdf.

Roy:2009:IIT

[1106] Olivier Roy. Intentions and interactive transformations of decision problems. Synthese, 169(2):335–349, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9553-5;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9553-
5.pdf.

Broersen:2009:DEM

[1107] Jan Broersen, Rosja Mastop, John-Jules Meyer, and Paolo Turrini. Determining the environment: a modal logic for closed interaction.

Synthese, 169(2):351–369, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9550-8;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9550-8.pdf.

Gierasimczuk:2009:BLT

[1108] Nina Gierasimczuk. Bridging learning theory and dynamic epistemic logic. Synthese, 169(2):371–384, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-009-9549-1;   http://link.springer.
com/content/pdf/10.1007/s11229-009-9549-1.pdf.

Alechina:2009:VTM

[1109] Natasha Alechina, Brian Logan, Hoang Nga Nguyen, and Abdur Rakib. Verifying time, memory and communication bounds in systems of reasoning agents. Synthese, 169(2):385–403, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9557-1;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9557-1.pdf.

vanDitmarsch:2009:IF

[1110] Hans van Ditmarsch, Andreas Herzig, Jérôme Lang, and Pierre Marquis. Introspective forgetting. Synthese, 169(2):405–423, July 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9554-4;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9554-4.pdf.

Frigg:2009:E

[1111] Roman Frigg, Stephan Hartmann, and Cyrille Imbert. Editorial. Synthese, 169(3):425, August 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9439-y;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9439-y.pdf.

Batterman:2009:IM

[1112] Robert W. Batterman. Idealization and modeling. Synthese, 169(3): 427–446, August 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9436-1;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9436-1.pdf.

Parker:2009:CUD

[1113] Matthew W. Parker. Computing the uncomputable; or, the discrete charm of second-order simulacra. Synthese, 169(3):447–463, August 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9441-4;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9441-4.pdf.

Mattingly:2009:PPA

[1114] James Mattingly and Walter Warwick. Projectible predicates in analogue and simulated systems. Synthese, 169(3):465–482, August 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9433-4;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9433-4.pdf.

Parker:2009:DMR

[1115] Wendy S. Parker. Does matter really matter? Computer simulations, experiments, and materiality. Synthese, 169(3):483–496, August 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9434-3;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9434-3.pdf.

Gelfert:2009:RRC

[1116] Axel Gelfert. Rigorous results, cross-model justification, and the transfer of empirical warrant: the case of many-body models in physics. Synthese, 169(3):497–519, August 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9431-6;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9431-
6.pdf.

Schlimm:2009:LEM

[1117] Dirk Schlimm. Learning from the existence of models: On psychic machines, tortoises, and computer simulations. Synthese, 169(3): 521–538, August 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9432-5;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9432-5.pdf.

Grune-Yanoff:2009:EPA

[1118] Till Grüne-Yanoff. The explanatory potential of artificial societies. Synthese, 169(3):539–555, August 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9429-0;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9429-0.pdf.

Barberousse:2009:CSE

[1119] Anouk Barberousse, Sara Franceschelli, and Cyrille Imbert. Computer simulations as experiments. Synthese, 169(3):557–574, August 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9430-7;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9430-7.pdf.

Winsberg:2009:TTM

[1120] Eric Winsberg. A tale of two methods. Synthese, 169(3):575–592, August 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9437-0;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9437-0.pdf.

Frigg:2009:PSH

[1121] Roman Frigg and Julian Reiss. The philosophy of simulation: hot new issues or same old stew? Synthese, 169(3):593–613, August 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9438-z;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9438-z.pdf.

Humphreys:2009:PNC

[1122] Paul Humphreys. The philosophical novelty of computer simulation methods. Synthese, 169(3):615–626, August 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9435-2;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9435-
2.pdf.

deSa:2009:RUT

[1123] Dan López de Sa. Relativizing utterance-truth? Synthese, 170(1): 1–5, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9345-3;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9345-3.pdf.

Williams:2009:UAU

[1124] Neil Edward Williams. The ungrounded argument is unfounded: a response to Mumford. Synthese, 170(1):7–19, September 2009. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9344-4;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9344-
4.pdf.

Pedroso:2009:TAA

[1125] Makmiller Pedroso. On three arguments against categorical structuralism. Synthese, 170(1):21–31, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-008-9346-2;   http://link.springer.
com/content/pdf/10.1007/s11229-008-9346-2.pdf.

Ferreiros:2009:HLM

[1126] José Ferreirós. Hilbert, logicism, and mathematical existence. Synthese, 170(1):33–70, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9347-1;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9347-1.pdf.

Baumgartner:2009:UDC

[1127] Michael Baumgartner. Uncovering deterministic causal structures: a Boolean approach. Synthese, 170(1):71–96, September 2009. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9348-0;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9348-
0.pdf.

Adams:2009:EEK

[1128] Marcus P. Adams. Empirical evidence and the knowledge-that/ knowledge-how distinction. Synthese, 170(1):97–114, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9349-z;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9349-z.pdf.

Leuenberger:2009:WGS

[1129] Stephan Leuenberger. What is global supervenience? Synthese, 170(1): 115–129, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9360-4;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9360-4.pdf.

Paez:2009:AEE

[1130] Andrés Páez. Artificial explanations: the epistemological interpretation of explanation in AI. Synthese, 170(1):131–146, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9361-3;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9361-3.pdf.

Turri:2009:GAA

[1131] John Turri. On the general argument against internalism. Synthese, 170(1):147–153, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9362-2;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9362-2.pdf.

Aakerman:2009:PP

[1132] Jonas Åkerman. A plea for pragmatics. Synthese, 170(1):155–167, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9365-z;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9365-z.pdf.

Matthewson:2009:STM

[1133] John Matthewson and Michael Weisberg. The structure of tradeoffs in model building. Synthese, 170(1):169–190, September 2009. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9366-y;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9366-
y.pdf.

Horowitz:2009:TZH

[1134] Amir Horowitz. Turning the zombie on its head. Synthese, 170(1): 191–210, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.

1007/s11229-008-9367-x;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9367-x.pdf.

Crane:2009:I

[1135] Tim Crane and Brian P. McLaughlin. Introduction. Synthese, 170(2): 211–215, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9579-8;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9579-8.pdf.

Loewer:2009:WTA

[1136] Barry Loewer. Why is there anything except physics? Synthese, 170(2): 217–233, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9580-2;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9580-2.pdf.

Schneider:2009:LCE

[1137] Susan Schneider. LOT, CTM, and the elephant in the room. Synthese, 170(2):235–250, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9581-1;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9581-1.pdf.

McLaughlin:2009:SR

[1138] Brian P. McLaughlin. Systematicity redux. Synthese, 170(2):251– 274, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9582-0;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9582-0.pdf.

Segal:2009:KMS

[1139] Gabriel Segal. Keep making sense. Synthese, 170(2):275–287, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9583-z;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9583-z.pdf.

Edwards:2009:WCD

[1140] Kevan Edwards. What concepts do. Synthese, 170(2):289–310, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

009-9584-y;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9584-y.pdf.

Balog:2009:JFN

[1141] Katalin Balog. Jerry Fodor on non-conceptual content. Synthese, 170(2): 311–320, September 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9585-x;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9585-x.pdf.

Linnebo:2009:I

[1142] Øystein Linnebo. Introduction. Synthese, 170(3):321–329, October 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9267-5;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9267-5.pdf.

Uzquiano:2009:BCG

[1143] Gabriel Uzquiano. Bad company generalized. Synthese, 170(3):331– 347, October 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9266-6;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9266-6.pdf.

Cook:2009:HBB

[1144] Roy T. Cook. Hume’s big brother: counting concepts and the bad company objection. Synthese, 170(3):349–369, October 2009. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9264-8;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9264-
8.pdf.

Linnebo:2009:BCT

[1145] Øystein Linnebo. Bad company tamed. Synthese, 170(3):371–391, October 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
007-9265-7;   http://link.springer.com/content/pdf/10.1007/
s11229-007-9265-7.pdf.

Eklund:2009:BCN

[1146] Matti Eklund. Bad company and neo-Fregean philosophy. Synthese, 170 (3):393–414, October 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-007-9262-x;   http://link.springer.com/content/
pdf/10.1007/s11229-007-9262-x.pdf.

Ebert:2009:GBU

[1147] Philip Ebert and Stewart Shapiro. The good, the bad and the ugly. Synthese, 170(3):415–441, October 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-007-9263-9;   http://link.springer.com/
content/pdf/10.1007/s11229-007-9263-9.pdf.

MacFarlane:2009:DVT

[1148] John MacFarlane. Double vision: two questions about the neo-Fregean program. Synthese, 170(3):443–456, October 2009. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9260-z;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9260-
z.pdf.

Hale:2009:FRC

[1149] Bob Hale and Crispin Wright. Focus restored: Comments on John MacFarlane. Synthese, 170(3):457–482, October 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-007-9261-y;
 http://link.springer.com/content/pdf/10.1007/s11229-007-9261-
y.pdf.

Moro:2009:NCF

[1150] Rodrigo Moro. On the nature of the conjunction fallacy. Synthese, 171 (1):1–24, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9377-8;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9377-8.pdf.

Pelczar:2009:KAO

[1151] Michael Pelczar. The knowledge argument, the open question argument, and the moral problem. Synthese, 171(1):25–45, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9378-7;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9378-7.pdf.

Cantwell:2009:CR

[1152] John Cantwell. Conditionals in reasoning. Synthese, 171(1):47–75, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9379-6;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9379-6.pdf.

Fischer:2009:PPS

[1153] Eugen Fischer. Philosophical pictures and secondary qualities. Synthese, 171(1):77–110, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9380-0;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9380-0.pdf.

Halpern:2009:EUL

[1154] Joseph Y. Halpern and Riccardo Pucella. Evidence with uncertain likelihoods. Synthese, 171(1):111–133, November 2009. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9381-z;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9381-
z.pdf.

Meyers:2009:PIP

[1155] C. D. Meyers and Sara Waller. Psychological investigations: the private language argument and inferences in contemporary cognitive science. Synthese, 171(1):135–156, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9382-y;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9382-
y.pdf.

Yap:2009:LSB

[1156] Audrey Yap. Logical structuralism and Benacerraf’s problem. Synthese, 171(1):157–173, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9383-x;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9383-x.pdf.

Collins:2009:LCL

[1157] John Collins. The limits of conceivability: logical cognitivism and the language faculty. Synthese, 171(1):175–194, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9391-x;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9391-x.pdf.

Walicki:2009:RPT

[1158] Michal Walicki. Reference, paradoxes and truth. Synthese, 171(1): 195–226, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9392-9;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9392-9.pdf.

vanDitmarsch:2009:LCC

[1159] Hans van Ditmarsch, Brian Hill, and Ondrej Majer. Logic of change, change of logic. Synthese, 171(2):227–234, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9648-z;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9648-z.pdf.

Bradley:2009:RIA

[1160] Richard Bradley. Revising incomplete attitudes. Synthese, 171(2):235– 256, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9638-1;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9638-1.pdf.

Hansson:2009:PBC

[1161] Sven Ove Hansson. Preference-based choice functions: a generalized approach. Synthese, 171(2):257–269, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9650-5;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9650-
5.pdf.

Genot:2009:GII

[1162] Emmanuel J. Genot. The game of inquiry: the interrogative approach to inquiry and belief revision theory. Synthese, 171(2):271–289, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9639-0;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9639-0.pdf.

Roy:2009:DEH

[1163] O. Roy. A dynamic-epistemic hybrid logic for intentions and information changes in strategic games. Synthese, 171(2):291–320, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9644-3;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9644-3.pdf.

Herzig:2009:DIA

[1164] Andreas Herzig, Tiago de Lima, and Emiliano Lorini. On the dynamics of institutional agreements. Synthese, 171(2):321–355, November 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9645-2;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9645-2.pdf.

Kallestrup:2009:I

[1165] Jesper Kallestrup and Duncan Pritchard. Introduction. Synthese, 171(3): 357–358, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9329-3;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9329-3.pdf.

Wright:2009:FBP

[1166] Crispin Wright. Foreword: on becoming a philosopher. Synthese, 171(3): 359–364, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9321-y;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9321-y.pdf.

Coliva:2009:SKC

[1167] Annalisa Coliva. Self-knowledge and commitments. Synthese, 171(3): 365–375, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9322-x;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9322-x.pdf.

Kallestrup:2009:CRC

[1168] Jesper Kallestrup. Conceivability, rigidity and counterpossibles. Synthese, 171(3):377–386, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9319-5;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9319-5.pdf.

Gundersen:2009:DCS

[1169] Lars Bo Gundersen. Disjunctivism, contextualism and the sceptical aporia. Synthese, 171(3):387–397, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9320-z;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9320-
z.pdf.

Greenough:2009:WIQ

[1170] Patrick Greenough. On what it is to be in a quandary. Synthese, 171(3): 399–408, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9317-7;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9317-7.pdf.

Hamilton:2009:MSC

[1171] Andy Hamilton. Memory and self-consciousness: immunity to error through misidentification. Synthese, 171(3):409–417, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9318-6;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9318-6.pdf.

Marion:2009:RAR

[1172] Mathieu Marion. Radical anti-realism, Wittgenstein and the length of proofs. Synthese, 171(3):419–432, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9315-9;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9315-
9.pdf.

Miller:2009:PQS

[1173] Alexander Miller. Primary qualities, secondary qualities and the truth about intention. Synthese, 171(3):433–442, December 2009. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9316-8;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9316-
8.pdf.

Pedersen:2009:EVR

[1174] Nikolaj Jang Pedersen. Entitlement, value and rationality. Synthese, 171 (3):443–457, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9330-x;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9330-x.pdf.

Philie:2009:ERI

[1175] Patrice Philie. Entitlement as a response to I–II–III scepticism. Synthese, 171(3):459–466, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9327-5;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9327-5.pdf.

Pritchard:2009:WCM

[1176] Duncan Pritchard. Wright contra McDowell on perceptual knowledge and scepticism. Synthese, 171(3):467–479, December 2009. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9328-4;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9328-
4.pdf.

Rosenkranz:2009:LEV

[1177] Sven Rosenkranz. Liberalism, entitlement, and verdict exclusion. Synthese, 171(3):481–497, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9331-9;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9331-9.pdf.

Tanney:2009:RR

[1178] Julia Tanney. Real rules. Synthese, 171(3):499–507, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9326-6;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9326-6.pdf.

Zalabardo:2009:OSR

[1179] José L. Zalabardo. One strand in the rule-following considerations. Synthese, 171(3):509–519, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9323-9;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9323-9.pdf.

Weiss:2009:MDI

[1180] Bernhard Weiss. Minimalism deflated: independence without substance. Synthese, 171(3):521–529, December 2009. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-008-9324-8;   http://link.springer.
com/content/pdf/10.1007/s11229-008-9324-8.pdf.